A Complete Bibliography of Publications in North American Journal of Fisheries

Management: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

05 June 2024 Version 1.21

Title word cross-reference

Y Y [ACRB22]. F_{1} [BDHP20]. × [HBT^{+}21, NLK21]. W_{s} [BBFQ21, BBMC23, GBJ21, MRW^{+}24].

10-Year [LKP^{+}21]. 12-mm [TRB^{+}21]. 19 [AWRL23, LTL^{+}23, TBG^{+}22].

2021 [Ano22a, SK23].

4-Year [BOB^{+}21]. 41 [Ano22a]. 4th [CP21a].

8- [TRB^{+}21].

9- [TRB^{+}21].

Abandoned [MSP^{+}22]. Abiotic [AWF23, CH22, PO22, BHB^{+}23]. above [JJD^{+}23]. Absolute [GRM^{+}21]. Abundance [BCS^{+}21, BGH^{+}20, BHT^{+}21, BWM22, CBLW21, CH22, CWB^{+}21, CDBC22, GRM^{+}21, GGWV22, HQWD20, HDQ22, Kep21, MBW21, MSLP22, Mic20a, MSV^{+}21, MST22, PS21, PA21, PREBK22, RAHS20, SSS^{+}22, SMP^{+}21, EPD^{+}24]. Acceleration [LDH^{+}22]. Acceleration-Logging [LDH^{+}22]. Accessibility [CLTJ21]. Acclimated [CLBL22]. Accuracy [Har20, MCB20, MS22, ORL^{+}22]. Accurately [EMRK20, LSB^{+}21, SHI22]. Achieve [MBW21, SLZ^{+}20]. Achieving [PST^{+}20]. Acidic [HKPP20]. Acknowledgments [Ano20g, Ano21i, Ano22h, Ano23h]. Acoustic [CHJ^{+}22, CCL^{+}22, HYM^{+}22, KHL^{+}24, MWR21, SGK^{+}22, GWRF24, SLD^{+}23]. across [PFGE23, PHW^{+}23, RSM^{+}23, WMB23]. Action [BGC^{+}21]. Actions [Mun22, STB^{+}21]. Active [MSL^{+}20, RPEF^{+}23]. Actively [WSK^{+}22]. Activities [HSW21]. Activity [YZJ^{+}21, SLD^{+}23]. Actual [MSP^{+}22]. Acute [GRA^{+}22]. Adaptation [BBB^{+}21]. Adaptive [TJ22]. addition [SSR^{+}24]. additives [LHB^{+}24]. Adfluvial [MBW21]. Adjust [PHG^{+}23]. admixture [LOV^{+}23]. Adult [BR21, CVD^{+}20, DPR^{+}20, DRPW21, EPB^{+}22, FM20, HRS^{+}20, IPZK21, KNC^{+}20, KES^{+}21, KJB^{+}22, MR21b, PTFZ23, PST^{+}20, RCD22, SFFW21, WRH^{+}21, CEB^{+}23, Gep23, HLJ^{+}23, LCVS23, MMSJ23, YBB^{+}23]. Advanced [TPPC20]. Advanced-Fingerlings [TPPC20]. Advancement [MS21a]. Advances [SGK^{+}22, SDB^{+}20]. Aerial [KDM^{+}23, TVG^{+}21]. Affect [BWM22, HRS^{+}20, KJB^{+}22, MSP21, CBKH23, KPD^{+}24, SWA^{+}24, SMB23]. Affecting [AWF23, HAP^{+}22, HEP^{+}23, KMS21, PO22, SMP^{+}21, ZWBS21]. Affiliated [CSQ^{+}21]. after [BAH^{+}20]. Age [AWF23, BGH^{+}20, CLM20, DAS20, DKEI20, DMBC23, ESCI21, GBJ^{+}22, GW20, GW21, HNM^{+}22, Har20, HMQM22, HGSS20, HJBP20, HMO^{+}21, KBC^{+}20, KSPP22, LCF20, LSSB23, LMO21, MVH^{+}22, MNA^{+}20, Mic20b, MCB20, ORL^{+}22, PG20, RCD20, SB21a, SHI22, SSPL20, WWK^{+}23, KHBF24, PHW^{+}23]. Age-0 [AWF23, BGH^{+}20, DKEI20, ESCI21, GBJ^{+}22, GW20, GW21, KSPP22, LCF20, SSPL20, WWK^{+}23]. Age-Based [HNM^{+}22]. Age-Specific [RCD20]. Age-Structured [HMO^{+}21]. Agencies [GDS^{+}22]. Agents [CGD^{+}21]. Ages [DMBC23, IPZK21, RCD22]. Aggregate [NMBM21, ZLM21]. Aggregations [WGKB20]. Aging [PSW23]. Agricultural [NHC^{+}20]. Aid [HCP^{+}23]. Alabama [LEHP21, AP20, BS21, ECC22, LSB^{+}21, LEHP21, PJD^{+}21, SDL^{+}23, SDR24]. Alaska [CLSM20, CFHS22, HG22, JWG^{+}21, MBP^{+}22, SFFW21]. Alaskan [AFBRS22, CFCN20]. Albemarle [MWR21]. Alberta [MWASW20, PCM^{+}21, RM21]. Alewife [SLD^{+}23, SBB23, WSG^{+}23]. Alligator [AKC20, AP20, BS20, DAS20, HDH^{+}20, KDB^{+}20, LBH21, LSP20, SDB^{+}20, SBD^{+}20, SLP20, SSSP22, TLSP20]. along [AWG^{+}20, LCIL23b]. Alter [SDL^{+}23]. Altered [HSP21]. Alternative [BTZ^{+}21, PS21]. Amberjack [GRM^{+}21]. America [HRK^{+}22, LCM^{+}23, WAKO21]. American [SGM^{+}20, BIE^{+}22, CBLW21, CVD^{+}20, CWB^{+}21, MWR21, MR21a, MGMP21, MBW21, MWMH22, MSD^{+}22, SWJ^{+}21]. among [AWO21, EPD^{+}24, GvRSS21, HSP21, HMO^{+}21, MML^{+}22, WA23]. Anadromous [CLM20, TCB^{+}21]. Anal [MCB20]. Analysis [CFCN20, DH21, FLW21, HRS^{+}20, KFG22, KDB^{+}20, MBP^{+}22, Mun22, RGHF23, SSPL20, SHFP20, ZJL^{+}21, vPM20, ACGF23, LOV^{+}23]. Analyzing [SHFP20]. Anchor [SPSH20]. Ancient [LHB^{+}23]. Anesthesia [LL21]. Anesthetic [KS20]. Angled [CRH^{+}22, TCL^{+}21, MLB^{+}23]. Angler [BM20, CLTJ21, CM22, ERT21, GET22, GGW21, JOH^{+}22, KQS^{+}20, MAB^{+}20, Mun22, PWLH21, PHG^{+}23, SDF^{+}21, TPC^{+}22, TBIJ21, vPM20, NKG23, SSM^{+}24, YBH24]. Anglers [GvRSS21, HBT^{+}21, KBDA21, MML^{+}22, STB^{+}21, RR23]. Angling [BDC22, CR23, ERT21, KS21, MKC^{+}23, MDA21, SLG^{+}23, SGH^{+}21, SCP^{+}23, SMB23]. Anguilla [SLL^{+}21]. Animals [HIF21]. Annual [ASDW23, GLSF21]. Annuli [BS20]. Anodic [BD22]. Antenna [FPR20]. Antennas [OTBB23, ZWBS21]. antipodarum [GET22]. Appalachian [HKPP20, SHWW22]. Apparent [FDG^{+}21]. Application [CSLW22, CGD^{+}21, FLW21, JSH^{+}21, KDM^{+}23, RMB^{+}20, CEC^{+}24, LKD23, RAH^{+}23]. Applications [DMBC23]. Applied [FB20, RAB^{+}20]. Applying [RJ22]. Apportioning [HL23]. Approach [BBB^{+}21, CVD^{+}20, HSTH20, ZJL^{+}21, PTFZ23]. Approaches [CLM20, FJH^{+}22, PS21, SLE21, SCW^{+}24]. Aquatic [BBO20, CBS21, GK21, HIF21, HMB21, KR23, HWCS24]. Arctic [GBJ21, HYM^{+}22, MRK^{+}22, MRM^{+}21, RM21, SYHF24, SCDS23]. Areas [HGM22, SFFW21]. Arizona [BM20, HSS^{+}20, HGMR21, NLB21, SPY^{+}20]. Arkansas [RKP^{+}22, EP21]. armored [AMR^{+}23]. Arrays [ZWBS21]. Artificial [BCW^{+}21, JCB^{+}21, PBD^{+}22]. Artisanal [QBA21]. Assay [Ano22a, WPGC21]. Assemblage [HGMR21, Rac21, SBF^{+}23]. Assemblages [CSQ^{+}21, RMWK23, SDL^{+}23, YZJ^{+}21]. Assess [AKC20, MLB21]. Assessed [DAS20, LTB^{+}22, MRK^{+}22]. Assessing [ERCB22, Har20, HBBS22, MBF^{+}23, NHC^{+}20, PTCZ22, SDR24, SKWR23, SFFW21, SW22]. Assessment [BAH^{+}20, FM20, FB20, GKRH21, HCC^{+}21, HSW21, HCM^{+}22, RAB^{+}20, RLS^{+}23, RRL^{+}21, SLE21, WD21, BSR^{+}23, HCO^{+}24]. Assessments [CP21b]. Assign [LLA^{+}22]. Assigning [DMBC23, RMB^{+}20]. Associated [BFLI22, GRA^{+}22, HTH^{+}21, HJBP20, LSB^{+}21, PREBK22, SVS^{+}21]. Associations [CP21b, HDQ22, RMR^{+}22, WD21]. Atlantic [OS20, AS20, BPL^{+}21, BSKSK24, FHK23, GSML23, HLDK21, HWCS24, HL23, Kep21, LOC23, MGNP22, NLW^{+}20, PA21, Rac21, RPJ^{+}21, SCW^{+}24, SCL^{+}20]. Attitudes [MAB^{+}20]. Attributes [DH21]. augmentation [ROP^{+}23]. August [Ano22a]. Authentication [ZJL^{+}21]. Automated [AFBRS22, SLD^{+}23]. Autonomous [DDT^{+}20]. Availability [SKWR23, WW20]. Avian [EPB^{+}22, HEP^{+}23, PFGE23]. avoidance [LOC23]. Avoiding [PBEK23].

Backcross [LMA23]. Backpack [LCVS23]. Bag [CM22, MSW22]. bait [SPM^{+}24]. Baited [TPC21b]. Baitfish [MML^{+}22]. Balance [vPM20]. Bald [CJN^{+}21]. Bank [DHWP21]. Bar [SPSH20, SCDS23]. Barotrauma [ERT21]. Barotraumatized [RSM^{+}23]. Barrier [TM21]. Bartram [PJD^{+}21]. Based [BHCC23, CLM20, DAS20, DH21, DC22, GWMP23, HNM^{+}22, HDP^{+}22, HJBP20, HBBS22, HJ21, JSH^{+}21, LSSB23, NLW^{+}20, RCD22, SL21, SR20, TBIJ21, WSK^{+}22, BSKSK24]. Baseline [KLR^{+}21]. Bases [SPY^{+}20]. Basic [NRJ^{+}20]. Basin [ELCM22, HTH^{+}21, JDO^{+}20, LSB^{+}21, LGQ^{+}21, CFGZ23, LJB^{+}24, WMB23, Ano23g, DTK20, ESCI21, FCMT21, HHH^{+}22, HHH^{+}23, KJB^{+}22, PE21]. Basins [CP21a]. basinwide [CFGZ23]. Bass [AZL^{+}21, ARK^{+}21, BTZ^{+}21, BMAM22, BDHP20, CLTJ21, CAZ^{+}21, EP21, ERT21, HDP^{+}22, HL23, KR24, LTB^{+}22, LSB^{+}21, LEHP21, LMA23, MSW22, MQK^{+}20, PJD^{+}21, RRB20, RSM^{+}23, SDL^{+}23, SR20, SW22, TPPC20, WD21, YBH24, CEC^{+}24, LHB^{+}24, SSM^{+}24, SWA^{+}24, KKMA23]. Batch [DWGR21, HKPP20, LCIL23a, LCIL23c]. Bay [FB20, HJB^{+}21, IDH^{+}23, SHI22, ZK22, FNT21]. Bayesian [ORWS21]. be [HDH^{+}20, LCVS23]. Bear [HDQ22, HDT^{+}22]. Beaver [RH20]. Been [KMS21]. before [BGH^{+}20]. Behavior [BLS23, CLM20, LTB^{+}22, MGNP22, Mun22, WWBM20, WW20, NKG23, PK24]. Behavioral [CRH^{+}22, KSPP22, LHB^{+}24]. Behaviors [MAB^{+}20]. Benefit [MKC^{+}23, SCMJF20]. Benefits [MDA21]. Benthic [PKTZ20, NMS^{+}23]. benthic-spawning [NMS^{+}23]. Bering [GLSF21]. Better [GW20]. between [CLSM20, CSQ^{+}21, CHJ^{+}22, GDS^{+}22, HGH^{+}23, KNK22, LMA23, PO24, RSO23, ZAK^{+}23]. Beverages [TCL^{+}21]. Beyond [PC20]. Bias [DH21, Har20, PSW23, ZSLG21, BSKSK24]. Biases [HJBP20]. Biennial [GLSF21]. Big [HG22]. Bigger [GW20]. Bighead [AAW^{+}23, LLA^{+}22]. Bigheaded [AWRL23, LTL^{+}23, LCIL23a, LCIL23b, LCIL23c, RPEF^{+}23, RZK^{+}21, SMCEF^{+}23]. Bighorn [HSH21]. Bight [OS20]. Bigmouth [ASDW23, PSW23, SWM22]. Bin [HJBP20]. Bioelectrical [CFCN20]. Bioenergetics [GW21, TCM20]. Biofilm [HSW21]. Biologgers [LTB^{+}22]. Biological [CGD^{+}21, SSF^{+}23, UHB^{+}23]. Biology [RG20, ZZWX20]. Biota [KR23]. Biotic [AWF23, BHB^{+}23, CH22, PO22]. Black [AWO21, Bar21, BFM^{+}22, CLTJ21, EP21, LTB^{+}22, LSB^{+}21, Mic20b, MSL22, RRB20, RSM^{+}23, SDL^{+}23, WAKO21, KHBF24, SSM^{+}24, KKMA23]. Bleeding [TCL^{+}21]. Blue [CSLW22, ERCB22, FNT21, HJB^{+}21, HBT^{+}21, MDA21, MSZ^{+}23, MS21b, MSP21, NLK21, NKK21, NF21, NMBM21, OTW^{+}22, PAC22, PTCZ22, SB21b, WM21, BSKSK24, HGMR21, NKG23]. Blueback [SLD^{+}23, HCO^{+}24]. bluefin [GSML23]. Bluegill [BDC22, GGW21, PSS21, PK24]. Bluehead [RHUM^{+}23]. Boat [CJP^{+}21, DDT^{+}20, MS21b, PTCZ22]. Boats [BD22]. Bodied [DWSM21, MB21]. Body [CHL22, HDH^{+}20, NLK21, FHB^{+}24]. Bolstering [Mic20a]. Bomb [DAS20]. Bonneville [HDQ22, HDT^{+}22]. Bottlenecks [GBJ^{+}22]. Bottom [CTJS^{+}22, GC21]. Bottom-Trawl [CTJS^{+}22]. Boulder [LKD23]. Bound [WRH^{+}21]. Bowfishing [MSSS23, YSS22]. Branch [KHB22]. Branched [BBKJ21]. Breakaway [FPR20]. Brief [HMB21]. Bringing [LHWS23]. British [HMO^{+}21, RPJ^{+}21]. Broadband [DS20]. Broaden [KBC^{+}20]. Broadscale [RRL^{+}21]. Bronx [CWB^{+}21]. brood [PK24]. Broodfish [AP20]. Broodstock [HHH^{+}22, HHH^{+}23, Ano23g]. Brook [Ano22a, ACRB22, HSF^{+}22, HKPP20, KLR^{+}21, LL21, MKK^{+}21, RH20, RRL^{+}21, SCP^{+}23, SG22, SMM^{+}22, TCM20, WPGC21, LOV^{+}23]. Brown [BR21, DWGR21, HW22, Mun22, RCD20]. Brunswick [ARK^{+}21, HL23]. Buffalo [ASDW23, LSSB23, PSW23, SWM22]. Bull [AS20, BVP^{+}21, BDG^{+}23, CHJ^{+}22, KSP^{+}24, LGOM22, RSK^{+}21]. Bullhead [Bar21, SVS^{+}21]. Burbot [BQR^{+}20, RHFU24]. Burden [HDH^{+}20]. burrowed [HLSW23]. Bycatch [HLDK21, MRK^{+}22, MSL22, PSS20, TJ22, TSJ20, LOC23].

C [Ano22a]. calcified [KHBF24]. California [ACGF23, Bar21, Boo20, DRPW21, KFG22, RG20]. Camera [HL23]. Cameras [DDT^{+}20, ECC22]. can [LCVS23]. Canada [CVD^{+}20, ARK^{+}21, HYM^{+}22, HL23, MNA^{+}20, RPJ^{+}21, RM21]. Canadian [MRK^{+}22]. Canals [MLB21]. Canyon [HSS^{+}20]. Capacity [BDG^{+}23, PC20, TCM20]. captively [CBKH23]. Capture [BLS23, CMBR20, DHWP21, GC21, HMP22, HSF^{+}22, LHR^{+}20, LGOM22, MSW22, OMAA22, RJ22, SGM^{+}20, dVB20a, dVB20b]. Captured [CJN^{+}21]. Carbon [FMM20]. Carbonated [TCL^{+}21]. Carcasses [PKTZ20]. Caribbean [CRNRFL24]. Carolina [RRB^{+}22, CKB^{+}21, KLR^{+}21, MWR21, RR23]. Carp [AAW^{+}23, AWRL23, ASDW23, CSWP23a, CSWP23b, CBS21, CCL^{+}22, DS20, FNB^{+}22, GWMP23, HBV^{+}21, HBR^{+}23, HCP^{+}23, LTL^{+}23, LLA^{+}22, LCIL23a, LCIL23b, LCIL23c, MSM21, RGMH20, RLS^{+}23, SMCEF^{+}23, SWM22, SHFP20, SM22, SGM^{+}20, WDPH22, WWK^{+}23, YBB^{+}23, ZZWX20, KHBF24, RAH^{+}23]. Carps [PBK^{+}23, RPEF^{+}23, RZK^{+}21]. Carrying [TCM20]. Cascade [WMJ^{+}24]. Case [CBC^{+}20, FNT21, HDH^{+}20, KBC^{+}20, CFGZ23]. Catch [BCS^{+}21, BM20, BBKJ21, BWM22, BDC22, CM22, ERT21, HG22, KQS^{+}20, MdAM21, MDBC21, MVH^{+}22, MSP21, Mun22, PWLH21, PSS20, PG20, RGMH20, SLG^{+}23, SB21b, SWM22, SGH^{+}21, LJB^{+}24, RR23, SSM^{+}24, SMB23]. Catch-and-Harvest [Mun22]. Catch-and-Release [BCS^{+}21, BBKJ21, BDC22, ERT21, MDBC21, SLG^{+}23, SGH^{+}21, LJB^{+}24, SMB23]. Catch-at-Age [MVH^{+}22]. Catch-Curve [MdAM21]. Catchability [HBT^{+}21, Pan21]. Categories [BBFQ21, BB21a, BBMC23, MRW^{+}24]. Catfish [BTZ^{+}21, BBKJ21, BOB^{+}21, CFPSW21, DWSM21, DHWP21, DSW^{+}24, ESCI21, FNT21, HTH^{+}21, HSP21, HGH^{+}21, HJB^{+}21, HBT^{+}21, Kep21, KS21, MDA21, MS21a, MS22, MSZ^{+}23, MS21b, MSP21, MK21, NLK21, NKK21, NF21, NMBM21, ORWS21, PKNS21, PTCZ22, SB21a, SL21, STB^{+}21, SB21b, Sin21, SMP^{+}21, SPSH20, TPC21a, TPC21b, WM21, AMR^{+}23, CHO24, SBH^{+}23, NKG23]. Catostomids [BBP20]. Caught [GC21]. Cause [PFGE23]. Cause-Specific [PFGE23]. caused [HLSW23]. CECs [HBTE22]. Ceded [LLE22]. Central [GGWV22, GKT^{+}22, HLM^{+}23, RMR^{+}22, WW20]. Centrarchids [GRA^{+}22]. centrifugal [WPM^{+}23]. Centropomus [AWG^{+}20]. Ceratonova [RAHS20]. cf. [PJD^{+}21]. Challenges [HIF21, SK23]. Chamber [DS20]. Champlain [HMQM22]. Change [BIE^{+}22, GRA^{+}22, WZG22]. Changes [CLRL22, CBS21, FHR21, HRK^{+}22, MSLP22, RSO23, SVS^{+}21, YZJ^{+}21, ZK22, GBM23]. Changing [BKE22, FHK23, HBBS22]. Channel [BTZ^{+}21, CFPSW21, DSW^{+}24, ESCI21, HGH^{+}21, HBT^{+}21, Kep21, NLK21, NKK21, NMBM21, SL21, Sin21, SPSH20, TPC21a, TPC21b, UHB^{+}23, ROP^{+}23]. Char [HYM^{+}22, MRK^{+}22]. Characteristic [LMA23]. Characteristics [BKM21, CLTJ21, GKRH20, HSP21, HQWD20, HJBP20, KNK22, LLE22, MBCH22, NF21, RCR21, SHWW22, SHFP20, TPC21a, ZLM21]. Characterization [AWG^{+}20]. Characterize [KDB^{+}20, TCB^{+}21]. Characterized [BWM22]. Characterizes [WSK^{+}22]. Characterizing [CP21b, CFHS22, SMM^{+}22]. Chase [MS21b, PTCZ22]. Chase-Boat [PTCZ22]. Chasing [SGH^{+}21]. Chemical [FHA^{+}21]. Chemistry [AWRL23]. Chesapeake [FNT21, HJB^{+}21]. China [CSQ^{+}21, LGQ^{+}21, SLL^{+}21, YZJ^{+}21, ZZWX20]. Chinese [ZJL^{+}21]. Chinook [Ano23g, BHT^{+}21, BB21b, BSC^{+}21, CEH^{+}20, CLBL22, CFCN20, CDBC22, ELCM22, FHR21, FMM20, FTLR23, GWRF24, HAP^{+}22, HLM^{+}23, HJ21, HHH^{+}22, HHH^{+}23, JSH^{+}21, KSJ^{+}20, NKC^{+}23, PFGE23, PO24, PST^{+}20, PC20, RJ22, RAHS20, TRB^{+}21, WRH^{+}21, ZAK^{+}23]. Chionoecetes [GLSF21, GBM23]. Chip [MB21]. Choice [TPC^{+}22]. Chub [BBFQ21, DEO20, HSS^{+}20, MDHQ22, MSD^{+}22, RBFQ20, SPY^{+}20, EFK^{+}23, MBF^{+}23, McK23, PAE^{+}23, PKSG23, WMB23, WA23, SBW^{+}23]. chubs [PBEK23]. Chubsucker [BB21a]. Chum [JWG^{+}21, MBP^{+}22]. Cisco [FB20, VF21, YDM^{+}20]. Clair [CBB^{+}23, JCB^{+}21, UHB^{+}23, BHT^{+}20]. clam [HA24]. Class [DNW^{+}22]. Classified [BBO20]. Cleaning [BD22]. Clear [PKNS21]. Climate [FHK23, CDG^{+}23]. climate-related [CDG^{+}23]. Clips [LCF20]. Clock [RMWK23]. clusters [RR23]. Co [GHA^{+}23, SCMJF20]. Co-Occurrence [SCMJF20]. Co-Occurring [GHA^{+}23]. Coarse [GWM21, SSR^{+}24]. coast [KKMA23]. Coastal [BPT21, CLM20, DRPW21, EMRK20, KBDA21, LFS20, MRK^{+}22, MSL22, RG20, LHPS^{+}24, MBP^{+}22]. Cod [SCL^{+}20]. coded [CEB^{+}23]. Coexistence [SCMJF20]. Coho [BBB^{+}21, FLW21, KHL^{+}24, SKBB24, SFFW21]. Cold [CRH^{+}22, MWASW20]. collect [LCVS23]. Collected [SB21b]. Collection [CDB22, HRS^{+}20, CRNRFL24]. Collections [AP20, WTW^{+}23]. Colonization [KHD^{+}22]. Colorado [SPY^{+}20, AWF23, HSS^{+}20, HUTB21, RFK^{+}20, RKP^{+}22]. Columbia [FCMT21, MST22, RPJ^{+}21, SG21, Ano23g, BHS^{+}23, BHT^{+}21, HLJ^{+}23, HHH^{+}22, HHH^{+}23, HMO^{+}21, PO24, PE21, PST^{+}20, SWTH21]. Combination [SGK^{+}22]. Combining [CDG^{+}23, HL23]. Commercial [HYM^{+}22, NLW^{+}20, PG20, SCL^{+}20, TSJ20, WM21]. Common [ASDW23, DS20, MSM21, PG20, SWM22, SM22, SGM^{+}20]. Communicate [PG20]. Communities [KMS21, Mic20a, WHA23]. Community [CM22, SVS^{+}21, VF21, KHL^{+}24]. community-run [KHL^{+}24]. Compact [WSG^{+}23]. Comparative [LL21]. compare [SCW^{+}24]. Comparing [Sin21]. Comparison [ERT21, FM20, HCS^{+}21, LTB^{+}22, LSSB23, MdAM21, NLK21, PSS21, RRL^{+}21, RPEF^{+}23, SB21a, SKBB24, DSW^{+}24]. Comparisons [CFCN20, PO24]. Competing [FNT21]. competition [WA23]. Complex [AWG^{+}20, BCW^{+}21, HYM^{+}22, MW20, SG21, LOV^{+}23]. Complexity [GWM21]. Complicated [HRK^{+}22]. compliments [WHA23]. Composition [BHCC23, HHH^{+}22, HHH^{+}23, Ano23g]. Concentration [CWB^{+}21]. Concentrations [MKG^{+}22]. Concern [HBTE22]. Concert [MCB20]. Condition [MLB^{+}23, SSH22, EPD^{+}24, FHB^{+}24]. Conditioning [FDG^{+}21]. Conditions [BFLI22, HTH^{+}21, HSP21, OS20, QBA21, RSO23, RBFS20, ZK22]. Conductivity [HCS^{+}21]. Confined [SM22]. Confirms [AWO21]. Connected [DBS21]. Connectivity [COT20]. Consequences [CAZ^{+}21]. Conservation [BDG^{+}23, CKB^{+}21, KLR^{+}21, LHWS23, LHB^{+}23, LGQ^{+}21, McK23, RG20, SDB^{+}20, SBW^{+}23, vPM20, PBEK23, SPM^{+}24]. Considerations [BPT21, SSF^{+}23, TPC^{+}22]. Considering [MSV^{+}21]. consistency [EPD^{+}24]. Conspecifics [DPR^{+}20]. Constraints [TBIJ21]. construction [SFM23]. Consumption [GW21, HGH^{+}21, MMSJ23]. Contact [RM21]. Contaminant [HDH^{+}20]. Contaminants [HBTE22]. Contemporary [HMQM22]. Contest [Mun22]. Context [NMW^{+}21]. Contrasting [WW20]. Contribution [NLK21]. Contributions [GDS^{+}22, HDP^{+}22, HDH^{+}20, KDB^{+}20, HS24]. Control [CGD^{+}21, RLS^{+}23, TPA^{+}21, CEC^{+}24]. controlled [HCO^{+}24]. Controls [FHA^{+}21, JDO^{+}20]. Cook [ZLM21]. coosae [PJD^{+}21]. copepods [SMB23]. Correct [DH21]. Correlates [RAHS20, SBD^{+}20]. Cost [BCW^{+}21, HSTH20]. Cost-Effective [BCW^{+}21, HSTH20]. Coster [Ano22a]. Costs [CLRL22, CR23]. Count [AFBRS22, PREBK22, SKBB24, SDF^{+}21]. Counting [EMRK20, HL23, KHJS22]. County [ZLM21]. Coupling [GET22]. Cove [RMWK23]. Cover [GKRH20]. Coves [PBD^{+}22]. COVID [TBG^{+}22]. COVID-19 [TBG^{+}22]. CPUE [GBM23]. Crab [ERCB22, GLSF21, OTW^{+}22, PAC22, ZJL^{+}21, BSKSK24, GBM23]. Crabs [OS20, OTW^{+}22]. Crappie [BFM^{+}22, HN21, MKKN22]. Crappies [Mic20b]. Crayfish [BOHA20]. Creates [HKPP20]. Creek [CJN^{+}21, LKD23, LEHP21, WCSL22]. Creel [DDT^{+}20, ECC22, NLK21, NMBM21]. Criteria [NLB21, RRL^{+}21]. Critical [BMAM22, TVG^{+}21, LHB^{+}23, RHUM^{+}23]. Cues [SM22]. Culling [CAZ^{+}21, GGW21]. culture [MMJW23]. Culvert [KHD^{+}22]. Culverts [RRL^{+}21]. Cured [SMM20]. Current [BD22, SDB^{+}20]. Curve [MdAM21]. Customer [STB^{+}21]. Customized [MLB21]. Cutthroat [BPT21, CLM20, HQWD20, HDQ22, HDT^{+}22, LKD23, LFS20, MDHQ22, MBW21, NMW^{+}21, SCN^{+}21]. Cycle [BBB^{+}21, KUV^{+}23, PHW^{+}23]. Cycles [GLSF21]. Cytological [FJH^{+}22].

D [KBDI21]. D-Frame [KBDI21]. Dace [RBFS20]. Dakota [BSKM20, DNW^{+}22, SL21]. Dam [HSTH20, KES^{+}21, LTL^{+}23, LBH21, PTFZ23]. Dammed [CBLW21]. Dams [PE21, SWTH21, JJD^{+}23]. Data [ASDW23, BW22, DC22, DWGR21, EMRK20, FB20, JOH^{+}22, LHR^{+}20, LMO21, MdAM21, NMBM21, NMW^{+}21, ORWS21, PWLH21, Rac21, RJ22, SKA21, SB21b, SLZ^{+}20, TPC^{+}22, CDG^{+}23, CRNRFL24, KPD^{+}24, RH23]. Data-Limited [BW22, ORWS21, SLZ^{+}20]. Dating [DAS20]. DC [CMBR20]. Deal [TJ22]. Decades [MWASW20, MS21a]. Decision [BDG^{+}23, vPM20, McK23]. decisions [KPD^{+}24]. Declines [NMW^{+}21]. Declining [HRM^{+}22, SDWM21]. Decontamination [GET22]. decreases [RHUM^{+}23]. Deepwater [LMN^{+}20, MLB21, RMB^{+}20]. Delaware [SMP^{+}21]. Delay [BPL^{+}21]. delayed [HWCS24]. Delta [AP20, BSC^{+}21]. Demand [CLRL22]. Demersal [RRB^{+}22]. Demographic [AWG^{+}20, AWO21, SSS^{+}22]. Demographics [SNLD21, WDPH22, TO23, YBH24]. Demography [WAKO21]. Denil [SBB23]. Densities [HLM^{+}23]. Density [CBS21, EMRK20, GKT^{+}22, KQS^{+}20, KDO^{+}22, RCD20, TPC21b, BHS^{+}23, RHUM^{+}23]. Density-Dependent [CBS21]. Dependence [GKT^{+}22]. Dependent [CBS21, SDWM21, SW22, ZSLG21]. Depends [HRK^{+}22]. Deposition [GWM21]. Depredation [RRB^{+}22]. Depths [RSM^{+}23]. Derived [ORL^{+}22, RCD22]. Descender [RRB20, RMB^{+}20, RRB^{+}22]. Desert [HGH^{+}21, HGH^{+}23]. Desiccation [TPA^{+}21]. Design [SFM23, RAH^{+}23]. Designed [WSG^{+}23]. Designing [BCW^{+}21]. Designs [CHS^{+}22]. Despite [MWASW20]. Detect [ZWBS21]. detecting [SKBB24]. Detection [PE21, ZWBS21, KPD^{+}24, SLD^{+}23]. Determination [FJH^{+}22, MSD^{+}22]. Determine [AWRL23, KKMA23, SYHF24]. Determining [WKM^{+}21]. Detroit [CBB^{+}23, JCB^{+}21]. Develop [BDG^{+}23]. Developing [BVP^{+}21, HW22, CEB^{+}23]. Development [BB21b, CSLW22, FPR20, KLR^{+}21, LSP20, NLB21, Owe21]. Devices [RRB20, RMB^{+}20, RRB^{+}22]. dewatered [HLSW23]. dewatering [LHB^{+}23]. Diameter [STP^{+}21]. DIDSON [HLM^{+}23]. Diel [RGMH20]. Diet [CDB22, GHA^{+}23, RWS^{+}21]. Dietary [FHR21]. Diets [STP^{+}21]. Differ [NF21]. Differences [AWO21, CEH^{+}20, HBTE22, LMA23, SHWW22]. Different [CAZ^{+}21, LTB^{+}22]. Differential [OTW^{+}22]. Differentiate [BDHP20]. Differentiates [SGM^{+}20]. Differentiation [RGHF23]. differs [HGH^{+}23]. Digested [STP^{+}21]. Digital [AFBRS22, DGJ^{+}22, ECC22, JOH^{+}22]. Dimensions [KBDA21]. Dimorphism [LKP^{+}21, AMR^{+}23]. diode [AJHB23]. Dip [KBDI21]. Diphenyl [HBTE22]. Direct [BD22, CLBL22, RAB^{+}20]. Direct-Current [BD22]. Direct-Stocked [CLBL22]. Directions [LCM^{+}23]. Discard [SSSP22, SCL^{+}20, WKM^{+}21]. Discharge [SPP^{+}21, SKGL22]. Discontinuities [BS20]. Discriminating [Ano22a, WPGC21]. Disease [ESF23, RAHS20, TPA^{+}21]. Disease-Resistant [ESF23]. Disparate [SPY^{+}20]. Dispersal [AZL^{+}21, BS21, DKEI20, MGNP22]. Disrupt [HSW21]. Distance [SGH^{+}21, CR23]. Distinct [BMAM22]. distinctiveness [EFK^{+}23]. Distribution [BHT^{+}21, CBLW21, CSQ^{+}21, FCMT21, HQWD20, MDHQ22, OBR^{+}20, PJD^{+}21, SLL^{+}21, BHS^{+}23, CDG^{+}23]. Distributional [SCMJF20]. Disturbance [PJD^{+}21, RSO23]. Diversity [CLSM20, CKB^{+}21, DBCS20, LGQ^{+}21, ACGF23]. DNA [BSR^{+}23, CBC^{+}20, CWB^{+}21, KSJ^{+}20, SKBB24, WHA23]. Do [HSW21, KJB^{+}22, LHB^{+}24, TCL^{+}21]. Does [BWM22, CBKH23, HSF^{+}22, YDM^{+}20, SMB23]. Domesticated [SG22]. donor [PO24]. Dosage [KS20]. Double [FPR20]. Double-Breakaway [FPR20]. Downstream [MGNP22, SPP^{+}21]. Drainage [RG20]. Drawdown [MVM^{+}23]. Dredge [WKM^{+}21]. Dreissenid [WZG22]. Driven [WZG22]. Drivers [Boo20, JOH^{+}22]. Drives [BGC^{+}21]. Drought [BBB^{+}21, NMS^{+}23]. Drum [AWO21, MSL22, WAKO21]. Dry [BBC21, WMB23]. Dual [EMRK20]. Dual-Frequency [EMRK20]. duality [NMS^{+}23]. Dummy [CCL^{+}22]. Duplex [OTBB23]. Duration [MS22, MS21b, SB21b]. during [GBM23, KSPP22, TM21, ZAK^{+}23]. Dwelling [RCD20]. Dwindling [KUV^{+}23]. Dynamics [BQR^{+}20, BW22, CFPSW21, DBS21, HDT^{+}22, HJB^{+}21, Kep21, MQK^{+}20, RSK^{+}21, SMCEF^{+}23, SL21, SSH22, SG22, KR24, MDD^{+}23].

Eagle [CJN^{+}21]. Early [AZL^{+}21, AWRL23, BBP20, DNW^{+}22, LSP20, PO22, WWK^{+}23, AJHB23, KHL^{+}24, WA23]. Early- [AZL^{+}21]. East [RSK^{+}21]. Eastern [WCSL22, AZL^{+}21, CTJS^{+}22, DNW^{+}22, GLSF21, WSK^{+}22]. Ecological [BGC^{+}21, MBCH22, SLE21, WZG22]. Ecology [BBP20, RBFQ20, SSPL20, ZZWX20, WTW^{+}23]. Economic [Pan21]. Ecosystem [BIE^{+}22, HRK^{+}22, SBF^{+}23, CHO24]. Ecosystems [KBDA21]. Edge [AWG^{+}20]. Eel [CBLW21, CWB^{+}21]. Eels [SLL^{+}21]. Effect [BD22, CLRL22, CMBR20, CM22, FB20, GKRH20, MKKN22, RZA^{+}21, RH20, SB21b]. Effective [BCW^{+}21, COT20, HSTH20, KS20, TVG^{+}21, WB24]. Effectiveness [KBDI21, RRB20, RSM^{+}23, TCB^{+}21, dVB20b]. Effects [BBC21, BIE^{+}22, BHT^{+}20, CCL^{+}22, CHL22, ESCI21, FFG20, GRA^{+}22, HRK^{+}22, HBTE22, KR24, KR23, LSP20, MSW22, MPM21, MS21b, NKG23, OTW^{+}22, PSS20, RCD20, RZK^{+}21, SL21, SLG^{+}23, SSH22, SPY^{+}20, SW22, TBG^{+}22, WW21, ZMA20, HFJ23, HWCS24]. Efficacy [DHWP21, HCP^{+}23, LKD23, OMAA22, TPA^{+}21, WGKB20, WGBK22, RAH^{+}23]. Efficiency [BOHA20, CMBR20, LGOM22, MS21b, SBB23, WCSL22]. Efficient [SMM20, TVG^{+}21]. Effort [DDT^{+}20, ERCB22, LMA23, TBG^{+}22, SDR24, SSM^{+}24]. Efforts [CBC^{+}20, HBV^{+}21, HCP^{+}23, SGK^{+}22]. Egg [AFBRS22, CEH^{+}20, GWM21, HRS^{+}20, MKG^{+}22]. Eggs [GWMP23]. Elastomer [BR21, LFS20, MB21]. Electric [LL21]. Electrical [BD22]. Electrified [RGMH20]. Electrofisher [MLB21]. Electrofishing [ASDW23, BOHA20, BD22, CDB22, DKEI20, HSW21, HCM^{+}22, HSF^{+}22, MS22, MSZ^{+}23, MS21b, MSP21, PTCZ22, SB21b, SWM22, HLSW23, LCVS23, SDR24, YBB^{+}23]. Electronic [MAB^{+}20, RJ22, GSML23]. Elemental [LMA20]. Elevated [KSPP22, SSS^{+}22]. Elevation [FFG20, KJB^{+}22]. Elicit [SPGP21]. Eliminate [TSJ20]. Eliminates [KDM^{+}23]. Elk [WRH^{+}21]. Embryo [HSF^{+}22]. Embryos [PKTZ20, SCN^{+}21]. emerged [HLSW23]. Emerging [AWRL23, BDC22, HBTE22]. Emigration [BPT21, MR21b, MW20, SPP^{+}21]. emitting [AJHB23]. Empirical [BDHP20]. Encounter [DKEI20, Kin22, LJB^{+}24]. Endangered [HAP^{+}22, HSS^{+}20, KJB^{+}22, LKP^{+}21, RBFS20, UHB^{+}23, CBKH23]. Endemic [CKB^{+}21, PJD^{+}21]. Endemism [LGQ^{+}21]. Energetic [CFCN20]. Engaging [BDG^{+}23]. Enhancement [HG22]. Enhancing [SDF^{+}21]. Entiat [PC20]. Entosphenus [ACGF23]. entrainment [WPM^{+}23]. Enumerating [BSC^{+}21]. Enumeration [LRM^{+}20]. Environment [BW22, EMRK20, LFS20]. Environmental [BFLI22, CBC^{+}20, CWB^{+}21, HTH^{+}21, HMP22, KSJ^{+}20, MSP21, PSS20, SM22, WHA23, PAE^{+}23, SKBB24]. Environments [AWRL23, GK21, WWK^{+}23]. Equation [BBFQ21, BB21a, MRW^{+}24]. Equations [BBMC23, GBJ21]. Eradicating [MSM21]. Eradication [CBC^{+}20, HBV^{+}21, SLE21, TM21]. Erie [EFK^{+}23, HBV^{+}21, MBF^{+}23, McK23, SM22]. Eriocheir [ZJL^{+}21]. Eroded [HDH^{+}20]. Erratum [Ano21a, Ano22a]. Escape [OTW^{+}22, SHFP20]. Escapement [LCM^{+}23, SKA21, WW21]. Escapements [JWG^{+}21]. Establishment [HBR^{+}23, HSS^{+}20, KSP^{+}24, NHC^{+}20]. Estimate [DEO20, GW21, Har20, HL23, LHR^{+}20, LMO21, RJ22, SKA21, STP^{+}21, CHB^{+}23, SDR24]. Estimates [BOHA20, BCS^{+}21, CEB^{+}23, CDBC22, DAS20, GRM^{+}21, HMQM22, HLM^{+}23, LSSB23, MCB20, ZSLG21]. Estimating [CLM20, CFB^{+}24, DDT^{+}20, GWRF24, HA24, HNM^{+}22, HSTH20, IPZK21, Kin22, LCIL23c, PFGE23, PSS20, PREBK22, SB21a, SCL^{+}20]. Estimation [HJBP20, LDH^{+}22, MdAM21, ORWS21, PS21, KHBF24, PHW^{+}23]. Estimator [DH21]. estuaries [CHO24]. Estuarine [BGC^{+}21]. Estuary [AWO21, WHA23, BHT^{+}21, BSC^{+}21]. Estuary-Level [AWO21]. Ethanol [CHL22]. Ethers [HBTE22]. Evaluate [HJBP20, MDA21, NMBM21, WWBM20]. Evaluated [HMQM22]. Evaluates [KUV^{+}23]. Evaluating [BBB^{+}21, LOC23, LMO21, MBW21, MRM^{+}21, PS21, PTCZ22, RKP^{+}22, SCDS23, TM21, TCB^{+}21]. Evaluation [AP20, ACRB22, BKE22, BBKJ21, BOB^{+}21, BDHP20, CLM20, CEC^{+}24, ESF23, EPB^{+}22, GWMP23, HN21, LMN^{+}20, LFS20, LMA23, MDBC21, MB21, MK21, NLB21, NLW^{+}20, ORWS21, OMAA22, PHW^{+}23, QBA21, RFK^{+}20, RBFQ20, SCN^{+}21, SLE21, AJHB23, FHB^{+}24, RAH^{+}23, SFM23, YBB^{+}23, GW20]. Event [RMWK23]. Events [RSO23]. Evidence [BSKSK24, LEHP21, SCP^{+}23]. Examination [LCIL23a, KMS21]. Example [JOH^{+}22, KS21]. Excluding [CTJS^{+}22]. Exclusion [KR23]. Exhibit [DPR^{+}20]. Expansion [DH21, SMP^{+}21]. Expenditure [PHG^{+}23]. Experiencing [ERT21]. experimental [HCO^{+}24]. expert [CDG^{+}23]. Explain [JOH^{+}22]. Explaining [TPC21a]. Exploitation [BCS^{+}21, DTK20, NRJ^{+}20, SSS^{+}22, WM21, GBM23, YBH24]. exploited [HFJ23]. Exploring [PHG^{+}23, Rac21]. Exposed [BOB^{+}21, MSZ^{+}23]. Exposure [HBTE22, RPJ^{+}21]. Extending [GK21]. External [CCL^{+}22, CFB^{+}24]. Externally [DPR^{+}20]. extraction [BSR^{+}23]. Extreme [BCS^{+}21]. Extrinsic [WD21]. Eye [RGHF23].

Facility [AWZ^{+}20]. Factors [AWF23, BM20, CH22, HAP^{+}22, HMP22, HEP^{+}23, PO22, SMP^{+}21, ZWBS21, dVB20a, BHB^{+}23, BSR^{+}23, HA24]. Failure [LSB^{+}21]. Fall [BGH^{+}20, PFGE23, TRB^{+}21]. Familiarity [GDS^{+}22]. Farms [RPJ^{+}21]. farther [PKSG23]. Fates [RMB^{+}20]. Features [GWMP23]. Fecundity [AFBRS22, LCIL23c, YDM^{+}20]. Federally [DEO20, GRM^{+}21]. Feeding [Owe21, PO22, SSPL20, MMSJ23]. Female [GKRH20, OS20, OTW^{+}22]. Females [HDH^{+}20]. fertile [HFJ23]. fidelity [SWA^{+}24]. Field [YBB^{+}23, AJHB23]. Fields [BD22]. Files [EMRK20]. Fin [IPZK21, LCF20, LLA^{+}22, LMA20, LSSB23, MR21a, MCB20]. Financial [GDS^{+}22]. Finds [QBA21]. Fine [BCW^{+}21, SM22]. Fine-Scale [BCW^{+}21, SM22]. Fingerling [CHL22, WWBM20]. Fingerlings [BTZ^{+}21, TPPC20]. first [MMSJ23]. first-feeding [MMSJ23]. Fish [AAW^{+}23, BBC21, BBB^{+}21, CBC^{+}20, CSQ^{+}21, CLM20, CP21b, CGD^{+}21, DBS21, DH21, ELCM22, EMRK20, GRM^{+}21, GDS^{+}22, HCS^{+}21, HGMR21, JWG^{+}21, KHB22, KNK22, KQG23, KMS21, KJB^{+}22, LL21, LCM^{+}23, LGQ^{+}21, MdAM21, Mic20a, MSL^{+}20, OBR^{+}20, PKTZ20, PBD^{+}22, RZA^{+}21, RFK^{+}20, RLS^{+}23, RMWK23, SPP^{+}21, SVS^{+}21, SBF^{+}23, SPGP21, TM21, TVG^{+}21, TBG^{+}22, TSJ20, WSG^{+}23, WSK^{+}22, YZJ^{+}21, ZK22, AJHB23, BSR^{+}23, EPD^{+}24, HFJ23, HS24, PKSG23, SSR^{+}24, WPM^{+}23, WHA23]. Fish-Tancing [TBG^{+}22]. Fished [WSK^{+}22]. Fisheries [BM20, CM22, CP21b, CDBC22, HBT^{+}21, KBC^{+}20, KQS^{+}20, MRK^{+}22, MKKN22, Mun22, NLK21, NLW^{+}20, PWLH21, PG20, SDF^{+}21, SLZ^{+}20, GBM23, SSM^{+}24]. Fisheries-Independent [CP21b]. Fishers [WM21, CRNRFL24]. Fishery [BCS^{+}21, BHT^{+}20, HYM^{+}22, HDT^{+}22, JSH^{+}21, MDBC21, MAB^{+}20, MSL22, ORWS21, Owe21, PAC22, QBA21, RKP^{+}22, SCW^{+}24, SDWM21, SCL^{+}20, TSBJ20, TJ22, WKM^{+}21, WSK^{+}22, ZK22, vPM20, CR23, FTLR23, GSML23, GBM23, LOC23]. Fishery-Dependent [SDWM21].

Fishery-Independent [WSK^{+}22]. Fishes [BSC^{+}21, BMAM22, HRM^{+}22, HMP22, HGH^{+}21, LRM^{+}20, LHB^{+}23, MR21a, MBCH22, NLB21, RRB^{+}22, YSS22, ZMA20, EPD^{+}24, HGH^{+}23, MSSS23, NMS^{+}23, SDR24, SPM^{+}24]. Fishing [AZL^{+}21, BOB^{+}21, CLRL22, CAZ^{+}21, DDT^{+}20, ERCB22, FM20, HYM^{+}22, MSP^{+}22, Mun22, TBIJ21, TBG^{+}22, YZJ^{+}21, ZLM21]. fishless [KSP^{+}24]. Fishway [RFK^{+}20, SBB23, PTFZ23, SFM23]. fixed [BSR^{+}23]. Flathead [BBKJ21, BOB^{+}21, MS21a, MS22, MK21, NKK21, PTCZ22, SL21, SMP^{+}21, WM21]. Flexible [OTBB23]. Flexible-Pass-Through [OTBB23]. Flood [FPR20]. Flood-Prone [FPR20]. Floodplain [SCMJF20, dVB20b]. Florida [CH22, ERCB22, LMA23, TPPC20]. Flounder [SDWM21]. Flow [BGC^{+}21, DBS21, FDG^{+}21, NRJ^{+}20, PAE^{+}23, ROP^{+}23, SMM^{+}22]. Flows [TPA^{+}21, PAE^{+}23]. Fluvial [HSP21]. Follow [BOB^{+}21]. Follow-Up [BOB^{+}21]. Following [AZL^{+}21, KHD^{+}22, LMA23, SGH^{+}21, WZG22, SMB23]. Food [MR21a, RZK^{+}21, SPY^{+}20, WA23]. Footrope [HLDK21]. Forage [AAW^{+}23, KDO^{+}22]. Foraging [SG22]. Fore [TCB^{+}21]. Forebay [AWZ^{+}20]. Forensic [WMJ^{+}24]. Forest [GWMP23]. Fork [RSK^{+}21]. Form [BDC22]. formalin [BSR^{+}23]. formalin-fixed [BSR^{+}23]. Formats [LTB^{+}22]. Fort [SNLD21]. Forward [MNA^{+}20]. Fouling [CJP^{+}21]. Four [HJB^{+}21, LHR^{+}20, RZK^{+}21, SVS^{+}21, KHBF24]. Fragmented [Sin21, LOV^{+}23]. Frame [KBDI21]. Francisco [BSC^{+}21]. Frequency [CMBR20, EMRK20, MS22, MSZ^{+}23, MSP21, PTCZ22, SB21b]. Frequent [CDB22]. Freshwater [HIF21, KBDA21, MR21a, QBA21, BSR^{+}23]. Front [LCIL23b]. Frozen [BBC21]. fuel [CR23]. Future [GDS^{+}22, KS21, LCM^{+}23, PKNS21, TPC^{+}22, PBEK23]. Fyke [TCB^{+}21]. Fyke-Net [TCB^{+}21].

Gafftopsail [CHO24]. Gag [CLRL22]. gains [GBM23]. Gar [AKC20, AP20, BS20, DAS20, HDH^{+}20, KS20, KDB^{+}20, LBH21, LSP20, SCMJF20, SDB^{+}20, SBD^{+}20, SSPL20, SLP20, SSSP22, TLSP20]. Gastrically [DPR^{+}20]. Gear [GET22, MDA21, PSS20, WSK^{+}22]. Gears [RPEF^{+}23]. Generating [CVD^{+}20]. generosa [HA24]. Genetic [AWG^{+}20, Ano22a, BAH^{+}20, BVP^{+}21, BHCC23, CHB^{+}23, CKB^{+}21, DC22, DRPW21, EFK^{+}23, HHH^{+}22, HHH^{+}23, JSH^{+}21, KLR^{+}21, LOV^{+}23, LHR^{+}20, LEHP21, MWASW20, MKK^{+}21, WPGC21, Ano23g]. Genetics [BSC^{+}21, BDHP20, TLSP20]. genomes [BSR^{+}23]. Genomic [AWO21]. genomics [LHPS^{+}24]. Genotyping [MBP^{+}22]. Genotyping-in-Thousands-by-Sequencing [MBP^{+}22]. geochemistry [WMJ^{+}24]. geoduck [HA24]. Geographical [ZJL^{+}21]. Geometric [ZJL^{+}21]. Geomorphic [HSP21]. Georgia [TBIJ21, YBH24]. Ghost [MSP^{+}22]. Gila [MSD^{+}22]. Gill [BWM22, HLDK21, Kin22, RAB^{+}20, SWJ^{+}21, TCL^{+}21]. Gill-Net [BWM22, HLDK21, Kin22]. Gillnetting [WGKB20]. Gizzard [CHL22, KNK22, STP^{+}21]. Glacial [BSKM20, DNW^{+}22, MW20, EPD^{+}24].

Gloves [LL21]. Gold [RKP^{+}22]. goldeye [MRW^{+}24]. good [EPD^{+}24]. Gowan [Ano22a]. Grade [BBO20, LRM^{+}20]. Gradients [NRJ^{+}20]. Grand [HSS^{+}20]. Grande [COT20]. Grandparent [DC22]. Grass [CBS21, FNB^{+}22, HBV^{+}21, HBR^{+}23, HCP^{+}23, YBB^{+}23, ZZWX20]. Grayling [GBJ21, MRM^{+}21, RM21, SYHF24]. Great [CFB^{+}24, FNB^{+}22, RZK^{+}21, YDM^{+}20, BKE22, NMS^{+}23, SDF^{+}21]. Greater [DAS20, GRM^{+}21]. Green [BQR^{+}20, ZK22, IDH^{+}23, SHI22]. Grounds [DPR^{+}20, WB24]. Groundwater [WB21]. Groundwater-Influenced [WB21]. Groupers [RMB^{+}20]. Grow [NLK21]. Growth [AOB^{+}22, BKM21, CH22, CLSM20, CBS21, CDB22, EP21, ESCI21, HMQM22, HGSS20, HJB^{+}21, JJD^{+}23, LMO21, NLK21, NHC^{+}20, RCD20, SHI22, SR20, SPY^{+}20, WD21, ZMA20, EPD^{+}24, MMJW23, TO23]. guarding [PK24]. Guatemala [QBA21]. Guide [STB^{+}21]. Guided [CVD^{+}20]. Gulf [AS20, AWG^{+}20, AOB^{+}22, CTJS^{+}22, SCL^{+}20, WSK^{+}22, CLRL22, GRM^{+}21, NLW^{+}20, PAC22, PG20].

Habitat [AKC20, BCW^{+}21, BBKJ21, BHT^{+}21, COT20, CTJS^{+}22, CP21b, GKRH21, GWM21, HQWD20, HDQ22, HCP^{+}23, JDO^{+}20, KR23, MDHQ22, MVM^{+}23, MSP21, MSV^{+}21, NLB21, PC20, PBD^{+}22, RH20, RWS^{+}21, RMR^{+}22, SSF^{+}23, SKWR23, SSH22, TVG^{+}21, WW20, WSK^{+}22, YZJ^{+}21, ZK22, CFGZ23, McK23, ROP^{+}23, SSR^{+}24]. Habitat-Based [WSK^{+}22]. Habitats [HBR^{+}23, KHD^{+}22, RGMH20, SCMJF20, SMM^{+}22, HLSW23]. Habituate [DS20]. hake [LHPS^{+}24]. Half [OTBB23]. Half-Duplex [OTBB23]. Halibut [LDH^{+}22]. Hand [BOB^{+}21]. Handling [CJN^{+}21, GRA^{+}22, LL21]. Hard [MGNP22]. Hard- [MGNP22]. Harris [WCSL22]. Harry [HBR^{+}23]. Harvest [ARK^{+}21, CFPSW21, CM22, HDT^{+}22, KBC^{+}20, MKKN22, Mun22, NMBM21, ORWS21, RJ22, SHWW22, SSSP22, WW21]. hatch [WMB23]. Hatchability [MKG^{+}22]. Hatchery [BHT^{+}21, BHCC23, COT20, CEH^{+}20, DC22, ELCM22, FDG^{+}21, HDP^{+}22, HHH^{+}22, HHH^{+}23, JWG^{+}21, KFG22, LMN^{+}20, LMA20, MNA^{+}20, MWMH22, PE21, RAHS20, SCN^{+}21, WRH^{+}21, Ano23g, HGH^{+}23, KHL^{+}24, PO24, WB24, ZAK^{+}23]. Hatchery- [WRH^{+}21, PO24, ZAK^{+}23]. Hatchery-Origin [RAHS20, SCN^{+}21]. Hatchery-Produced [MWMH22]. Hatchery-Reared [FDG^{+}21, LMN^{+}20, LMA20, PE21, HGH^{+}23]. Hatching [ESCI21, GKRH20, LSP20]. Haul [LGOM22]. Head [LBH21]. Headwater [Bar21, CDB22, HSW21, HCS^{+}21, RRL^{+}21]. Health [KHB22]. Heavily [CBLW21]. heavy [GBM23]. Help [JOH^{+}22, TSJ20]. Henrys [MVH^{+}22]. Herding [RLS^{+}23, RAH^{+}23]. herring [HCO^{+}24, SLD^{+}23]. Hidden [RMB^{+}20]. Hierarchical [PS21]. High [EMRK20, KNC^{+}20, MSSS23]. High-Density [EMRK20]. Highland [CLTJ21, RWS^{+}21]. Highly [MDA21]. Historical [LCM^{+}23, MSLP22, NMW^{+}21, WAKO21]. historically [LOV^{+}23]. History [AWO21, BBP20, CLSM20, DNW^{+}22, DBCS20, DRPW21, HNM^{+}22, KFG22, SG21]. Holding [SFFW21]. home [RHFU24].

Homeward [WRH^{+}21]. Hooked [WSK^{+}22]. Hooking [CJN^{+}21, GC21]. Hoop [TPC21b]. Hormones [HBTE22]. Hudson [PA21]. Hull [BD22]. Hulls [CJP^{+}21]. Human [KBDA21]. Humpback [HSS^{+}20, SPY^{+}20]. Huron [PO22]. Hybridization [LEHP21, PJD^{+}21, TLSP20]. Hybrids [HBT^{+}21, LSB^{+}21, NLK21]. Hydraulic [WCSL22, SFM23]. Hydro [SMM^{+}22]. Hydro-Regulated [SMM^{+}22]. Hydroacoustic [MSV^{+}21]. Hydrodynamic [SKWR23]. Hydroelectric [CVD^{+}20]. Hydrography [DGJ^{+}22]. Hydrokinetic [CFHS22]. Hydrologic [ESCI21, SBD^{+}20]. Hydrologically [DBS21]. Hydrology [EP21]. Hydropeaking [SKGL22]. Hydropower [AWZ^{+}20, BPL^{+}21, WSG^{+}23]. Hyperstability [SWM22]. hyperstable [GBM23]. hypotheses [PKSG23].

Ice [GGW21, KSJ^{+}20, SGH^{+}21]. Ichthyoplankton [KBDI21]. Ictalurid [Rac21]. Idaho [HDQ22, HDT^{+}22, GGWV22, GKT^{+}22, MVH^{+}22, RSK^{+}21, RMR^{+}22, WD21, WMJ^{+}24]. Identification [BS20, DRPW21, EMRK20, JSH^{+}21, LMA20]. identifies [WMJ^{+}24]. Identify [BB21b, DC22, GWMP23, LSB^{+}21, MWR21, RR23]. Identifying [KSJ^{+}20, STB^{+}21, SM22, SLZ^{+}20, TVG^{+}21]. Identity [GDS^{+}22]. illegal [WMJ^{+}24]. Illinois [AAW^{+}23, PBK^{+}23, SMCEF^{+}23, ZLM21]. Image [HJ21]. Image-Based [HJ21]. Imagery [AFBRS22, LRM^{+}20]. Imaging [CHJ^{+}22, HL23, NKG23]. Immersion [LL21, SLP20]. Impact [TSBJ20]. Impacts [CFPSW21]. Impair [HSF^{+}22]. Impairment [CRH^{+}22]. Impart [ZSLG21]. Impassable [KES^{+}21]. Impedance [CFCN20]. impeller [WPM^{+}23]. Imperfect [KPD^{+}24, HFJ23]. Imperiled [BBB^{+}21, OBR^{+}20, HGH^{+}23]. Implant [BR21, LFS20, MB21]. Implantation [CCL^{+}22, LL21]. Implementation [KR23, SSF^{+}23]. Implemented [ORWS21]. Implications [AMR^{+}23, BQR^{+}20, CTJS^{+}22, CKB^{+}21, ESCI21, HBV^{+}21, HBR^{+}23, HLM^{+}23, Kep21, KBDA21, LCF20, MRM^{+}21, MKK^{+}21, NRJ^{+}20, RG20, RLS^{+}23, RAHS20, RCR21, SPM^{+}24, SMM^{+}22, CHO24, JJD^{+}23, LHB^{+}23, MSSS23, PAE^{+}23, RAH^{+}23]. Importance [GvRSS21, GWM21, MDD^{+}23, MSV^{+}21]. Important [NMW^{+}21, Gep23]. Impounded [HRM^{+}22]. Impoundment [KNK22, Mic20a]. Impoundments [MKKN22, Mic20b, PSS21, CEC^{+}24]. Imprinting [CLBL22]. Improve [WGBK22, CRNRFL24, MMJW23, SBW^{+}23]. Improved [EMRK20, MdAM21]. Improvement [KQG23]. improves [ROP^{+}23]. Improving [CDBC22, GSML23, KHJS22, MK21, PBEK23]. In-River [WRH^{+}21]. In-Stream [CEH^{+}20, PC20]. Incidental [MDBC21]. Including [BMAM22, HBTE22]. Incorporating [DWGR21, ECC22, RAB^{+}20]. Increase [PJD^{+}21, WGKB20]. Increased [ELCM22, RHUM^{+}23]. Increases [FDG^{+}21, PC20]. Increasing [PST^{+}20, RRB20, RSM^{+}23]. Incubate [SCN^{+}21]. Incubation [SMM^{+}22]. Incubators [CEH^{+}20, MRM^{+}21, SCN^{+}21]. Independent [CP21b, MCB20, WSK^{+}22]. Index [CSLW22]. Indexing [BGH^{+}20, TPC21b]. Indicates [LCIL23a, EFK^{+}23]. Indicators [KHB22]. Indices [BWM22, DNW^{+}22, PHG^{+}23, SDWM21]. Indirect [RAB^{+}20]. industry [SPM^{+}24]. inexpensive [RH23]. Infection [RAHS20, SMB23]. Infer [Rac21]. Inference [DC22]. inferences [KPD^{+}24]. Inferred [HRS^{+}20, MBCH22, SSPL20]. Influence [BKM21, DBS21, FMM20, GGW21, KQS^{+}20, LLE22, MR21b, ZLM21, BHB^{+}23, CR23, LHB^{+}24]. Influenced [Mun22, WB21]. Influences [BFM^{+}22, HMP22, RGMH20]. Influencing [BM20, dVB20a, BSR^{+}23]. Inform [FNB^{+}22]. Information [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, KLR^{+}21, SWJ^{+}21]. Informative [SLZ^{+}20]. Informing [MVH^{+}22, KPD^{+}24]. Informs [CBC^{+}20, KFG22]. Infrastructure [AWZ^{+}20]. Ingestion [PE21]. Initially [DS20]. Initiative [LHWS23]. injection [HCO^{+}24]. Injuries [TCL^{+}21]. Injury [CMBR20]. Inland [RSO23, TBG^{+}22]. inlet [KHL^{+}24]. Inshore [TJ22]. Insight [KFG22, WTW^{+}23]. Insights [NRJ^{+}20, SDF^{+}21]. Instream [NRJ^{+}20]. Intake [AWZ^{+}20]. Integrated [CJP^{+}21, CGD^{+}21, DWSM21, DWGR21, FHA^{+}21, GK21, HIF21, HMB21, KFG22, MVH^{+}22, OTBB23, SG21, SBB23, TRB^{+}21]. Integrity [MWASW20]. Interactions [CLSM20, CHJ^{+}22, CFHS22, HMO^{+}21, WB24]. Interbasin [MW20]. intercept [CRNRFL24]. Interconnected [MW20, SGM^{+}20]. Intermittent [HUTB21, RZA^{+}21]. internal [CFB^{+}24, DSW^{+}24]. Interpopulation [CFCN20]. Interpretations [ORL^{+}22]. Interspecific [EPD^{+}24]. Interventions [ERT21]. Intra [ASDW23]. Intra-Annual [ASDW23]. intraperitoneal [HCO^{+}24]. Introduced [KMS21, LEHP21, RWS^{+}21, RBFQ20, RCR21, SDL^{+}23, SLE21]. Introduction [CP21a, GK21, HMB21, SK23, SDB^{+}20, VF21, WMJ^{+}24]. Invasion [LCIL23b, PJD^{+}21]. Invasive [CSWP23a, CSWP23b, FNT21, FNB^{+}22, GET22, GvRSS21, GWMP23, HMP22, HJB^{+}21, HIF21, HMO^{+}21, KQG23, NF21, PBK^{+}23, PJD^{+}21, PKTZ20, RPEF^{+}23, SGK^{+}22, SMP^{+}21, SWJ^{+}21, WGBK22, AMR^{+}23, PK24, SPM^{+}24, SCP^{+}23]. Investigation [PFGE23, UHB^{+}23]. Investments [BDG^{+}23]. Invisible [HMP22]. Iowa [KS21, MK21, WWBM20, WW20]. Irrigation [MLB21]. Isolated [SNLD21]. Isotope [FMM20, RGHF23]. Isotopes [MR21a, SSPL20, WZG22]. Issue [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b].

James [SL21, MWMH22]. Japanese [PK24, SLL^{+}21]. japonica [SLL^{+}21]. Jinsha [LGQ^{+}21, YZJ^{+}21]. Jonah [OS20]. Juvenile [AAW^{+}23, BS20, BPT21, BDC22, DMBC23, EPB^{+}22, FMM20, HJ21, HEP^{+}23, HBTE22, IPZK21, KSJ^{+}20, MGNP22, PFGE23, PA21, PE21, PO22, RPJ^{+}21, RCD22, SWTH21, WSG^{+}23, CBKH23, GWRF24, SKBB24, WA23].

Juveniles [CEH^{+}20].

Kansas [DHWP21, MKKN22, NLK21, SWJ^{+}21]. Keep [HMP22]. Keys [DMBC23]. Klamath [CBKH23, KJB^{+}22, KJB^{+}22, TPA^{+}21]. Klickitat [CHB^{+}23, ZAK^{+}23]. Knowledge [SDB^{+}20, CDG^{+}23]. Known [HGSS20, ORL^{+}22, SHI22]. Known-Age [HGSS20, ORL^{+}22, SHI22]. Kokanee [AWZ^{+}20, BWM22, HMO^{+}21, KQS^{+}20, MVM^{+}23]. Kvichak [CFHS22].

Laboratory [AJHB23, LSP20, WPM^{+}23, WA23, MMJW23]. Labrador [GBM23]. Lacustrine [GKRH21]. LaGrange [AAW^{+}23]. Lake [AZL^{+}21, BAH^{+}20, BSKM20, BB21a, BHT^{+}20, CBB^{+}23, HNM^{+}22, HDQ22, HDT^{+}22, HMQM22, HBTE22, HMO^{+}21, IPZK21, IDH^{+}23, KDM^{+}23, LMN^{+}20, LLE22, LMA20, MKG^{+}22, MDD^{+}23, MVH^{+}22, MWASW20, MNA^{+}20, MW20, MBCH22, MPM21, ORL^{+}22, PKTZ20, PO22, RH20, SSS^{+}22, SLG^{+}23, SHI22, SGK^{+}22, SSH22, SLE21, SGM^{+}20, TCM20, WGKB20, WGBK22, ZK22, CEB^{+}23, CR23, KSP^{+}24, PK24, CBKH23, CEB^{+}23, CLBL22, EFK^{+}23, FFG20, FHR21, HBV^{+}21, HMQM22, KDM^{+}23, KJB^{+}22, LHR^{+}20, MKG^{+}22, MBF^{+}23, McK23, PKTZ20, SM22, TLSP20, WMJ^{+}24, WGBK22]. Lakes [ASDW23, BGH^{+}20, BFLI22, BFM^{+}22, CH22, CSQ^{+}21, DKEI20, DNW^{+}22, GHA^{+}23, GBJ^{+}22, GKT^{+}22, HGSS20, LHR^{+}20, LLE22, SSF^{+}23, SVS^{+}21, SPGP21, TPPC20, TBG^{+}22, WD21, WWBM20, ZMA20, BHB^{+}23, EPD^{+}24, CFB^{+}24, FNB^{+}22, PO22, RZK^{+}21, YDM^{+}20]. Lampetra [ACGF23]. Lamprey [JDO^{+}20, LHWS23, MLB21, ACGF23, LHB^{+}23, LCVS23, PTFZ23, CFGZ23, HLJ^{+}23, MMJW23, PHW^{+}23, WTW^{+}23]. Lampreys [ACGF23, CHS^{+}22, KNC^{+}20, LHB^{+}23, RG20, BHS^{+}23, HLSW23, LHWS23, Gep23, MMSJ23, SCDS23]. Lands [RSO23]. Landscape [HW22, MML^{+}22]. Landscape-Scale [HW22]. Landscapes [DBS21]. Lapse [ECC22]. Large [BPL^{+}21, BGC^{+}21, GGWV22, HRM^{+}22, HDH^{+}20, HCP^{+}23, MGMP21, MQK^{+}20, MK21, OTBB23, PFGE23, PTCZ22, RWS^{+}21, RPEF^{+}23, SLL^{+}21, SBF^{+}23, WWBM20, WDPH22, dVB20b, HS24]. Large-River [RPEF^{+}23, HS24]. Large-Scale [BGC^{+}21]. Largemouth [AZL^{+}21, BTZ^{+}21, BDHP20, CAZ^{+}21, HDP^{+}22, LMA23, MSW22, SW22, TPPC20, WD21, CEC^{+}24, LHB^{+}24]. larger [PKSG23]. Largest [HYM^{+}22]. Larvae [JDO^{+}20, dVB20b, RHUM^{+}23]. Larval [CHS^{+}22, CHL22, HRS^{+}20, KBDI21, MGMP21, MLB21, RPEF^{+}23, BHS^{+}23, HLSW23, LHB^{+}23, MMJW23, MMSJ23]. Late [AZL^{+}21]. Late-Season [AZL^{+}21]. Latitude [MPM21]. Laurentian [CFB^{+}24, FNB^{+}22, RZK^{+}21]. Leading [AWG^{+}20]. Leads [DGJ^{+}22]. Leap [SHFP20]. Learned [CBC^{+}20, CRNRFL24]. Leech [LHR^{+}20]. Legacy [Rac21]. Lemon [HCC^{+}21]. Length [BBC21, BBFQ21, BKE22, BB21a, BBMC23, DGJ^{+}22, DMBC23, HBBS22, KQS^{+}20, MSW22, Mic20b, MDA21, SL21, STP^{+}21, MRW^{+}24].

Length-Based [HBBS22, SL21]. Lens [RGHF23]. Lentic [RGMH20]. Less [GC21]. Lessons [CBC^{+}20, CRNRFL24]. Level [AWO21, SW22, TBIJ21, SSM^{+}24, SMB23]. Levels [ARK^{+}21, MBW21, ZSLG21]. Leverage [SCMJF20]. License [TBG^{+}22, ZLM21]. Licensing [RJ22]. Life [AWO21, AWRL23, BBP20, BBB^{+}21, CLSM20, DNW^{+}22, DBCS20, DRPW21, HNM^{+}22, KUV^{+}23, KFG22, SG21, WWK^{+}23, AJHB23, PHW^{+}23]. Life-History [DNW^{+}22]. Lifespan [LKP^{+}21]. Light [KBDI21, LTL^{+}23, dVB20a, dVB20b, AJHB23]. light-emitting-diode-lit [AJHB23]. Limestone [HKPP20]. Limitations [MSL^{+}20]. Limited [BW22, DNW^{+}22, ORWS21, SLZ^{+}20]. Limits [BKE22, CM22, KBC^{+}20, MSW22, MDA21]. Line [GC21]. lit [AJHB23]. Little [SPY^{+}20]. littoral [SSR^{+}24]. Live [CTJS^{+}22, MML^{+}22, SPM^{+}24, LHB^{+}24, NKG23]. Live-bait [SPM^{+}24]. Live-Bottom [CTJS^{+}22]. live-imaging [NKG23]. live-well [LHB^{+}24]. Livestock [KR23]. Lobster [SCL^{+}20]. Location [HBTE22, MR21b, SGM^{+}20, ZSLG21]. Location-Dependent [ZSLG21]. Locations [IDH^{+}23, KSJ^{+}20, SFFW21, MBF^{+}23]. Lock [DS20, LBH21, LTL^{+}23]. Logging [LDH^{+}22]. Long [CBS21, LMA23, Mun22, PA21, SGH^{+}21, GBM23, HA24]. Long-Distance [SGH^{+}21]. Long-Term [CBS21, LMA23, Mun22, PA21, GBM23, HA24]. Longline [LDH^{+}22]. Longlines [GC21]. Loons [MRK^{+}22]. Lost [EPB^{+}22, KJB^{+}22]. Lotic [SR20]. Louisiana [SDWM21]. Low [AWZ^{+}20, HCS^{+}21, LBH21, MS22, MSZ^{+}23, MSP21, PTCZ22, SB21b, SMB23, KHL^{+}24]. Low-Conductivity [HCS^{+}21]. Low-Frequency [MS22, MSZ^{+}23, MSP21, PTCZ22, SB21b]. Low-level [SMB23]. Lower [DPR^{+}20, AKC20, AP20, HCM^{+}22, SL21]. Lowland [HCS^{+}21].

M [ACRB22]. macabre [PBEK23]. Macrhybopsis [PBEK23]. Macrohabitat [LBH21]. Macroscopic [KHB22]. Mad [KFG22]. Made [MCB20]. Madtom [CKB^{+}21, JCB^{+}21, UHB^{+}23]. magellanicus [LOC23]. Magnitude [PBK^{+}23]. Main [CSQ^{+}21, PE21, BHS^{+}23]. Main-Stem [PE21, BHS^{+}23]. Maine [PTFZ23, KKMA23, SCL^{+}20]. Male [MSM21, OS20, OTW^{+}22, PK24, BSKSK24, SWA^{+}24]. Manage [BSKM20, MSL^{+}20]. Managed [GRM^{+}21]. Management [BGC^{+}21, BQR^{+}20, BDG^{+}23, CBS21, CDBC22, CJP^{+}21, CGD^{+}21, ESCI21, FNT21, FHA^{+}21, GK21, HDT^{+}22, HIF21, HMB21, KLR^{+}21, Kep21, KFG22, KBDA21, MDBC21, MBW21, MVH^{+}22, MKK^{+}21, Mun22, NRJ^{+}20, PAC22, RLS^{+}23, RAHS20, SNLD21, STB^{+}21, SDF^{+}21, SMM^{+}22, SDB^{+}20, SLZ^{+}20, TSBJ20, TJ22, TPA^{+}21, YSS22, AMR^{+}23, GBM23, LHB^{+}23, MSSS23, PBEK23, SCW^{+}24, SBW^{+}23]. Managers [BPT21]. Managing [DBCS20, FHK23, HMP22, PWLH21, PAE^{+}23]. Manitoba [MNA^{+}20]. Maps [BBO20]. Marine [CLSM20, KBC^{+}20, LFS20, KHL^{+}24, WTW^{+}23]. Mark [DWGR21, HKPP20, SHI22, SWA^{+}24, ZSLG21, FTLR23, HS24]. mark-recapture [HS24]. mark-selective [FTLR23]. Markers [LSB^{+}21]. Marking [LCF20, LMA20, NKK21, SLP20]. Markov [RMB^{+}20]. Maryland [HGM22, MKK^{+}21, WCSL22]. mask [GBM23]. Mass [RLS^{+}23]. Massachusetts [TCB^{+}21]. Maternal [HDH^{+}20]. Maternally [FMM20]. Maternally-Transferred [FMM20]. mature [MMSJ23]. Maturity [OS20]. Maxillary [HMQM22]. Maximum [BMAM22, Pan21, RHUM^{+}23]. May [HDH^{+}20, PA21, SPGP21]. Medal [RKP^{+}22]. Medium [OTBB23]. Menhaden [NLW^{+}20]. Mercury [ARK^{+}21]. Metabolic [GRA^{+}22]. Metabolism [BLS23]. Method [AFBRS22, EMRK20, LMN^{+}20, LCIL23c, MdAM21, SMM20, TPC^{+}22, RH23]. Methods [CAZ^{+}21, CHS^{+}22, FM20, FNB^{+}22, HRS^{+}20, KS20, LMO21, MB21, NKK21, PSS21, PREBK22, SB21a, SCMJF20, TPC^{+}22, DSW^{+}24, HLSW23]. Metrics [HNM^{+}22, HBBS22]. Mexico [AS20, AWG^{+}20, AOB^{+}22, CLRL22, CTJS^{+}22, GRM^{+}21, HTH^{+}21, PAC22, PG20, WSK^{+}22]. Michigan [PO22, BHB^{+}23, CEB^{+}23, IDH^{+}23, MKG^{+}22, MDD^{+}23, MRM^{+}21, SHI22, ZK22]. Microchemical [HKPP20]. Microchemistry [SGM^{+}20, WWK^{+}23, KKMA23]. Microfishing [BDC22]. Micropterus [LEHP21, PJD^{+}21]. Microsatellite [FLW21]. Microstructure [WWK^{+}23]. Mid [AZL^{+}21, Kep21]. Mid- [AZL^{+}21]. Mid-Atlantic [Kep21]. Middle [CSQ^{+}21, OS20]. Midwest [SGM^{+}20]. Midwestern [KDO^{+}22, MDA21, PKSG23]. Migrating [CVD^{+}20, CHJ^{+}22, LGOM22, HAP^{+}22]. Migration [AS20, BPL^{+}21, Boo20, BW22, KNC^{+}20, MWR21, MSLP22, MST22, SG21, SM22, CHB^{+}23]. Migratory [CLM20]. Milford [PTFZ23]. minimal [HLSW23]. Minimize [GET22]. Minimum [BKE22, MDA21]. Mining [RKP^{+}22]. Minnesota [GHA^{+}23, HBBS22, KMS21, Sin21]. Minnow [COT20, MB21, OBR^{+}20]. Miramichi [ARK^{+}21, HL23]. misidentification [KPD^{+}24]. Mississippi [AKC20, AWRL23, CSWP23a, CSWP23b, CP21a, LTL^{+}23, WWK^{+}23, WM21, DTK20]. Missouri [FFG20, SNLD21, BAH^{+}20, HSP21, HBR^{+}23, Mic20b, WA23, WM21]. Mitochondrial [ACGF23]. Mitten [ZJL^{+}21]. Mixed [DH21, FLW21, HYM^{+}22, JSH^{+}21, MBP^{+}22]. Mixed-Stock [DH21, FLW21, HYM^{+}22, JSH^{+}21, MBP^{+}22]. Mixing [IDH^{+}23]. mm [TRB^{+}21]. Moapa [RBFS20]. Mobile [AP20, HLM^{+}23, MDA21, ZWBS21, LSB^{+}21]. Model [BB21b, BDG^{+}23, CFPSW21, DWGR21, GWMP23, KUV^{+}23, MVH^{+}22, OMAA22, RAB^{+}20, SKA21, SCW^{+}24, SFM23]. Modeled [KDO^{+}22, SHWW22]. Modeling [BBB^{+}21, CFCN20, MGMP21, PS21, PJD^{+}21, SSSP22, TCM20, HS24, SBW^{+}23]. Models [OBR^{+}20, RMB^{+}20, SCMJF20, SHI22]. Modern [TPC^{+}22]. Modification [HLDK21, KHD^{+}22]. Modified [TSJ20]. Molecular [FJH^{+}22, MSD^{+}22]. Monitor [HHH^{+}23]. Monitoring [Ano23g, BVP^{+}21, HMP22, HHH^{+}22, KHJS22, KR23, RPEF^{+}23, SG21, CHB^{+}23, WHA23]. Monkfish [WKM^{+}21].

Monongahela [CFPSW21]. Montana [LKD23, SNLD21, VF21]. Montane [SLE21]. Moon [BWM22]. Morphometric [BB21b, GWMP23, OS20, ZJL^{+}21]. Morphometrics [HJ21, SYHF24]. Morphometry [Kin22]. Mortality [AOB^{+}22, BTZ^{+}21, CJN^{+}21, CBS21, GGW21, GC21, LDH^{+}22, MSW22, MdAM21, NKC^{+}23, PFGE23, SSSP22, SCL^{+}20, SW22, TSJ20, WW21, WKM^{+}21, CEB^{+}23, DSW^{+}24, FTLR23, HLSW23, LJB^{+}24, MLB^{+}23, MSSS23]. mosaic [LOV^{+}23]. Motor [DS20]. Mountain [RMR^{+}22, SG21]. Movement [AWRL23, BCW^{+}21, BQR^{+}20, EMRK20, HBV^{+}21, HGH^{+}23, HDQ22, HGM22, Kep21, MQK^{+}20, MW20, MPM21, WZG22, WB21, SPM^{+}24, SSR^{+}24]. Movements [DTK20, IDH^{+}23, LBH21, SGH^{+}21, WRH^{+}21]. Moving [HMP22]. Mud [GET22]. Mudflats [NHC^{+}20]. Multi [EPB^{+}22]. Multi-Predator [EPB^{+}22]. multidecadal [CRNRFL24]. Multidimensional [SBW^{+}23]. Multigeneration [HHH^{+}22, HHH^{+}23, Ano23g]. Multipass [BOHA20]. Multiple [HQWD20, LTL^{+}23, SKA21, WWK^{+}23, CFB^{+}24, PKSG23]. Multiscale [TO23]. Multispecies [RFK^{+}20]. Multistate [OMAA22]. Muskellunge [BCS^{+}21, BLS23, DKEI20, GHA^{+}23, HGM22, HGSS20, KMS21, KSPP22, KBDI21, LHR^{+}20, MW20, MCB20, PWLH21, SHI22, WW20]. Mussels [CGD^{+}21, WZG22]. mykiss [KHD^{+}22]. Mysids [HMO^{+}21]. Myxospore [ESF23].

N [Ano22a]. Natal [RGHF23, KKMA23]. Native [HRM^{+}22, HGH^{+}21, HGMR21, LEHP21, LHB^{+}23, MDHQ22, MBCH22, NLB21, OBR^{+}20, SDL^{+}23, TVG^{+}21, LHWS23, SPM^{+}24, SCP^{+}23]. Natural [BFM^{+}22, CEH^{+}20, DH21, DRPW21, HSH21, LMA20, RG20, WW21, WRH^{+}21, CEB^{+}23, PO24, ZAK^{+}23]. Natural-Origin [DH21, WRH^{+}21, PO24, ZAK^{+}23]. Naturally [BHT^{+}21]. Nature [HYM^{+}22]. Near [DPR^{+}20]. Nearshore [BCW^{+}21, BHT^{+}21]. Nebraska [BBKJ21, HSP21, RMWK23]. Necessarily [SPGP21]. Negotiations [TBIJ21]. nest [SWA^{+}24]. Net [BWM22, HLDK21, Kin22, RAB^{+}20, RPJ^{+}21, SWJ^{+}21, TCB^{+}21, CBKH23]. net-pens [CBKH23]. Nets [KBDI21, TPC21b]. Network [JDO^{+}20]. Newly [BTZ^{+}21]. nexus [SBW^{+}23]. Niche [GHA^{+}23, SBW^{+}23]. niches [SBW^{+}23]. Nighttime [HCM^{+}22]. Nitrogen [FMM20]. No [SCP^{+}23]. Nocturnal [KHJS22]. Nongame [YSS22, MSSS23]. Nonlethal [Ano22a, MR21a, WPGC21, SYHF24]. Nonlinear [Pan21]. Nonnative [BQR^{+}20, HGH^{+}21, MDHQ22, MBCH22, OBR^{+}20, TM21, SPM^{+}24]. Nontarget [TSBJ20]. Nontidal [HGM22]. nonwadeable [SDR24]. North [MWR21, MSD^{+}22, SSS^{+}22, BIE^{+}22, HRK^{+}22, KLR^{+}21, LCM^{+}23, MR21a, RH20, RR23, RRB^{+}22, SWJ^{+}21, SPSH20, SGM^{+}20, UHB^{+}23, WAKO21]. Northeast [ERCB22]. Northern [GBJ^{+}22, JCB^{+}21, LCF20, MKG^{+}22, PCM^{+}21, SGH^{+}21, SM22, TCL^{+}21, TBG^{+}22, UHB^{+}23, ZMA20, AS20, BGH^{+}20, BFM^{+}22, MLB^{+}23, MDD^{+}23, SVS^{+}21, WD21, ZK22]. Northwest [RSO23]. Northwestern [AWG^{+}20, HGSS20]. Norway [JOH^{+}22]. Novel [DHWP21, HBR^{+}23, SMM20, GWRF24]. Nucleotide [BVP^{+}21, MBP^{+}22]. Number [HL23]. Numbers [HJBP20, HRS^{+}20]. Nurse [AS20]. Nurseries [SGM^{+}20]. Nursery [GKT^{+}22]. Nutrient [HRM^{+}22].

Oahe [FFG20]. Oak [BBKJ21]. Objectives [FNT21, MBW21]. observation [HA24]. Observations [DEO20]. Observes [CHJ^{+}22]. occupancy [CDG^{+}23]. Occurrence [CH22, HDQ22, SCMJF20]. Occurrences [HTH^{+}21]. Occurring [GHA^{+}23]. ocean [KHL^{+}24, WTW^{+}23]. Oceanic [TSBJ20]. Odell [MBCH22]. off [HCC^{+}21, KKMA23, ROP^{+}23, RRB^{+}22]. off-channel [ROP^{+}23]. Offs [HG22]. Offshore [ERCB22, GRM^{+}21, PBD^{+}22]. Offspring [DC22, HLJ^{+}23]. Ohio [OBR^{+}20, HBV^{+}21, PWLH21, TPC21a]. Oklahoma [CLTJ21, LSSB23, YSS22]. Old [SK23]. Oligotrophic [HRM^{+}22]. Olympic [MSLP22]. Oncorhynchus [KHD^{+}22]. one [LKD23]. one-time [LKD23]. Online [PWLH21]. Ontario [AZL^{+}21, CLBL22, FB20, FHR21]. Ontogenetic [FHR21]. Open [RPJ^{+}21]. Open-Net-Pen [RPJ^{+}21]. Operation [RZA^{+}21, KPD^{+}24]. Opercula [LSSB23]. opilio [GLSF21, GBM23]. opposing [NMS^{+}23]. Optimal [MS22, SBH^{+}23]. Options [KUV^{+}23]. Oregon [CFGZ23, BHS^{+}23, HSW21, JDO^{+}20, MBCH22, WRH^{+}21]. Organic [KDM^{+}23]. Organisms [CJP^{+}21]. orientations [RR23]. Origin [CLBL22, DH21, RAHS20, RGHF23, SCN^{+}21, WRH^{+}21, ZJL^{+}21, KKMA23, PO24, ZAK^{+}23]. Other [BSC^{+}21, GRM^{+}21]. Otolith [AWRL23, DAS20, DRPW21, HNM^{+}22, ORL^{+}22, SB21a, SR20, SMM20, SGM^{+}20, WWK^{+}23, KKMA23]. Otolith-Based [DAS20]. Otolith-Derived [ORL^{+}22]. Otoliths [BBP20, BS20, CLM20, LKP^{+}21, LSSB23, SLP20]. Out-Migrating [CVD^{+}20, LGOM22, HAP^{+}22]. Outboard [DS20]. Outcomes [KDO^{+}22]. Ovaries [LCIL23a]. Overlap [GHA^{+}23, ZAK^{+}23]. Overshoot [MST22]. Overview [BIE^{+}22]. Overwintering [FCMT21, KSJ^{+}20]. Owhi [TCM20]. Oxygen [MMSJ23]. Oxytetracycline [BDHP20, SLP20]. Oyster [WCSL22]. Ozark [CLTJ21, RWS^{+}21, ZWBS21].

p [MB21]. p-Chip [MB21]. Pacific [CFGZ23, HA24, Har20, HLJ^{+}23, KNC^{+}20, KUV^{+}23, LHWS23, LDH^{+}22, LHPS^{+}24, MMJW23, MMSJ23, PHW^{+}23, RG20, RSO23, WTW^{+}23]. Paddlefish [DTK20, SNLD21]. Pahrump [LKP^{+}21]. pains [GBM23]. Paired [CP21b]. Pandemic [TBG^{+}22]. Panel [BVP^{+}21, MBP^{+}22]. Panfish [PSS20]. Panopea [HA24]. Paradigm [GK21, RJ22]. Parameter [ZSLG21]. Parameters [LMO21]. Parasitic [TPA^{+}21, SMB23]. Parentage [DH21, DC22, HDP^{+}22, JSH^{+}21]. Parentage-Based [DH21, DC22, HDP^{+}22, JSH^{+}21]. Parental [RCD22]. Parental-Based [RCD22]. Participation [PWLH21]. Partly [STP^{+}21]. partners [LHWS23]. party [CR23]. Pass [FPR20, OTBB23]. Pass-Through [FPR20]. Passage [HSTH20, OTBB23, RFK^{+}20, RRL^{+}21, SWTH21, SBB23, WSG^{+}23, CFGZ23, PTFZ23]. Passive [DWSM21, OTBB23, PREBK22, RPEF^{+}23, SG21, SBB23, TRB^{+}21]. Patent [WCSL22]. Path [MNA^{+}20]. pathway [SPM^{+}24]. Patterns [AZL^{+}21, BHS^{+}23, Boo20, CSQ^{+}21, DHWP21, EP21, GHA^{+}23, HGM22, KFG22, MWR21, MML^{+}22, MWMH22, SCMJF20, SLL^{+}21, SG21, SM22, WD21, WB21]. Paupier [RGMH20]. PBDEs [HBTE22]. Pearl [SLL^{+}21]. Peck [SNLD21]. Pectoral [LLA^{+}22, LSSB23]. Pedigree [HRS^{+}20, KFG22]. Pedigrees [HHH^{+}22, HHH^{+}23, Ano23g]. Pelagic [NMS^{+}23]. Pelagic- [NMS^{+}23]. Pelagically [PBK^{+}23]. Pelican [MBW21]. Pellets [KDM^{+}23]. Pen [CLBL22, RPJ^{+}21]. Pen-Acclimated [CLBL22]. Penaeid [TSBJ20, TJ22]. Peninsula [MSLP22]. Pennsylvania [SMP^{+}21, CJN^{+}21, RRL^{+}21]. Penobscot [PTFZ23]. pens [CBKH23]. Perceived [TBIJ21]. Percent [BBC21]. percentages [PO24]. perception [NKG23]. Perceptions [GvRSS21]. Perch [BFLI22, DNW^{+}22, GWM21, HRK^{+}22, HBBS22, MSV^{+}21, SSH22, SWJ^{+}21]. Percid [BIE^{+}22]. Performance [ELCM22, LMO21, TRB^{+}21, JJD^{+}23]. Period [DPR^{+}20, GBM23]. Permeable [TM21]. Permeable-Style [TM21]. Persistence [BS21, CR23, HFJ23]. Perspectives [HG22, YSS22]. Pest [CJP^{+}21, CGD^{+}21, FHA^{+}21, GK21, HIF21, HMB21]. Petroleum [GRM^{+}21]. pH [KSPP22]. Phase [BWM22]. Phenology [CSWP23a]. Phenotypic [LSB^{+}21]. phylogenetic [ACGF23]. Phylogeographic [RM21]. Phylogeography [WAKO21]. Physical [SSF^{+}23]. Physiological [CAZ^{+}21, GGW21, KSPP22, LHB^{+}24]. Physiology [ELCM22]. Pike [LCF20, PCM^{+}21, SGH^{+}21, SM22, TCL^{+}21, MLB^{+}23]. Piscivore [Mic20a, STP^{+}21]. Piscivores [GHA^{+}23, GW21]. PIT [FPR20, HCO^{+}24, MB21, PE21, ZWBS21]. PIT-Tagged [ZWBS21]. placement [DSW^{+}24]. Placopecten [LOC23]. Plains [BKE22, NMS^{+}23]. Planning [KS21]. Plant [CBS21]. Plantings [NHC^{+}20]. Plasticity [LCIL23b, SSS^{+}22]. Platforms [GRM^{+}21]. Ploidy [FJH^{+}22]. Poles [DHWP21]. Policy [HIF21]. Polybrominated [HBTE22]. Polymorphism [BVP^{+}21, MBP^{+}22]. Pond [CM22]. Poolfish [LKP^{+}21]. Pools [AWRL23]. poor [KR24]. Population [AWRL23, BOHA20, BBKJ21, BOB^{+}21, COT20, CHO24, CFPSW21, DGJ^{+}22, DBCS20, DEO20, DWGR21, FJH^{+}22, FLW21, HSP21, HDH^{+}20, HSS^{+}20, HDT^{+}22, HJBP20, HSH21, KUV^{+}23, Kep21, KQS^{+}20, LHPS^{+}24, LMA23, MDA21, MSV^{+}21, RBFQ20, RSK^{+}21, RMR^{+}22, RCR21, SL21, SHWW22, SPGP21, SSSP22, SW22, TLSP20, TCB^{+}21, TPC21a, TPC21b, WDPH22, WAKO21, YBH24, AMR^{+}23, CBB^{+}23, CEB^{+}23, EFK^{+}23, KR24, MDD^{+}23, SCW^{+}24, TO23]. Population-Level [SW22]. Populations [AWO21, ACRB22, BIE^{+}22, BMAM22, HRK^{+}22, HBBS22, MSM21, MVM^{+}23, MSL^{+}20, MS22, MKK^{+}21, MLB21, NKC^{+}23, NF21, PCM^{+}21, PTCZ22, SPP^{+}21, Sin21, SHWW22, WW21, WSK^{+}22, ZK22, HFJ23, HS24]. Port [NLW^{+}20]. Portable [MLB21]. Postflood [TJ22]. Postocular [BR21].

Postrelease [CRH^{+}22, CLBL22, FTLR23, LTB^{+}22, LDH^{+}22, RRB20, RSM^{+}23, MLB^{+}23]. Postspawn [FCMT21]. Poststocking [ESF23, WWBM20, WW20]. pot [BSKSK24]. pot-based [BSKSK24]. Potamopyrgus [GET22]. Potential [Boo20, CTJS^{+}22, DC22, HSH21, IDH^{+}23, LCIL23b, MDA21, MK21, RZK^{+}21, STB^{+}21, SLE21, SFFW21, ZJL^{+}21, JJD^{+}23, MBF^{+}23, WA23]. Potomac [KHB22, HGM22]. Practices [BKM21, GRA^{+}22, GGW21, LLE22]. prairie [WMB23, SBW^{+}23]. Prairies [CVD^{+}20]. Precision [KHBF24, MS22, ORL^{+}22, PSW23]. Predation [BTZ^{+}21, BSC^{+}21, BKM21, EPB^{+}22, GW20, HEP^{+}23, PFGE23, GWRF24]. Predator [CHJ^{+}22, EPB^{+}22, HLM^{+}23]. predators [PK24]. Predatory [AAW^{+}23, MSL^{+}20]. Predict [SPP^{+}21, SHI22]. Predicted [BLS23]. Predicting [DNW^{+}22]. Prediction [HJ21]. Predictive [CFCN20]. Predictor [HW22]. predictors [SSM^{+}24]. preliminary [SFM23]. Preparation [SMM20]. Preservation [CHL22]. Prespawn [NKC^{+}23, Gep23, SCDS23]. Prevalence [RAHS20]. Prevent [TM21]. previously [ACGF23, KSP^{+}24]. Prey [BTZ^{+}21, CHJ^{+}22, HMO^{+}21, PBD^{+}22]. Price [Ano22a, PHG^{+}23]. Primary [KDM^{+}23]. Prior [MNA^{+}20]. Prioritize [BDG^{+}23]. Probability [DKEI20, HMP22, Kin22, LLE22]. Probiotics [MMJW23]. Process [LRM^{+}20]. Processing [SB21a]. Produce [BBO20, CEH^{+}20]. Produced [BHT^{+}21, MWMH22]. Productivity [PST^{+}20, CHB^{+}23]. Profiles [PHG^{+}23]. Program [FHA^{+}21, HG22, HW22, KFG22, LGOM22, PWLH21]. Programs [DC22, NLW^{+}20]. Project [TM21]. projection [CDG^{+}23]. Promise [IPZK21]. Promote [KBC^{+}20, GBM23]. Prone [FPR20]. Proportions [JWG^{+}21]. Proposed [BBFQ21, BB21a, BBMC23, GBJ21, MRW^{+}24]. Protocol [LCF20]. Protocols [Har20]. Protracted [WWK^{+}23]. Proves [TVG^{+}21]. Provide [NMW^{+}21]. Provides [KFG22]. Proximity [BHCC23, PJD^{+}21]. Public [NLK21, RSO23]. Puerto [CRNRFL24]. Puget [CLM20]. Pulsed [CMBR20]. pump [WPM^{+}23]. Put [BM20, CM22, HBT^{+}21, NLK21]. Put-and-Take [BM20, CM22, HBT^{+}21].

Quagga [CGD^{+}21]. Quality [FB20, KJB^{+}22, Owe21, TPC^{+}22, BSR^{+}23]. Quantified [HDP^{+}22]. Quantifies [EMRK20]. Quantify [RBFS20]. Quantifying [HS24, HGH^{+}21, RHFU24, SFFW21, WGBK22]. Quantity [FB20, BSR^{+}23]. Quatrefoil [KBDI21, AJHB23]. Québec [CVD^{+}20]. Question [TPA^{+}21]. Quick [LCIL23c].

Radiocarbon [DAS20]. Radiotelemetry [SFFW21, WWBM20, SCDS23]. Rainbow [AWF23, BS21, BR21, CRH^{+}22, ESF23, Owe21, RWS^{+}21, TCB^{+}21, WB21, JJD^{+}23, SMB23]. Raised [HLDK21]. Raised-Footrope [HLDK21]. Ramp [RFK^{+}20, SFM23]. Ramps [DDT^{+}20]. Random [GWMP23, HW22]. Range [AWG^{+}20, RSM^{+}23, SDL^{+}23, SMP^{+}21, RHFU24]. Rangewide [BVP^{+}21]. Rapid [KNC^{+}20, KHD^{+}22, LCF20, MSD^{+}22]. Rapidly [DS20]. Rate [DH21, GRA^{+}22, KQS^{+}20, LHR^{+}20, LSP20, SB21b, MDD^{+}23]. Rates [BM20, CM22, DPR^{+}20, MdAM21, MSP21, RGMH20, RRB20, SPP^{+}21, SWM22, SKGL22, SPY^{+}20, GSML23, GWRF24, LJB^{+}24]. Ratios [FMM20]. Ray [IPZK21]. Rays [LMA20, LSSB23, MCB20]. Razorback [FDG^{+}21, HUTB21, dVB20a, dVB20b]. Re [KMS21]. Re-examination [KMS21]. Reach [JDO^{+}20, AAW^{+}23]. Reach-Scale [JDO^{+}20]. Reaches [CSQ^{+}21]. Reared [FDG^{+}21, LMN^{+}20, LMA20, PE21, CBKH23, HGH^{+}23]. Rearing [MNA^{+}20]. Recapture [DWGR21, OMAA22, RJ22, SHI22, ZSLG21, HS24, SWA^{+}24]. Recently [ORWS21]. Recolonization [RG20, TM21]. Recommendations [KR23]. Recompression [RSM^{+}23]. Reconnection [RMWK23]. Records [MSLP22]. Recover [HRM^{+}22, PST^{+}20]. Recovery [BDG^{+}23, KBC^{+}20, KUV^{+}23, KS20, MNA^{+}20, PCM^{+}21, PA21, LHB^{+}24, RH23, SCP^{+}23]. Recreation [BBO20, GDS^{+}22]. Recreation-Grade [BBO20]. Recreational [BHT^{+}20, BDC22, CLRL22, DDT^{+}20, GvRSS21, JSH^{+}21, KBC^{+}20, KBDA21, LRM^{+}20, MML^{+}22, MAB^{+}20, NLK21, PG20, WM21, vPM20, CR23, FTLR23, GSML23]. Recreational-Grade [LRM^{+}20]. Recreationist [GDS^{+}22]. Recruitment [BFLI22, BFM^{+}22, BKM21, DBS21, FFG20, GBJ^{+}22, HSH21, KDM^{+}23, MGMP21, BHB^{+}23, CEC^{+}24, KR24, PAE^{+}23]. Red [BCW^{+}21, GRM^{+}21, MAB^{+}20, PG20, SDF^{+}21, SCW^{+}24, SPSH20, WMB23]. Redbreast [BBMC23]. Redear [CH22, PSS21]. Redeye [LEHP21]. Redistribution [PC20]. Redside [SLE21]. Reduce [TCL^{+}21, WB24]. Reduced [ELCM22]. Reduces [BTZ^{+}21, CRH^{+}22, HLDK21]. Reducing [GvRSS21, TPA^{+}21]. Reduction [RCD20]. Reef [BCW^{+}21, GKRH21, KDM^{+}23, RRB^{+}22, WSK^{+}22]. Reefs [BCW^{+}21, JCB^{+}21, WCSL22]. Reentrainment [AWZ^{+}20]. Reestablished [SHWW22]. Reflects [SGM^{+}20]. Refuge [HGM22, MKG^{+}22, MDD^{+}23]. Regeneration [LCF20]. Regime [SMM^{+}22]. Regimes [BKE22]. Region [FNT21]. Regional [AOB^{+}22]. Regulated [BBKJ21, ESCI21, MML^{+}22, SMM^{+}22]. Regulation [BSKM20, CFPSW21, KDO^{+}22]. Regulations [CLRL22, MKKN22, MK21, NMBM21, ORWS21, SL21, SHWW22]. Reintroduced [HSH21, HMO^{+}21]. Reintroduction [MRM^{+}21]. Related [CH22, CDG^{+}23]. Relating [CWB^{+}21, HIF21]. Relation [EP21, HQWD20, LBH21]. Relationship [COT20, CDB22, GDS^{+}22, RSO23]. Relationships [KNK22, OS20, PAE^{+}23, WMB23]. Relative [CR23, HBT^{+}21, KBDI21, PG20, FHB^{+}24]. Release [BCS^{+}21, BBKJ21, BDC22, ERT21, LMN^{+}20, MDBC21, MML^{+}22, SLG^{+}23, SGH^{+}21, CBKH23, LJB^{+}24, MSSS23, SMB23]. Released [MGNP22, RMB^{+}20, RRB^{+}22, KHL^{+}24]. Releases [BAH^{+}20]. reliable [RH23]. Relies [PST^{+}20]. Remediation [GKRH21, HKPP20]. Remote [AKC20, MRM^{+}21, SCN^{+}21]. Removal [Bar21, HCP^{+}23, MSL^{+}20, RLS^{+}23, ZMA20]. Removals [SVS^{+}21, SPGP21]. Repelled [DS20]. Reporting [MAB^{+}20, PWLH21, TPC^{+}22, GSML23]. Reproducing [HSS^{+}20]. Reproduction [ACRB22, CSWP23a, CSWP23b, HBR^{+}23, PBK^{+}23, RPEF^{+}23, CEB^{+}23]. Reproductive [GLSF21, KDB^{+}20, LCIL23b, NF21, SMCEF^{+}23, SBD^{+}20, SWA^{+}24]. required [SDR24]. Rescue [BBB^{+}21]. Research [LHB^{+}23, MS21a]. Researchers [BPT21]. Reservoir [BPL^{+}21, BBKJ21, BKM21, DBS21, FFG20, HBR^{+}23, HCP^{+}23, Kin22, KDO^{+}22, LCM^{+}23, MSP^{+}22, MVM^{+}23, MS22, NHC^{+}20, OTBB23, PBD^{+}22, RMWK23, SDL^{+}23, SK23, SNLD21, SKWR23, SPP^{+}21, SBF^{+}23, TPC21b, VF21, WW21, JJD^{+}23, SSM^{+}24, SSPL20, SSSP22]. Reservoirs [ECC22, HN21, KNC^{+}20, MKC^{+}23, MK21, NMBM21, PSS20, SK23, SSF^{+}23, SWJ^{+}21, TPC21a, WW20]. Resetting [RMWK23]. Reshaped [RMWK23]. Residency [AS20]. Resident [HMO^{+}21]. Resiliency [DBCS20]. Resins [SMM20]. Resistance [BD22]. Resistant [ESF23]. Resource [BPT21, GBM23]. Resources [PBD^{+}22]. Response [BGC^{+}21, CVD^{+}20, FNB^{+}22, MSZ^{+}23, RKP^{+}22, SHWW22, SSSP22, VF21, YZJ^{+}21, SCP^{+}23, YBB^{+}23]. Responses [GGW21, KES^{+}21, KSPP22, SSS^{+}22, SPGP21, SKGL22, LHB^{+}23, NMS^{+}23, SSR^{+}24, TO23]. Restoration [BAH^{+}20, GGWV22, HRM^{+}22, HGMR21, PC20, RKP^{+}22, CFGZ23, Gep23, LKD23]. Restore [PST^{+}20]. Restored [WCSL22]. Restrictive [BSKM20]. Restructure [Mic20a]. Results [FJH^{+}22]. retained [MLB^{+}23]. Retention [BR21, CRH^{+}22, DWSM21, OTW^{+}22, SPSH20, ZSLG21, dVB20a, CFB^{+}24, DSW^{+}24]. Retracted} [BM20]. Retraction [Ano23g]. Return [TPC^{+}22, HLJ^{+}23]. Returning [BB21b]. Returns [MR21b, MWMH22]. Reveal [BBP20, LKP^{+}21, MSLP22, WWK^{+}23]. Revealed [AFBRS22, DRPW21]. Reveals [HYM^{+}22, KUV^{+}23, LTL^{+}23, SG21, ACGF23, KHL^{+}24, LOV^{+}23]. Review [HRK^{+}22, HEP^{+}23, KR23, LCM^{+}23, NKK21, ZWBS21, ZZWX20]. Reviewer [Ano20g, Ano21i, Ano22h, Ano23h]. Revisited [BBC21]. Reward [HG22]. richness [SDR24]. Rico [CRNRFL24]. Ridges [LLA^{+}22]. Ring [OTW^{+}22]. Rio [COT20]. Risk [AWZ^{+}20, BDG^{+}23, GvRSS21, HG22, PCM^{+}21, RAHS20]. Risks [HG22]. River [ARK^{+}21, Ano23g, BHS^{+}23, BHT^{+}21, BQR^{+}20, CBLW21, CLSM20, CP21a, CSQ^{+}21, CFPSW21, CBB^{+}23, CFGZ23, CFHS22, DBS21, DTK20, EP21, ESCI21, FM20, FTLR23, FCMT21, GGWV22, HTH^{+}21, HSP21, HRM^{+}22, HBV^{+}21, HGH^{+}21, HL23, HGM22, HGMR21, HSH21, HHH^{+}22, HHH^{+}23, JDO^{+}20, KFG22, KJB^{+}22, LSB^{+}21, LGQ^{+}21, MWR21, MGMP21, MQK^{+}20, MNA^{+}20, MWMH22, PA21, PE21, PTFZ23, PC20, PTCZ22, Rac21, RKP^{+}22, RSK^{+}21, RPEF^{+}23, SL21, SCMJF20, SLL^{+}21, SMM^{+}22, SBF^{+}23, SPSH20, SKGL22, SPY^{+}20, TO23, TPA^{+}21, TCB^{+}21, WTW^{+}23, WDPH22, WRH^{+}21, ZAK^{+}23, HS24, SLD^{+}23, AKC20, AAW^{+}23, AWRL23, BB21b, CSWP23a, CSWP23b, CWB^{+}21, CHB^{+}23, CDBC22, DHWP21, ELCM22, EPB^{+}22, FFG20, FLW21, GKT^{+}22, HAP^{+}22, HSS^{+}20, HLJ^{+}23, HCM^{+}22, JJD^{+}23, JCB^{+}21, Kep21, KHB22, KJB^{+}22, LTL^{+}23, LJB^{+}24, MST22, NKC^{+}23, OBR^{+}20, PST^{+}20]. River [SNLD21, SMCEF^{+}23, Sin21, SG21, TRB^{+}21, TPA^{+}21, UHB^{+}23, WMB23, WA23, WWK^{+}23, YZJ^{+}21]. Riverine [BCS^{+}21, KR24]. Rivers [FPR20, KHJS22, MSLP22, OTBB23, SMP^{+}21, WM21, SDR24, SKBB24, YBH24, SWTH21]. riverscapes [PKSG23]. Rivière [CVD^{+}20]. Rivière-des [CVD^{+}20]. Road [RRL^{+}21]. Rock [RFK^{+}20, SFM23]. Rock-Ramp [RFK^{+}20]. Rockfish [AFBRS22]. Rocky [GKRH21]. Roles [MBCH22]. rotenone [CEC^{+}24, LKD23]. Round [BOB^{+}21]. Run [FHK23, HAP^{+}22, JWG^{+}21, SFFW21, KHL^{+}24].

S [Ano22a, HBR^{+}23]. Sacramento [HAP^{+}22]. Safety [WSG^{+}23]. Sagittal [BS20]. Sales [TBG^{+}22, ZLM21]. Salmon [BPL^{+}21, BBB^{+}21, BHT^{+}21, BB21b, BSC^{+}21, BW22, CHJ^{+}22, CEH^{+}20, CLBL22, CFCN20, CFHS22, CDBC22, DPR^{+}20, ELCM22, FM20, FHR21, FLW21, FMM20, FHK23, GWRF24, GKT^{+}22, Har20, HG22, HAP^{+}22, HWCS24, HL23, HLM^{+}23, HJ21, HHH^{+}22, HHH^{+}23, HMO^{+}21, JSH^{+}21, JWG^{+}21, KHL^{+}24, KUV^{+}23, KSJ^{+}20, KES^{+}21, MBP^{+}22, MGNP22, NKC^{+}23, PFGE23, PO24, PST^{+}20, PC20, RPJ^{+}21, RJ22, RAHS20, RSK^{+}21, SKA21, SKBB24, SFFW21, TRB^{+}21, WRH^{+}21, Ano23g, FTLR23, ZAK^{+}23, ELCM22]. Salmonid [KHJS22, PS21, SWTH21, TSJ20, TPA^{+}21, ROP^{+}23]. Salmonids [HEP^{+}23, NRJ^{+}20]. Salvage [HLSW23]. Salvaged [AWZ^{+}20]. Sample [MS21b, SB21b]. Sampling [BOHA20, CBC^{+}20, CHS^{+}22, DHWP21, HCS^{+}21, HW22, MR21a, MS22, NLW^{+}20, PSS21, RPEF^{+}23, SWJ^{+}21, TCB^{+}21, AJHB23, BSKSK24, WPM^{+}23, WHA23]. San [BSC^{+}21]. Sandbar [AS20]. Sandusky [HBV^{+}21]. Satisfaction [BM20, CM22, JOH^{+}22, SDF^{+}21]. Saugeye [BKM21, Kin22]. Saves [GC21]. Scale [BCW^{+}21, BGC^{+}21, HW22, JDO^{+}20, RCD22, SR20, SM22]. Scale-Based [SR20]. Scale-Derived [RCD22]. Scales [CLM20, Har20, HQWD20, PFGE23]. Scallop [WKM^{+}21, LOC23]. Scan [BBO20, LRM^{+}20, SGK^{+}22]. Scenarios [RBFQ20]. science [LHB^{+}23]. Screen [RZA^{+}21]. Screening [HW22]. screw [WPM^{+}23]. Sea [KKMA23, RRB20, RSM^{+}23, WKM^{+}21, LOC23, PTFZ23, GLSF21, Gep23]. Search [HSTH20]. Season [AZL^{+}21, MPM21, SPP^{+}21]. Seasonal [HGM22, RWS^{+}21, SSM^{+}24, HA24, RHFU24]. Seatrout [AOB^{+}22]. Second [IPZK21]. Section [BIE^{+}22, CP21a, GK21, SDB^{+}20, PBEK23]. Sectioned [LKP^{+}21]. Sediment [GKRH20]. Sedimentation [PKTZ20]. Segments [HSP21, STB^{+}21]. Selection [HCP^{+}23, MDHQ22]. Selective [GW20, OTBB23, FTLR23]. Selectivity [RAB^{+}20]. Self [MAB^{+}20]. Self-Reporting [MAB^{+}20]. Semiautomated [LRM^{+}20]. Sensing [AKC20]. Sensitive [KUV^{+}23]. Sequencing [MBP^{+}22]. Sets [Rac21]. Setting [LSP20]. Sex [Ano22a, BB21b, LLA^{+}22, MSD^{+}22, WPGC21, SYHF24].

Sexual [AMR^{+}23, LKP^{+}21]. sexually [MMSJ23]. Shad [CVD^{+}20, CHL22, KNK22, MWR21, MGMP21, MWMH22, STP^{+}21]. Shallow [ASDW23, SSH22]. Shark [HCC^{+}21]. Sharks [AS20, GC21]. Sharpnose [AS20]. shasta [RAHS20]. Shell [OS20]. Shift [RJ22, SSH22]. Shifts [WZG22]. Shiners [SLE21]. shoal [PAE^{+}23]. shoot [MSSS23]. Shore [RH20]. shoreline [CEC^{+}24]. Short [AWF23, CSQ^{+}21, GBM23, HWCS24, LTB^{+}22, MSSS23, SMB23]. Short-Term [AWF23, CSQ^{+}21, LTB^{+}22, GBM23, HWCS24, MSSS23, SMB23]. Shortnose [EPB^{+}22, KS20, KJB^{+}22]. Shovelnose [HSH21]. show [HA24, NMS^{+}23]. show-factors [HA24]. Shows [IPZK21]. Shrimp [PG20]. Sibship [KDB^{+}20]. Side [BBO20, LRM^{+}20, SGK^{+}22]. Side-Scan [BBO20, LRM^{+}20, SGK^{+}22]. Signal [PA21]. Signatures [LMA20]. Significantly [HLDK21]. Silver [AAW^{+}23, CCL^{+}22, RGMH20, RLS^{+}23, RAH^{+}23, SHFP20, WDPH22, WWK^{+}23, EFK^{+}23, MBF^{+}23, McK23, PKSG23]. Silvery [COT20]. Similar [BFLI22]. Similarities [CEH^{+}20]. Simple [BB21b, Kin22, MW20]. Simulated [COT20, HFJ23, KSPP22, MSM21, MSP^{+}22, SL21]. Simulation [HJBP20, NLW^{+}20]. Simulation-Based [NLW^{+}20]. Simulations [TPA^{+}21]. sinensis [ZJL^{+}21]. Single [BVP^{+}21, MBP^{+}22, PTCZ22]. Single- [PTCZ22]. Site [CGD^{+}21, MRM^{+}21, SCN^{+}21, SWA^{+}24]. Site-Specific [CGD^{+}21]. Situ [CP21b, SHFP20]. Situk [CLSM20]. Size [BLS23, BWM22, CSLW22, CH22, COT20, CLTJ21, CHL22, DGJ^{+}22, GW20, HJBP20, MR21b, MS21b, MK21, OTBB23, OS20, SB21b, SW22, TPC21b, WM21, CBKH23, FHB^{+}24]. Size-Based [HJBP20]. Size-Dependent [SW22]. Size-Selective [GW20, OTBB23]. Size-Specific [WM21]. Sizes [BR21, OTW^{+}22]. Skagit [JJD^{+}23]. Skaha [HMO^{+}21]. Skein [GWM21]. Slope [Rac21]. Slot [KBC^{+}20]. Small [CMBR20, DWSM21, DRPW21, GGWV22, HN21, HBR^{+}23, Mic20a, Mic20b, MB21, PSS21, SFFW21, TPPC20, CEC^{+}24, SKBB24]. Small-Bodied [DWSM21, MB21]. Smallmouth [AZL^{+}21, BMAM22, ERT21, KR24, LSSB23, MQK^{+}20, SR20, SWA^{+}24]. Smelt [BSC^{+}21, TCB^{+}21]. Smolt [Boo20, CFHS22, FHK23, PS21, PST^{+}20]. Smolt-To-Adult [PST^{+}20]. Smolts [CHJ^{+}22, SWTH21, HWCS24, KHL^{+}24]. Snails [GET22]. Snake [SWTH21, CDBC22, GKT^{+}22, LJB^{+}24, TRB^{+}21]. Snapper [BCW^{+}21, GRM^{+}21, MAB^{+}20, PG20, SDF^{+}21, SCW^{+}24]. Snook [AWG^{+}20]. Snoqualmie [TO23]. Snow [GLSF21, GBM23]. Soak [GC21]. Social [TBG^{+}22]. Sociodemographic [ZLM21]. Sockeye [CHJ^{+}22, CFHS22, DPR^{+}20, FM20, GKT^{+}22, HMO^{+}21, KES^{+}21, RPJ^{+}21]. Soft [MGNP22]. Soft-Released [MGNP22]. Sonar [BBO20, EMRK20, FM20, HL23, LRM^{+}20, SGK^{+}22, NKG23]. Sound [DS20, MWR21, CLM20]. Source [PJD^{+}21]. Sources [SKA21]. South [DNW^{+}22, SL21, BSKM20, CLM20, KHB22]. Southeast [HCC^{+}21, JWG^{+}21]. Southeastern [SKGL22, SCW^{+}24]. Southern [BCS^{+}21, HN21, SDWM21, CEB^{+}23, NMS^{+}23, Bar21, Boo20, JOH^{+}22, RG20, SCL^{+}20]. Southwest [YZJ^{+}21]. sp [PJD^{+}21]. Space [HBV^{+}21, SKA21, RHFU24]. Spanning [Mun22]. sparse [CDG^{+}23]. Spatial [BHCC23, FFG20, GLSF21, GBM23, HQWD20, MDHQ22, PFGE23, Rac21, SMCEF^{+}23, SLL^{+}21, SBF^{+}23, TSBJ20, TJ22, WDPH22, WGBK22, ZAK^{+}23, SCW^{+}24, SSM^{+}24]. Spatiotemporal [HAP^{+}22, PBK^{+}23, SDWM21, WD21]. Spawn [BB21b]. Spawning [AKC20, BBP20, DPR^{+}20, GKRH21, HUTB21, HSF^{+}22, HRS^{+}20, IDH^{+}23, KDM^{+}23, LCIL23a, MWR21, PBK^{+}23, RBFS20, SM22, SFFW21, WWK^{+}23, MBF^{+}23, NMS^{+}23, SLD^{+}23, WB24, ZAK^{+}23]. Spawns [LTL^{+}23]. spear [CRNRFL24]. Special [BIE^{+}22, CP21a, GK21, SDB^{+}20, PBEK23]. Specialization [TBIJ21]. Specialized [KS20]. Species [AWG^{+}20, EPB^{+}22, GvRSS21, HMP22, HL23, KNK22, KBDA21, LSB^{+}21, MNA^{+}20, MB21, MKK^{+}21, OBR^{+}20, SB21a, SDL^{+}23, SCMJF20, TSBJ20, SDR24, SPM^{+}24, WA23]. Specific [CGD^{+}21, PFGE23, RCD20, WM21]. Specimen [SMM20]. specimens [BSR^{+}23]. Spill [CVD^{+}20]. Spinal [CMBR20]. Spores [RAHS20]. Sport [OBR^{+}20, SPP^{+}21, SSR^{+}24, HFJ23]. Spotfin [DEO20]. Spotted [AOB^{+}22, BBMC23, LSP20]. spp [AWG^{+}20, ACGF23, LEHP21, MSD^{+}22]. Spread [GET22]. Spring [CDBC22, HHH^{+}22, HHH^{+}23, LGOM22, NKC^{+}23, Ano23g, FTLR23]. Spring/Summer [HHH^{+}22, HHH^{+}23, Ano23g]. St [CBB^{+}23, UHB^{+}23]. St. [BHT^{+}20, JCB^{+}21]. Stabilization [GRA^{+}22]. Stable [FMM20, MR21a, RGHF23, SSPL20, WZG22]. Stages [KUV^{+}23, AJHB23]. Stakeholders [BDG^{+}23]. Standard [BBFQ21, BB21a, BBMC23, GBJ21, RAB^{+}20, SWJ^{+}21, MRW^{+}24]. Standardized [BVP^{+}21]. Standards [SR20]. Standing [HSW21]. State [DEO20, ERCB22, GDS^{+}22, LHB^{+}23, MSLP22, SKA21]. State-Space [SKA21]. States [HCC^{+}21, PKSG23, PTFZ23]. Statewide [HW22, TBG^{+}22]. Station [CVD^{+}20]. statoliths [PHW^{+}23]. Status [CFCN20, FHR21, PAC22]. Steelhead [Boo20, BHCC23, CLSM20, DBCS20, DMBC23, DRPW21, FCMT21, KFG22, MR21b, MDBC21, MSLP22, MST22, PE21, PC20, RCD22, CHB^{+}23, JJD^{+}23, LJB^{+}24, PO24, ZAK^{+}23]. Stem [CSQ^{+}21, PE21, BHS^{+}23]. sterile [HFJ23]. Stock [DH21, DRPW21, FB20, FLW21, HCC^{+}21, HYM^{+}22, IDH^{+}23, JSH^{+}21, KBC^{+}20, MBP^{+}22, SNLD21, SFFW21]. Stocked [BS21, BTZ^{+}21, BDHP20, CLBL22, DKEI20, GW21, HMQM22, HGSS20, KMS21, LHR^{+}20, LLE22, TPPC20, WB21, BHB^{+}23, GWRF24]. Stocking [BKM21, HN21, HW22, KSPP22, LLE22, LMA23, MR21b, MSM21, MWASW20, MNA^{+}20, OMAA22, SKGL22, ZMA20, HFJ23, MDD^{+}23, SBH^{+}23]. Stocks [HSW21]. Storage [BBC21, ROP^{+}23]. Strain [SG22]. Strains [AWF23, ESF23, LHR^{+}20]. Stranded [AWZ^{+}20]. Strategic [KS21]. Strategies [MDBC21, SCW^{+}24]. Strategy [BBB^{+}21, Gep23, SBH^{+}23]. stray [PO24]. Straying [DRPW21, CHB^{+}23]. Strays [DC22]. Stream [Bar21, BOHA20, BM20, BMAM22, CLTJ21, CDB22, CEH^{+}20, DGJ^{+}22, DRPW21, HSW21, HKPP20, KR23, NRJ^{+}20, PST^{+}20, PC20, RCD20, RFK^{+}20, RKP^{+}22, RWS^{+}21, RRL^{+}21, RSO23, WB21, ZWBS21, NMS^{+}23, RH23]. Stream-Dwelling [RCD20]. Stream-Type [PST^{+}20]. Streams [AWF23, CLTJ21, CMBR20, HSW21, HCS^{+}21, MRM^{+}21, NLB21, RH20]. Streamwide [ACRB22]. Strength [DNW^{+}22]. Stress [ELCM22, GGW21]. stringers [MLB^{+}23]. Striped [ARK^{+}21, HL23]. Structural [SSF^{+}23]. Structure [CSLW22, CH22, CKB^{+}21, FLW21, KBC^{+}20, KQS^{+}20, MS21b, MKK^{+}21, MK21, SB21b, SFFW21, TPC21b, WGBK22, WAKO21, EFK^{+}23]. Structured [HMO^{+}21]. Structures [PBD^{+}22, PSW23, KHBF24]. Studies [BPT21, CBC^{+}20, MR21a, RMB^{+}20, TPC^{+}22, WPM^{+}23, WA23]. Study [AWZ^{+}20, BDC22, FNT21, HDH^{+}20, HJBP20, KBC^{+}20, LL21, OTBB23]. Sturgeon [BAH^{+}20, BHT^{+}20, HLDK21, HSH21, HBTE22, IPZK21, LMN^{+}20, LMA20, MNA^{+}20, MPM21, PA21, SLG^{+}23, CBB^{+}23, WA23]. Style [TM21]. Subadult [HLDK21]. subalpine [KSP^{+}24]. Subestuaries [HJB^{+}21]. Sublethal [CCL^{+}22]. Submerged [BBO20]. Submontane [SGK^{+}22]. Subsampled [LMO21]. Subsampling [HJBP20]. Subsequent [PE21]. Substantial [RSO23]. substrate [WA23]. Subtropical [SLL^{+}21]. Subyearling [TRB^{+}21]. Success [BFLI22, ESCI21, GKRH20, MST22, PO22, SBD^{+}20, NKG23]. Successful [HGMR21]. Successfully [WSK^{+}22]. Sucker [CSLW22, FDG^{+}21, HUTB21, KJB^{+}22, RHUM^{+}23, ZMA20, dVB20a, dVB20b]. suckermouth [AMR^{+}23]. Suckers [EPB^{+}22, ZWBS21, CBKH23]. Suffocation [PKTZ20]. Suggestions [KQG23]. Suitability [AKC20, CSWP23b, NLB21]. Suite [GW21]. Summer [CDBC22, HHH^{+}22, HHH^{+}23, JWG^{+}21, Ano23g]. Summer-Run [JWG^{+}21]. Summertime [HGM22]. Sunfish [BBMC23, CH22, PSS21]. Superior [RH20]. Supplemental [HN21, MWASW20, OMAA22, Owe21]. Supplementation [COT20]. Support [BDG^{+}23, KLR^{+}21, McK23]. Suppression [KQG23, MSM21, RBFQ20, RCR21, SGK^{+}22, WGKB20, WGBK22, SCP^{+}23]. Surface [KJB^{+}22]. Surfacing [MSZ^{+}23]. surgery [HWCS24]. Surgical [KS20, LL21]. Surrogate [FNB^{+}22]. Survey [CTJS^{+}22, CHS^{+}22, WSK^{+}22, CRNRFL24]. Surveying [WCSL22]. Surveys [DDT^{+}20, ECC22, GET22, HLM^{+}23, MSV^{+}21, SKBB24, SWA^{+}24]. Survival [ACRB22, AWF23, CLBL22, DKEI20, DPR^{+}20, ESF23, FDG^{+}21, FCMT21, GWM21, HSTH20, HAP^{+}22, HMQM22, HLM^{+}23, HSF^{+}22, KNC^{+}20, KJB^{+}22, LLE22, MW20, PFGE23, PST^{+}20, RZA^{+}21, RRB20, RSM^{+}23, SWTH21, TPPC20, WWBM20, WW20, WB21, CBKH23, FHB^{+}24, HGH^{+}23, HCO^{+}24, KHL^{+}24, MMJW23, ROP^{+}23, SMB23]. Susceptibility [HEP^{+}23]. Susquehanna [SMP^{+}21]. Sustainable [SLZ^{+}20]. sustained [SCP^{+}23]. Suwannee [YBH24]. Sympatric [BBP20, KNK22, NRJ^{+}20, WA23]. symposium [LHWS23, CP21a, SK23]. Synchrony [FFG20]. Synthesis [BPT21, KQG23, LHB^{+}23, SDB^{+}20, ZZWX20]. System [BHT^{+}20, FPR20, JCB^{+}21, LBH21, MQK^{+}20, RJ22, SKWR23, SGM^{+}20, CBB^{+}23, BQR^{+}20]. Systems [ESCI21, HMB21, KDB^{+}20].

T [SCDS23, SPSH20]. T-Bar [SPSH20, SCDS23]. Tactic [CGD^{+}21]. Tag [CCL^{+}22, DH21, FPR20, SG21, TPC^{+}22, ZWBS21, ZSLG21, CFB^{+}24, DSW^{+}24, HCO^{+}24]. Tag-Return [TPC^{+}22]. Tagged [DPR^{+}20, TRB^{+}21, ZWBS21]. Tagger [HWCS24]. Tagging [DH21, DC22, HDP^{+}22, JSH^{+}21, MB21, NKK21, RCD22, DSW^{+}24]. Tags [CCL^{+}22, DWSM21, DRPW21, LDH^{+}22, LFS20, OTBB23, PE21, SPSH20, SBB23, TRB^{+}21, CEB^{+}23, SCDS23]. Tailwater [BS21, SKGL22]. Take [BM20, CM22, HBT^{+}21, NLK21]. Tallapoosa [ESCI21]. Tancing [TBG^{+}22]. Tandem [TPC21b]. Target [MBW21]. Targeted [OMAA22]. Targeting [KBDA21, WGKB20]. Taxa [FNB^{+}22]. Taxonomic [AWG^{+}20]. Technique [TVG^{+}21]. Techniques [GKRH21, HCS^{+}21, RLS^{+}23, RJ22]. Technologies [PREBK22]. Technology [TVG^{+}21]. Telemetered [WGKB20]. Telemetry [AWZ^{+}20, HYM^{+}22, HCP^{+}23, MWR21, RMB^{+}20, SGK^{+}22, HWCS24, KHL^{+}24, RAH^{+}23]. Temperate [SSS^{+}22]. Temperature [GRA^{+}22, LSP20, MPM21, RH23]. Temperatures [SPY^{+}20]. Temporal [DHWP21, GLSF21, KHB22, LGOM22, MDHQ22, Rac21, RSK^{+}21, RMR^{+}22, ZAK^{+}23]. temporarily [MLB^{+}23]. Temporary [CRH^{+}22]. Tensaw [AP20]. Term [AWF23, CBS21, CSQ^{+}21, LTB^{+}22, LMA23, Mun22, PA21, GBM23, HA24, HWCS24, MSSS23, SMB23]. Terrestrial [GK21]. Territory [LLE22]. Test [FM20]. Testing [PKSG23]. Texas [AWO21, CHO24, KDB^{+}20, NMBM21, STB^{+}21, SPM^{+}24, SSM^{+}24, WHA23]. Texoma [SSPL20, SSSP22, TLSP20]. Their [RG20, SDL^{+}23, YSS22, HMO^{+}21, PHW^{+}23]. Thermal [BMAM22, HGM22, RHUM^{+}23]. Thiamine [FHR21, MKG^{+}22]. Thin [LKP^{+}21]. Thin-Sectioned [LKP^{+}21]. Those [GC21]. Thousands [MBP^{+}22]. Threatened [DEO20, HTH^{+}21, MKK^{+}21, CDG^{+}23]. Threats [LGQ^{+}21]. Three [ECC22, GKT^{+}22, NKC^{+}23, NLB21, OTW^{+}22, PSW23, SB21a, SWJ^{+}21, WWBM20]. Thresholds [MSZ^{+}23, YBB^{+}23]. throughout [DPR^{+}20]. Thunder [FB20]. Thyroid [HBTE22]. Tiber [VF21]. Tidal [FM20]. Time [ECC22, GC21, PHG^{+}23, RPJ^{+}21, LKD23]. Time-Lapse [ECC22]. Times [GC21]. Timing [FHK23, MSLP22, SFFW21, CHB^{+}23, SWA^{+}24]. together [LHWS23]. Tongs [WCSL22]. Tool [DHWP21, HW22, Owe21, TPA^{+}21, SYHF24, WB24]. Tournament [CJN^{+}21, CAZ^{+}21, LTB^{+}22, MSW22, SW22]. Tournaments [AZL^{+}21, ERT21]. Towed [KBDI21]. Town [LEHP21]. tracking [KHL^{+}24]. Trade [HG22]. Trade-Offs [HG22]. traditional [WHA23]. Trajectories [SHI22]. Transfer [HDH^{+}20, MKC^{+}23]. Transferred [FMM20]. Transition [RFK^{+}20]. translocation [KSP^{+}24]. Translocations [HSS^{+}20, HLJ^{+}23]. transmitter [CFB^{+}24]. Transmitters [LL21, GWRF24]. Transplantation [Gep23]. Transplanting [KES^{+}21]. Transponder [DWSM21, OTBB23, SG21, SBB23, TRB^{+}21]. Transport [GvRSS21, GGWV22]. Trap [LGOM22, OTW^{+}22, TSJ20, dVB20b, AJHB23]. Trap-and-Haul [LGOM22]. Trapping [LTL^{+}23]. Traps [KBDI21, dVB20a]. travel [CR23]. Traveling [RZA^{+}21]. Trawl [CTJS^{+}22, TSBJ20, TJ22]. Trends [BHT^{+}20, KHB22, Mun22, PWLH21, PA21, Rac21, RSK^{+}21, RMR^{+}22, TPC^{+}22, CHO24, CBB^{+}23]. Trials [GET22]. Tributaries [CSWP23a, HBR^{+}23, HDQ22, RPEF^{+}23, SMCEF^{+}23, SG21, HS24, JJD^{+}23]. Tributary [CSWP23b, HSS^{+}20, HUTB21, WDPH22, YZJ^{+}21]. Trip [CLRL22]. triploid [FHB^{+}24]. Trophic [MBCH22]. Trophy [SL21]. Trotline [MSL22]. Trotlines [MSP^{+}22]. Trout [Ano22a, ACRB22, AWF23, BS21, Bar21, BM20, BVP^{+}21, BDG^{+}23, BPT21, BR21, CJN^{+}21, CHJ^{+}22, CRH^{+}22, CMBR20, CLM20, CDB22, DGJ^{+}22, DWGR21, ESF23, HNM^{+}22, HQWD20, HDQ22, HDT^{+}22, HMQM22, HW22, HSF^{+}22, HKPP20, KLR^{+}21, KSP^{+}24, KDM^{+}23, LL21, LKD23, LGOM22, LFS20, MKG^{+}22, MR21b, MDHQ22, MBW21, MWASW20, MKK^{+}21, Mun22, NMW^{+}21, ORL^{+}22, Owe21, PKTZ20, RCD20, RH20, RWS^{+}21, RRL^{+}21, RSK^{+}21, SCN^{+}21, SGK^{+}22, SG22, SMM^{+}22, SKGL22, TCM20, TBIJ21, TM21, WPGC21, WGKB20, WGBK22, WB21, CDG^{+}23, CEB^{+}23, CR23, JJD^{+}23, LOV^{+}23, MDD^{+}23, RR23, SCP^{+}23, SMB23, TO23]. Truman [HBR^{+}23]. tuna [GSML23]. Turbid [FM20]. Turbine [AWZ^{+}20, CFHS22, WSG^{+}23]. Two [AWF23, BR21, GKRH21, GRA^{+}22, HSW21, HCS^{+}21, KDB^{+}20, MS21a, NLK21, NF21, PSS21, SB21a, WW20, DSW^{+}24, PKSG23, YBH24]. Type [PST^{+}20, PSS20, ZSLG21].

U.S. [CRNRFL24, GSML23, LOC23, MVM^{+}23, SCW^{+}24, SKGL22]. Ultra [HRM^{+}22]. Ultra-Oligotrophic [HRM^{+}22]. Ultrasound [CVD^{+}20]. Umpqua [JDO^{+}20]. Unanchored [GGWV22]. uncertainties [LHB^{+}23]. Uncertainty [PREBK22]. Uncommon [RRB^{+}22]. Undecided [HG22]. Under-Ice [KSJ^{+}20]. Underestimate [MSV^{+}21]. Underestimates [DGJ^{+}22]. Underfished [QBA21]. Underrepresents [LGOM22]. Understanding [BW22, DBCS20, NMW^{+}21, SDF^{+}21]. Underwater [DEO20, HL23, RBFS20, SKBB24]. United [HCC^{+}21, PKSG23, PTFZ23]. Units [PG20]. Unlocking [BSR^{+}23]. Unmanned [TVG^{+}21]. unrecognized [ACGF23]. Updated [EPB^{+}22]. upon [GC21]. Upper [LGQ^{+}21, PO24, TO23, WMB23, AWRL23, BQR^{+}20, CBKH23, CSWP23a, CSWP23b, FCMT21, KJB^{+}22, LTL^{+}23, MNA^{+}20, MST22, RKP^{+}22, SNLD21, SG21, WWK^{+}23, YDM^{+}20]. Upstream [BPL^{+}21, KHD^{+}22, KES^{+}21, LTL^{+}23, MGNP22, TM21, PKSG23]. Urban [CBLW21, HBT^{+}21, KS21]. USA [CFGZ23, BBKJ21, HSP21, HJB^{+}21, LEHP21, MSLP22, PA21, RRB^{+}22, SDR24, SLG^{+}23, SMP^{+}21, TCB^{+}21, WWBM20, ZAK^{+}23]. Use [BCW^{+}21, BSKM20, BBKJ21, HBV^{+}21, HGM22, JCB^{+}21, LBH21, MWR21, MSM21, MML^{+}22, MDA21, MS21b, MSP21, MSV^{+}21, PHG^{+}23, RWS^{+}21, RBFS20, STP^{+}21, SWJ^{+}21, ZWBS21, RHFU24, SCDS23]. Used [LMO21, OTBB23, LCVS23]. Usefulness [HNM^{+}22]. Using [AKC20, AWRL23, BBB^{+}21, BBO20, BSC^{+}21, CFPSW21, CRNRFL24, CFCN20, DAS20, DH21, DEO20, DDT^{+}20, FNB^{+}22, GW21, Har20, HMQM22, HLM^{+}23, HBBS22, HJ21, KBC^{+}20, KSJ^{+}20, KKMA23, LHR^{+}20, LLA^{+}22, LMA20, LDH^{+}22, MR21a, MKC^{+}23, MLB21, NMBM21, OMAA22, PG20, RCD22, RGMH20, RR23, RGHF23, RMB^{+}20, SKA21, SHFP20, SMM20, SBB23, TCM20, WWBM20, WSK^{+}22, ZWBS21, vPM20, GWRF24, HLSW23, MVH^{+}22, PKSG23, WPM^{+}23]. Utah [BBFQ21, HDQ22, HDT^{+}22, MDHQ22, RBFQ20]. Utility [DNW^{+}22, vPM20]. Utilized [RBFS20]. UV [SMM20]. UV-Cured [SMM20].

Validating [SYHF24]. Validation [BS20, LSSB23, RCD22]. Valley [HLM^{+}23]. Values [BMAM22, GvRSS21, YSS22]. Valuing [MKC^{+}23]. Variability [ASDW23, FHK23, SDWM21, SBF^{+}23, WDPH22]. Variable [Mic20b]. Variables [HW22, PSS20]. Variation [AWO21, AOB^{+}22, CLTJ21, FLW21, GLSF21, HAP^{+}22, HHH^{+}22, HHH^{+}23, LGOM22, MSP21, PBK^{+}23, WMB23, Ano23g]. variations [RHFU24]. Varied [ZSLG21]. Vary [YDM^{+}20]. Varying [HJBP20, KDO^{+}22]. Vegetation [BBO20]. Vehicle [TVG^{+}21]. Velocities [SHFP20]. Velocity [MS21b, NRJ^{+}20]. Venting [RRB20, RSM^{+}23]. Versus [PTCZ22, CLM20, DDT^{+}20, GLSF21, MCB20]. via [HDP^{+}22]. viability [SCDS23]. Video [SHFP20]. Videography [RBFS20]. Virginia [CFPSW21, MWMH22]. Visible [BR21, LFS20]. Vision [PKNS21]. visitation [CR23]. Visual [MB21, SKBB24]. Voluntary [PWLH21]. vulnerability [SMB23].

Wabash [HCM^{+}22, SMCEF^{+}23]. Wadeable [ZWBS21]. Walleye [BKE22, BGH^{+}20, BFLI22, BFM^{+}22, CFB^{+}24, CHL22, FJH^{+}22, FFG20, GKRH20, GBJ^{+}22, GW20, GW21, HRK^{+}22, IDH^{+}23, Kin22, KDO^{+}22, LLE22, PCM^{+}21, RAB^{+}20, RCR21, SSS^{+}22, SHWW22, WWBM20, WW21, WZG22, WMJ^{+}24, BHB^{+}23, FHB^{+}24]. Walleyes [BSKM20]. Warmwater [WB21]. Washington [BHS^{+}23, ZAK^{+}23, FTLR23, MSLP22, PC20, TCM20, TO23]. Water [BKE22, CRH^{+}22, KSPP22, KJB^{+}22, MS21b, RZA^{+}21, SHFP20, SPY^{+}20, KPD^{+}24, SSM^{+}24]. Waters [ERCB22, KMS21]. Watershed [HKPP20, SFFW21, LOV^{+}23, PO24, TO23]. Waterway [PBK^{+}23]. Web [MR21a]. Webs [RZK^{+}21]. Webster [Ano22a]. Weigh [LTB^{+}22]. Weigh-In [LTB^{+}22]. Weight [BBC21, BBFQ21, BB21a, BBMC23, GBJ21, HNM^{+}22, HJ21, NLK21, STP^{+}21, MRW^{+}24]. Weirs [WB24]. well [LHB^{+}24]. West [CFPSW21, KQG23]. Western [HRM^{+}22, MBP^{+}22, AOB^{+}22, HSW21, MQK^{+}20, MVM^{+}23, MSD^{+}22, OBR^{+}20, WTW^{+}23].

Westslope [HQWD20, SCN^{+}21, LKD23]. Wet [BBC21, WMB23]. Wetland [EMRK20, SM22, dVB20b]. Where [KMS21]. While [RBFS20]. Whirling [ESF23]. White [MBW21, MKKN22, Mic20b, SWJ^{+}21, ZMA20]. Whitefish [PO22, RMR^{+}22, SG21]. Wild [ACRB22, BHCC23, BDHP20, DBCS20, FJH^{+}22, FDG^{+}21, HMQM22, KLR^{+}21, LFS20, Owe21, RPJ^{+}21, RCD22, SCN^{+}21, FHB^{+}24, HGH^{+}23, WB24]. Wild- [SCN^{+}21]. Wildlife [GDS^{+}22]. Willamette [CFGZ23, NKC^{+}23]. Winter [HAP^{+}22, HCP^{+}23, MSLP22]. Winter-Run [HAP^{+}22]. wire [CEB^{+}23]. Wisconsin [BGH^{+}20, BFLI22, BFM^{+}22, DKEI20, GBJ^{+}22, HGSS20, LHR^{+}20, LLE22, SLG^{+}23, SVS^{+}21, TBG^{+}22]. within [DTK20, ERCB22, HKPP20, RHFU24, SSM^{+}24, SKGL22, SPY^{+}20]. without [PBD^{+}22]. Wood [SPGP21]. Woody [GWM21, SSR^{+}24]. Wyoming [HSH21].

Yakima [FTLR23]. Yangtze [CP21a, LGQ^{+}21, CSQ^{+}21]. Yaqui [HTH^{+}21, SBH^{+}23]. Year [BOB^{+}21, CSQ^{+}21, DNW^{+}22, LKP^{+}21, PC20]. Year-Class [DNW^{+}22]. Yearling [WW20, MMSJ23]. Years [BAH^{+}20, DAS20, Mun22, CFB^{+}24, WMB23]. Yellow [BFLI22, DNW^{+}22, GWM21, HRK^{+}22, HBBS22, MSV^{+}21, SSH22]. Yelloweye [AFBRS22]. Yellowstone [KDM^{+}23, MDHQ22, MBW21, PKTZ20, WGBK22]. Yield [Pan21]. York [PA21]. Young [CSQ^{+}21, PC20]. Young-of-the-Year [CSQ^{+}21, PC20]. Yukon [BB21b, FLW21]. YY [MSM21].

Zealand [GET22]. Zone [RM21, RFK^{+}20, SSR^{+}24]. Zooplankton [HMO^{+}21, SSH22, VF21].

References

Anderson:2023:JSC

[AAW^{+}23] Cory A. Anderson, Rebekah L. Anderson, Jun Wang, Neil Gillespie, Eli G. Lampo, Nerissa N. McClelland, Levi E. Solomon, Rich Pendleton, and James T. Lamer. Juvenile silver carp and bighead carp as forage for predatory fish in the LaGrange Reach of the Illinois River. North American Journal of Fisheries Management, 43 (1):164–175, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Auringer:2023:LCL

[ACGF23] Grace Auringer, Matthew A. Campbell, Pascale A. L. Goertler, and Amanda J. Finger. Lampreys in California (Lampetra spp. and Entosphenus spp.): Mitochondrial phylogenetic analysis reveals previously unrecognized lamprey diversity. North American Journal of Fisheries Management, 43(6):1511–1530, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Armstrong:2022:SES

[ACRB22] Benjamin A. W. Armstrong, Colleen A. Caldwell, Michael E. Ruhl, and Justin H. Bohling. Streamwide evaluation of survival and reproduction of M_{Y Y} and wild brook trout populations. North American Journal of Fisheries Management, 42(6): 1398–1413, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arthur:2022:AYR

[AFBRS22] Donald E. Arthur, Jeffrey A. Falke, Brittany J. Blain-Roth, and Trent M. Sutton. Alaskan yelloweye rockfish fecundity revealed through an automated egg count and digital imagery method. North American Journal of Fisheries Management, 42(4):828– 838, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Adams:2023:LFE

[AJHB23] Catherine M. Adams, M. Tildon Jones, Robert Hatch, and Kevin R. Bestgen. Laboratory and field evaluation of a new lightemitting-diode-lit quatrefoil light trap for sampling fish early life stages. North American Journal of Fisheries Management, 43(5): 1337–1348, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:2020:URS

[AKC20] Yvonne Allen, Kayla Kimmel, and Glenn Constant. Using remote sensing to assess alligator gar spawning habitat suitability in the Lower Mississippi River. North American Journal of Fisheries Management, 40(3):580–594, June 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arend:2023:SDI

[AMR^{+}23] Wesley A. Arend, Rebecca D. Mangold, Christopher L. Riggins, Collin Garoutte, Yeyetzi Rodriguez, Thomas C. Heard, Nick Menchaca, Janaye Williamson, Dusty McDonald, Daniel Daugherty, Monica McGarrity, Kevin W. Conway, and Joshuah S. Perkin. Sexual dimorphism in an invasive population of suckermouth armored catfish: Implications for management. North American Journal of Fisheries Management, 43(6):1735–1749, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. North American Journal of Fisheries Management, 40(1):1–2, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. North American Journal of Fisheries Management, 40(2):305–306, April 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. North American Journal of Fisheries Management, 40(3):525–526, June 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. North American Journal of Fisheries Management, 40(4):805–806, August 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. North American Journal of Fisheries Management, 40(5):1053–1054, October 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. North American Journal of Fisheries Management, 40(6):1353–1354, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2020:RA

[Ano20g] Anonymous. Reviewer acknowledgments. North American Journal of Fisheries Management, 40(6):1553–1556, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:E

[Ano21a] Anonymous. Erratum. North American Journal of Fisheries Management, 41(4):1225, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:IIa

[Ano21b] Anonymous. Issue information. North American Journal of Fisheries Management, 41(1):1–2, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:IIb

[Ano21c] Anonymous. Issue information. North American Journal of Fisheries Management, 41(2):259–260, April 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:IIc

[Ano21d] Anonymous. Issue information. North American Journal of Fisheries Management, 41(3):557–558, June 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:IId

[Ano21e] Anonymous. Issue information. North American Journal of Fisheries Management, 41(4):887–888, August 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:IIe

[Ano21f] Anonymous. Issue information. North American Journal of Fisheries Management, 41(5):1227–1228, October 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:IIf

[Ano21g] Anonymous. Issue information. North American Journal of Fisheries Management, 41(S1):i–iv, October 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:IIg

[Ano21h] Anonymous. Issue information. North American Journal of Fisheries Management, 41(6):1617–1618, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2021:RA

[Ano21i] Anonymous. Reviewer acknowledgments. North American Journal of Fisheries Management, 41(6):1932–1935, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2022:ENG

[Ano22a] Anonymous. Erratum: A Nonlethal Genetic Assay for Discriminating Sex in Brook Trout, by N. Webster, S. Price, C. Gowan, and S. S. Coster, 41(4) 1120–1123, August 2021. North American Journal of Fisheries Management, 42(4):1094, August 2022. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic). See [WPGC21].

Anonymous:2022:IIa

[Ano22b] Anonymous. Issue information. North American Journal of Fisheries Management, 42(1):1–2, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2022:IIb

[Ano22c] Anonymous. Issue information. North American Journal of Fisheries Management, 42(2):237–238, April 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2022:IIc

[Ano22d] Anonymous. Issue information. North American Journal of Fisheries Management, 42(3):475–476, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2022:IId

[Ano22e] Anonymous. Issue information. North American Journal of Fisheries Management, 42(4):813–814, August 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2022:IIe

[Ano22f] Anonymous. Issue information. North American Journal of Fisheries Management, 42(5):1095–1096, October 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2022:IIf

[Ano22g] Anonymous. Issue information. North American Journal of Fisheries Management, 42(6):1379–1380, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2022:RA

[Ano22h] Anonymous. Reviewer acknowledgments. North American Journal of Fisheries Management, 42(6):1665–1667, December 2022.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:IIa

[Ano23a] Anonymous. Issue information. North American Journal of Fisheries Management, 43(1):1–2, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:IIb

[Ano23b] Anonymous. Issue information. North American Journal of Fisheries Management, 43(2):285–286, April 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:IIc

[Ano23c] Anonymous. Issue information. North American Journal of Fisheries Management, 43(3):601–602, June 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:IId

[Ano23d] Anonymous. Issue information. North American Journal of Fisheries Management, 43(4):891–892, August 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:IIe

[Ano23e] Anonymous. Issue information. North American Journal of Fisheries Management, 43(5):1143–1144, October 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:IIf

[Ano23f] Anonymous. Issue information. North American Journal of Fisheries Management, 43(6):1447–1448, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:RMH

[Ano23g] Anonymous. Retraction: Monitoring hatchery broodstock composition and genetic variation of spring/summer Chinook salmon in the Columbia River Basin with multigeneration pedigrees. North American Journal of Fisheries Management, 43(3):889, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2023:RA

[Ano23h] Anonymous. Reviewer acknowledgments. North American Journal of Fisheries Management, 43(6):1799–1801, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2024:IIa

[Ano24a] Anonymous. Issue information. North American Journal of Fisheries Management, 44(1):1–2, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2024:IIb

[Ano24b] Anonymous. Issue information. North American Journal of Fisheries Management, 44(2):245–246, April 2024. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2022:RVG

[AOB^{+}22] Joel Anderson, Zachary Olsen, Nicolette Beeken, Roberta Weixelman, and Mark Fisher. Regional variation in growth and mortality of spotted seatrout in the Western Gulf of Mexico. North American Journal of Fisheries Management, 42(6):1381–1397, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Armstrong:2020:EAG

[AP20] David L. Armstrong, Jr. and Tommy R. Purcell. Evaluation of alligator gar broodfish collections in the Lower Mobile–Tensaw Delta, Alabama. North American Journal of Fisheries Management, 40(3):651–660, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Andrews:2021:HML

[ARK^{+}21] S. N. Andrews, D. H. Roth, K. A. Kidd, T. Linnansaari, and R. A. Curry. Harvest and mercury levels of striped bass in Miramichi River, New Brunswick, Canada. North American Journal of Fisheries Management, 41(5):1490–1498, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Altobelli:2020:MRS

[AS20] Ashley N. Altobelli and Stephen T. Szedlmayer. Migration and residency of sandbar, Atlantic sharpnose, bull, and nurse sharks in the Northern Gulf of Mexico. North American Journal of Fisheries Management, 40(5):1324–1343, October 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Annear:2023:IAV

[ASDW23] Andrew R. Annear, Martin A. Simonson, Philip M. Dixon, and Michael J. Weber. Intra-annual variability of common carp and bigmouth buffalo electrofishing data in shallow lakes. North American Journal of Fisheries Management, 43(2):556–568, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Avila:2023:BAF

[AWF23] Brian W. Avila, Dana L. Winkelman, and Eric R. Fetherman. Biotic and abiotic factors affecting short-term survival of two age-0 rainbow trout strains in Colorado streams. North American Journal of Fisheries Management, 43(3):786–793, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2020:DTG

[AWG^{+}20] Joel Anderson, Damon Williford, Alin González, Chris Chapa, Fernando Martinez-Andrade, and R. Deborah Overath. Demographic, taxonomic, and genetic characterization of the snook species complex (Centropomus spp.) along the leading edge of its range in the Northwestern Gulf of Mexico. North American Journal of Fisheries Management, 40(1):190–208, February 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2021:ELG

[AWO21] Joel Anderson, Damon Williford, and Zachary Olsen. Estuarylevel genomic variation confirms demographic and life history differences among black drum populations in Texas. North American Journal of Fisheries Management, 41(4):1040–1052, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2023:UOC

[AWRL23] Charmayne L. Anderson, Gregory W. Whitledge, Neil P. Rude, and James T. Lamer. Using otolith chemistry to determine early life environments and movement of the emerging bigheaded carp population in Pools 16–19 of the Upper Mississippi River. North American Journal of Fisheries Management, 43(1):126– 140, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Algera:2020:SKS

[AWZ^{+}20] Dirk A. Algera, Taylor Ward, Randy Zemlak, James Crossman, Philip Harrison, Alf Leake, Michael Power, and Steven J. Cooke. Stranded Kokanee salvaged from turbine intake infrastructure are at low risk for reentrainment: a telemetry study in a hydropower facility forebay. North American Journal of Fisheries Management, 40(6):1545–1552, December 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Abrams:2021:DPL

[AZL^{+}21] Alice E. I. Abrams, A. J. Zolderdo, Elodie J. I. Lédée, Michael J. Lawrence, Peter E. Holder, and Steven J. Cooke. Dispersal patterns of largemouth bass and smallmouth bass following early-, mid-, and late-season fishing tournaments in an Eastern Ontario lake. North American Journal of Fisheries Management, 41(5): 1454–1464, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Berkman:2020:GAM

[BAH^{+}20] Leah K. Berkman, Michelle R. Anderson, David P. Herzog, Travis L. Moore, and Lori S. Eggert. A genetic assessment of Missouri’s lake sturgeon after 30 years of restoration releases. North American Journal of Fisheries Management, 40(3):700–712, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barabe:2021:BBR

[Bar21] Russell Barabe. Black bullhead removal from a headwater trout stream in Southern California. North American Journal of Fisheries Management, 41(S1):S64–S70, October 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2021:PSW

[BB21a] Timothy F. Bonvechio and Kimberly I. Bonvechio. Proposed standard weight equation and standard length categories for lake chubsucker. North American Journal of Fisheries Management, 41(5):1375–1383, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bradley:2021:DSM

[BB21b] Catherine A. Bradley and Randy J. Brown. Development of a simple morphometric model to identify sex in Chinook salmon returning to spawn in the Yukon River. North American Journal of Fisheries Management, 41(5):1538–1548, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beebe:2021:EFR

[BBB^{+}21] Brittany A. Beebe, Kale T. Bentley, Thomas W. Buehrens, Russell W. Perry, and Jonathan B. Armstrong. Evaluating fish rescue as a drought adaptation strategy using a life cycle modeling approach for imperiled Coho salmon. North American Journal of Fisheries Management, 41(1):3–18, February 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baltasar:2021:EFS

[BBC21] Regan Q. Baltasar, Erin J. Burge, and Derek P. Crane. Effects of frozen storage on fish wet weight, percent dry weight, and length revisited. North American Journal of Fisheries Management, 41 (6):1744–1751, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Black:2021:PSW

[BBFQ21] Aaron R. Black, Zachary S. Beard, Jon M. Flinders, and Michael C. Quist. Proposed standard weight (W_{s}) equation and length categories for Utah chub. North American Journal of Fisheries Management, 41(5):1299–1308, October 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blank:2021:EHU

[BBKJ21] Aaron J. Blank, Tony J. Barada, Jordan D. Katt, and Jeffrey J. Jackson. Evaluation of habitat use by a catch-and-release regulated flathead catfish population in Branched Oak Reservoir, Nebraska, USA. North American Journal of Fisheries Management, 41(S1):S95–S102, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2023:PSW

[BBMC23] Timothy F. Bonvechio, Kimberly I. Bonvechio, Jennifer Moran, and Steve Crawford. Proposed standard weight (W_{s}) equations and standard length categories for redbreast sunfish and spotted sunfish. North American Journal of Fisheries Management, 43 (2):539–546, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bennett:2020:URG

[BBO20] Daniel L. Bennett, Timothy J. Bister, and Richard A. Ott, Jr. Using recreation-grade side-scan sonar to produce classified maps of submerged aquatic vegetation. North American Journal of Fisheries Management, 40(1):145–153, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barkalow:2020:ORS

[BBP20] Stephani L. Clark Barkalow, Mary A. Brandenburg, and Steven P. Platania. Otoliths reveal spawning ecology and early life history of sympatric catostomids. North American Journal of Fisheries Management, 40(2):415–426, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bauerlien:2021:EAE

[BCS^{+}21] Cory J. Bauerlien, Derek P. Crane, Scott Smith, George Palmer, Tyler Young, and Daniel B. Goetz. Estimates of abundance and extreme catch-and-release exploitation in a southern riverine muskellunge fishery. North American Journal of Fisheries Management, 41(5):1602–1615, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Banks:2021:DCE

[BCW^{+}21] Kesley Gibson Banks, Judson M. Curtis, Jason A. Williams, Jennifer J. Wetz, and Gregory W. Stunz. Designing cost-effective artificial reefs: Fine-scale movement and habitat use of red snapper around a nearshore artificial reef complex. North American Journal of Fisheries Management, 41(6):1850–1862, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bruce:2022:EHC

[BD22] Lewis J. Bruce and Jan C. Dean. Effect of hull cleaning on electrical resistance and anodic fields of direct-current electrofishing boats. North American Journal of Fisheries Management, 42(5): 1296–1300, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burton:2022:MJB

[BDC22] Declan Burton, Andy J. Danylchuk, and Steven J. Cooke. Microfishing for juvenile bluegill: A catch-and-release study on an emerging form of recreational angling. North American Journal of Fisheries Management, 42(2):369–374, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brignon:2023:ESD

[BDG^{+}23] William R. Brignon, M. Brian Davis, Stephanie Gunckel, Jason Dunham, Michael H. Meeuwig, Chris Allen, and Shaun Clements. Engaging stakeholders to develop a decision support model of conservation risk and management capacity to prioritize investments in bull trout recovery. North American Journal of Fisheries Management, 43(3):821–838, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bunch:2020:EEO

[BDHP20] Aaron J. Bunch, Brady P. Donovan, Frederick J. Hoogakker, and Eric Peatman. Empirical evaluation of oxytetracycline and F_{1} genetics to differentiate stocked from wild largemouth bass. North American Journal of Fisheries Management, 40(3):713–717, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brandt:2022:SEC

[BFLI22] Ethan J. Brandt, Zachary S. Feiner, Alexander W. Latzka, and Daniel A. Isermann. Similar environmental conditions are associated with walleye and yellow perch recruitment success in Wisconsin lakes. North American Journal of Fisheries Management, 42(3):630–641, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Broda:2022:BCI

[BFM^{+}22] Steven P. Broda, Zachary S. Feiner, Joseph T. Mrnak, Stephanie L. Shaw, and Greg G. Sass. Black crappie influences on walleye natural recruitment in Northern Wisconsin lakes. North American Journal of Fisheries Management, 42(5):1202–1214, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beakes:2021:LSF

[BGC^{+}21] Michael P. Beakes, Cory Graham, J. Louise Conrad, James R. White, Michael Koohafkan, John Durand, and Ted Sommer. Large-scale flow management action drives estuarine ecological response. North American Journal of Fisheries Management, 41 (1):64–77, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boehm:2020:IAW

[BGH^{+}20] Hadley I. A. Boehm, Jason C. Gostiaux, Gretchen J. A. Hansen, Joseph M. Hennessy, and Daniel A. Isermann. Indexing age-0 walleye abundance in Northern Wisconsin lakes before fall. North American Journal of Fisheries Management, 40(4):910–921, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bopp:2023:BAF

[BHB^{+}23] Justin Bopp, Seth J. Herbst, Travis O. Brenden, Kevin Wehrly, and Jan-Michael Hessenauer. Biotic and abiotic factors that influence walleye recruitment in stocked lakes in Michigan. North American Journal of Fisheries Management, 43(6):1673–1686, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowersox:2023:GCW

[BHCC23] Brett J. Bowersox, John S. Hargrove, Timothy Copeland, and Matthew R. Campbell. The genetic composition of wild steelhead based on spatial proximity to a hatchery. North American Journal of Fisheries Management, 43(2):431–450, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blanchard:2023:PDD

[BHS^{+}23] Monica R. Blanchard, Julianne E. Harris, Joseph J. Skalicky, Gregory S. Silver, and Jeffrey C. Jolley. Patterns in distribution and density of larval lampreys in the main-stem Columbia River, Washington–Oregon. North American Journal of Fisheries Management, 43(6):1458–1474, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Briggs:2020:TER

[BHT^{+}20] Andrew S. Briggs, Jan-Michael Hessenauer, Michael V. Thomas, Brad E. Utrup, and Todd C. Wills. Trends and effects of a recreational lake sturgeon fishery in the St. Clair system. North American Journal of Fisheries Management, 40(3):752–761, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bottom:2021:ADH

[BHT^{+}21] Daniel L. Bottom, Susan A. Hinton, David J. Teel, G. Curtis Roegner, Lyndal L. Johnson, and Benjamin P. Sandford. The abundance and distribution of hatchery and naturally produced Chinook salmon in Columbia River Estuary nearshore habitat. North American Journal of Fisheries Management, 41(5):1549– 1571, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boehm:2022:SSO

[BIE^{+}22] Hadley I. A. Boehm, Daniel A. Isermann, Mark J. Ermer, Lawrence D. Eslinger, Gretchen J. A. Hansen, and Dale E. Logsdon. Special section overview: Effects of ecosystem change on North American percid populations. North American Journal of Fisheries Management, 42(3):477–483, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2022:EWM

[BKE22] Brian G. Blackwell, Todd M. Kaufman, and Mark J. Ermer. Evaluation of walleye minimum length limits with changing water regimes in the Great Plains. North American Journal of Fisheries Management, 42(3):585–596, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Budnik:2021:ISP

[BKM21] Richard R. Budnik, Jahn L. Kallis, and Elizabeth A. Marschall. Influence of stocking practices, predation, and reservoir characteristics on the growth and recruitment of saugeye. North American Journal of Fisheries Management, 41(4):1159–1169, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bieber:2023:CPB

[BLS23] John F. Bieber, Michael J. Louison, and Cory D. Suski. Capture is predicted by behavior and size, not metabolism, in muskellunge. North American Journal of Fisheries Management, 43(1):231– 243, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beard:2020:RFI

[BM20] Zachary S. Beard and Ryan D. Mann. Retracted: Factors influencing angler satisfaction and catch rates in Arizona stream put-and-take trout fisheries. North American Journal of Fisheries Management, 40(2):521, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brewer:2022:CTM

[BMAM22] Shannon K. Brewer, Robert Mollenhauer, Justin Alexander, and Desiree Moore. Critical thermal maximum values for stream fishes, including distinct populations of smallmouth bass. North American Journal of Fisheries Management, 42(2):352–360, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodine:2021:RYF

[BOB^{+}21] Kristopher A. Bodine, Richard A. Ott, Daniel L. Bennett, Jacob D. Norman, and J. Warren Schlechte. Round 2: a 4-year follow-up evaluation of a flathead catfish population exposed to hand fishing. North American Journal of Fisheries Management, 41(S1):S255–S258, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barnett:2020:MES

[BOHA20] Zanethia C. Barnett, Clifford A. Ochs, Jason D. Hoeksema, and Susan B. Adams. Multipass electrofishing sampling efficiency for stream crayfish population estimates. North American Journal of Fisheries Management, 40(4):840–851, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Booth:2020:PPD

[Boo20] Michael T. Booth. Patterns and potential drivers of steelhead smolt migration in Southern California. North American Journal of Fisheries Management, 40(4):1032–1050, August 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Babin:2021:ASU

[BPL^{+}21] Amanda B. Babin, Stephan Peake, Tommi Linnansaari, R. Allen Curry, Mouhamed Ndong, Katy Haralampides, and Ross Jones. Atlantic salmon upstream migration delay in a large hydropower reservoir. North American Journal of Fisheries Management, 41 (1):158–175, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burns:2021:SJC

[BPT21] Christopher W. Burns, Ronald A. Ptolemy, and Peter J. Tschaplinski. A synthesis of juvenile coastal cutthroat trout emigration studies: Considerations for researchers and resource managers. North American Journal of Fisheries Management, 41(5):1522– 1537, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brauer:2020:MDN

[BQR^{+}20] Tucker A. Brauer, Michael C. Quist, Darren T. Rhea, Troy W. Laughlin, and Erik Waring. Movement dynamics of nonnative burbot in the Upper Green River System and implications for management. North American Journal of Fisheries Management, 40(5):1161–1173, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bushon:2021:RPV

[BR21] Amanda M. Bushon and Jacob M. Rash. Retention of postocular visible implant elastomer in two sizes of adult brown trout and rainbow trout. North American Journal of Fisheries Management, 41(5):1384–1389, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2020:VAI

[BS20] David L. Buckmeier and Nathan G. Smith. Validation of annuli and identification of discontinuities in sagittal otoliths of juvenile alligator gar. North American Journal of Fisheries Management, 40(3):607–612, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baker:2021:PDS

[BS21] Sarah M. Baker and Steven M. Sammons. Persistence and dispersal of stocked rainbow trout in an Alabama tailwater. North American Journal of Fisheries Management, 41(3):820–833, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brandl:2021:EPC

[BSC^{+}21] Scott Brandl, Brian Schreier, J. Louise Conrad, Bernie May, and Melinda Baerwald. Enumerating predation on Chinook salmon, delta smelt, and other San Francisco Estuary fishes using genetics. North American Journal of Fisheries Management, 41(4):1053– 1065, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2020:URR

[BSKM20] Brian G. Blackwell, Bradley J. Smith, Todd M. Kaufman, and Tyrel S. Moos. Use of a restrictive regulation to manage walleyes in a new glacial lake in South Dakota. North American Journal of Fisheries Management, 40(5):1202–1215, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brunson:2024:EMB

[BSKSK24] Jeff F. Brunson, Kimberly A. Sitta, Peter R. Kingsley-Smith, and Michael R. Kendrick. Evidence for a male bias in Atlantic blue crab pot-based sampling. North American Journal of Fisheries Management, 44(1):101–112, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brino:2023:UGF

[BSR^{+}23] Erin E. Brino, David A. Schumann, Calvin R. Rezac, Robbie J. Ellwanger, and Todd W. Osmundson. Unlocking the genomes of formalin-fixed freshwater fish specimens: an assessment of factors influencing DNA extraction quantity and quality. North American Journal of Fisheries Management, 43(6):1596–1609, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Becher:2021:APR

[BTZ^{+}21] Cory Becher, Stephen M. Tyszko, Richard D. Zweifel, Joseph D. Conroy, Jeremy J. Pritt, and Stuart A. Ludsin. Alternative prey reduces largemouth bass predation mortality on newly stocked channel catfish fingerlings. North American Journal of Fisheries Management, 41(5):1322–1333, October 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bohling:2021:DSS

[BVP^{+}21] Justin Bohling, Jennifer Von Bargen, Matthew Piteo, Amelia Louden, Maureen Small, Thomas A. Delomas, and Ryan Kovach. Developing a standardized single nucleotide polymorphism panel for rangewide genetic monitoring of bull trout. North American Journal of Fisheries Management, 41(6):1920–1931, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buchanan:2022:USM

[BW22] Rebecca A. Buchanan and Steven L. Whitlock. Understanding salmon migration dynamics in a data-limited environment. North American Journal of Fisheries Management, 42(5):1111– 1133, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Branigan:2022:DMP

[BWM22] Philip R. Branigan, Emily F. Washburne, and Kevin A. Meyer. Does Moon phase affect indices of Kokanee size and abundance as characterized by gill-net catch? North American Journal of Fisheries Management, 42(3):787–794, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chong:2021:PCD

[CAZ^{+}21] Daniel R. Chong, Alice E. I. Abrams, Aaron J. Zolderdo, Michael J. Lawrence, Connor H. Reid, and Steven J. Cooke. Physiological consequences of different fishing tournament culling methods on largemouth bass. North American Journal of Fisheries Management, 41(1):26–34, February 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chiotti:2023:LSP

[CBB^{+}23] Justin A. Chiotti, James C. Boase, Andrew S. Briggs, Chris Davis, Richard Drouin, Darryl W. Hondorp, Lloyd Mohr, Edward F. Roseman, Michael V. Thomas, and Todd C. Wills. Lake sturgeon population trends in the st. clair–detroit river system, 2001–2019. North American Journal of Fisheries Management, 43(4):1066–1080, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carim:2020:EDS

[CBC^{+}20] K. J. Carim, N. J. Bean, J. M. Connor, W. P. Baker, M. Jaeger, M. P. Ruggles, K. S. McKelvey, T. W. Franklin, M. K. Young, and M. K. Schwartz. Environmental DNA sampling informs fish eradication efforts: Case studies and lessons learned. North American Journal of Fisheries Management, 40(2):488–508, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caldwell:2023:DRS

[CBKH23] John M. Caldwell, Summer M. Burdick, Jacob R. Krause, and Alta C. Harris. Does release size into net-pens affect survival of captively reared juvenile endangered suckers in Upper Klamath Lake? North American Journal of Fisheries Management, 43(5): 1322–1336, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Camhi:2021:ADA

[CBLW21] Merry Camhi, Michael Bednarski, Jake LaBelle, and John Waldman. Abundance and distribution of American eel in a heavily dammed urban river. North American Journal of Fisheries Management, 41(4):1131–1140, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caves:2021:DDC

[CBS21] Stradder Caves, Jessica R. Baumann, and Daniel S. Stich. Density-dependent changes in grass carp growth and mortality in long-term aquatic plant management. North American Journal of Fisheries Management, 41(2):355–365, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coulter:2022:SEA

[CCL^{+}22] Alison A. Coulter, David P. Coulter, Ellen Lark, Mark W. Fritts, Cameron Davis, David Glover, and Reuben Goforth. Sublethal effects of acoustic dummy tag implantation and external tags in silver carp. North American Journal of Fisheries Management, 42 (6):1502–1512, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clancy:2022:RTG

[CDB22] Niall G. Clancy, James L. Dunnigan, and Phaedra Budy. Relationship of trout growth to frequent electrofishing and diet collection in a headwater stream. North American Journal of Fisheries Management, 42(1):109–114, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coykendall:2022:IAE

[CDBC22] D. Katharine Coykendall, Thomas A. Delomas, Matthew Belnap, and Matthew R. Campbell. Improving abundance estimates of spring–summer Snake River Chinook salmon for fisheries management. North American Journal of Fisheries Management, 42 (6):1454–1464, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chelgren:2023:CEK

[CDG^{+}23] Nathan D. Chelgren, Jason B. Dunham, Stephanie L. Gunckel, David P. Hockman-Wert, and Chris S. Allen. Combining expert knowledge of a threatened trout distribution with sparse occupancy data for climate-related projection. North American Journal of Fisheries Management, 43(3):839–858, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2023:EAL

[CEB^{+}23] Jr. Richard D. Clark, Mark P. Ebener, James R. Bence, Matthew S. Kornis, Charles R. Bronte, Theodore J. Treska, Jory L. Jonas, Charles P. Madenjian, and Iyob W. Tsehaye. Estimates of adult lake trout mortality from coded wire tags in a population with developing natural reproduction in southern Lake Michigan. North American Journal of Fisheries Management, 43(4):1035–1051, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coleman:2024:ESR

[CEC^{+}24] Tyler Steven Coleman, Robert W. Eckelbecker, Andrew K. Carlson, Dennis R. DeVries, Russell A. Wright, Benjamin A. Staton, Stephen W. Parker, Collin R. Chittam, Richard G. Lovell, and Matthew J. Catalano. Evaluation of shoreline rotenone application to control largemouth bass recruitment in small impoundments. North American Journal of Fisheries Management, 44 (1):57–69, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Conley:2020:SEI

[CEH^{+}20] Keats R. Conley, Jonathan D. Ebel, John S. Hargrove, Wyatt Petersen, and Lytle P. Denny. In-stream egg incubators produce hatchery Chinook salmon with similarities to and differences from natural juveniles. North American Journal of Fisheries Management, 40(1):256–277, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colborne:2024:EIT

[CFB^{+}24] S. F. Colborne, M. D. Faust, T. O. Brenden, T. A. Hayden, J. M. Robinson, T. M. MacDougall, H. A. Cook, D. A. Isermann, D. J.

Dembkowski, M. Haffley, and C. S. Vandergoot. Estimating internal transmitter and external tag retention by walleye in the Laurentian Great Lakes over multiple years. North American Journal of Fisheries Management, 44(2):377–393, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Courtney:2020:ESA

[CFCN20] Kristin R. Courtney, Jeffrey A. Falke, M. Keith Cox, and Jeff Nichols. Energetic status of Alaskan Chinook salmon: Interpopulation comparisons and predictive modeling using bioelectrical impedance analysis. North American Journal of Fisheries Management, 40(1):209–224, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clemens:2023:CBP

[CFGZ23] Benjamin J. Clemens, Thomas A. Friesen, Stanley V. Gregory, and Courtney L. Zambory. The case for basinwide passage and habitat restoration for Pacific Lamprey in the Willamette River basin (Oregon, USA). North American Journal of Fisheries Management, 43(6):1567–1583, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Courtney:2022:CSS

[CFHS22] Michael B. Courtney, Austin J. Flanigan, Mary Hostetter, and Andrew C. Seitz. Characterizing sockeye salmon smolt interactions with a hydrokinetic turbine in the Kvichak River, Alaska. North American Journal of Fisheries Management, 42(4):1054– 1065, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chestnut-Faull:2021:UPD

[CFPSW21] Kristen L. Chestnut-Faull, Quinton E. Phelps, Dustin M. Smith, and David I. Wellman. Using population dynamics to model harvest regulation impacts to channel catfish in the Monongahela River, West Virginia. North American Journal of Fisheries Management, 41(S1):S259–S267, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Culver:2021:IPMb

[CGD^{+}21] Carolynn S. Culver, Samuel C. Ginther, Daniel Daft, Leigh T. Johnson, and Andrew J. Brooks. An integrated pest management tactic for quagga mussels: Site-specific application of fish biological control agents. North American Journal of Fisheries Management, 41(2):329–343, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlson:2022:RSO

[CH22] Andrew K. Carlson and Mark V. Hoyer. Redear sunfish occurrence, abundance, growth, and size structure as related to abiotic and biotic factors in Florida lakes. North American Journal of Fisheries Management, 42(3):775–786, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:2023:GMS

[CHB^{+}23] Erin E. Collins, Jon E. Hess, Shawn Bechtol, Nicolas Romero, Shawn R. Narum, and Joseph S. Zendt. Genetic monitoring of steelhead in the Klickitat River to estimate productivity, straying, and migration timing. North American Journal of Fisheries Management, 43(4):1000–1016, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cheng:2022:AIO

[CHJ^{+}22] Matthew L. H. Cheng, Scott G. Hinch, Francis Juanes, Stephen J. Healy, Andrew G. Lotto, Sydney J. Mapley, and Nathan B. Furey. Acoustic imaging observes predator–prey interactions between bull trout and migrating sockeye salmon smolts. North American Journal of Fisheries Management, 42(6):1494–1501, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cristan:2022:EEP

[CHL22] Erik T. Cristan, Adam G. Hansen, and Jesse M. Lepak. Effects of ethanol preservation on larval and fingerling walleye and gizzard shad body size. North American Journal of Fisheries Management, 42(4):874–881, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cates:2024:GCT

[CHO24] Lydia Cates, Stephen Hale, and Zachary Olsen. Gafftopsail catfish in Texas estuaries: Population trends and ecosystem implications. North American Journal of Fisheries Management, 44 (1):145–161, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clemens:2022:SMS

[CHS^{+}22] Benjamin J. Clemens, Julianne E. Harris, Steven J. Starcevich, Thomas M. Evans, Joseph J. Skalicky, Fraser Neave, and Ralph T. Lampman. Sampling methods and survey designs for larval lampreys. North American Journal of Fisheries Management, 42(2): 455–474, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carline:2021:HHM

[CJN^{+}21] Robert F. Carline, Mark F. Jackson, Mark A. Nale, Dhanushi A. Wijeyakulasuriya, and Xiaoyue Niu. Hooking and handling mortality of trout captured in the Bald Eagle Creek Trout Tournament, Pennsylvania. North American Journal of Fisheries Management, 41(5):1465–1472, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Culver:2021:IPMa

[CJP^{+}21] Carolynn S. Culver, Leigh T. Johnson, Henry M. Page, Jenifer E. Dugan, and Christen A. Santschi. Integrated pest management for fouling organisms on boat hulls. North American Journal of Fisheries Management, 41(2):301–328, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cope:2021:GSD

[CKB^{+}21] W. Robert Cope, Thomas J. Kwak, Tyler R. Black, Krishna Pacifici, Sheila C. Harris, Caitlin M. Miller, Morgan E. Raley, and Eric M. Hallerman. Genetic structure and diversity of the endemic Carolina madtom and conservation implications. North American Journal of Fisheries Management, 41(S1):S27–S41, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connerton:2022:OPS

[CLBL22] Michael J. Connerton, Jana R. Lantry, Charles R. Bronte, and Steve R. Lapan. Origin, postrelease survival, and imprinting of pen-acclimated and direct-stocked Chinook salmon in Lake Ontario. North American Journal of Fisheries Management, 42(3): 713–740, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Claiborne:2020:EMB

[CLM20] Andrew M. Claiborne, James P. Losee, and Jessica A. Miller. Estimating migratory behavior and age for anadromous coastal cutthroat trout in South Puget Sound: Evaluation of approaches based on fish scales versus otoliths. North American Journal of Fisheries Management, 40(5):1313–1323, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carter:2022:ECT

[CLRL22] David W. Carter, Sabrina Lovell, David Records, and Christopher Liese. The effect of changes in trip costs and gag regulations on recreational fishing demand in the Gulf of Mexico. North American Journal of Fisheries Management, 42(6):1465–1476, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catterson:2020:IBM

[CLSM20] Matthew R. Catterson, David C. Love, Trent M. Sutton, and Megan V. McPhee. Interactions between marine growth and life history diversity of steelhead from the Situk River, Alaska. North American Journal of Fisheries Management, 40(1):242– 255, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chapagain:2021:VBB

[CLTJ21] Binod P. Chapagain, James M. Long, Andrew T. Taylor, and Omkar Joshi. Variation in black bass angler characteristics by stream size and accessibility in Oklahoma’s Ozark highland streams. North American Journal of Fisheries Management, 41 (3):585–599, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chiaramonte:2022:EBL

[CM22] Luciano V. Chiaramonte and Kevin A. Meyer. Effect of bag limits on angler harvest, catch rates, and satisfaction at put-and-take community pond fisheries. North American Journal of Fisheries Management, 42(6):1550–1560, December 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chiaramonte:2020:EPD

[CMBR20] Luciano V. Chiaramonte, Kevin A. Meyer, Philip R. Branigan, and James B. Reynolds. Effect of pulsed DC frequency on capture efficiency and spinal injury of trout in small streams. North American Journal of Fisheries Management, 40(3):691–699, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carson:2020:RES

[COT20] Evan W. Carson, Megan J. Osborne, and Thomas F. Turner. Relationship of effective size to hatchery supplementation and habitat connectivity in a simulated population of Rio Grande silvery minnow. North American Journal of Fisheries Management, 40 (4):922–938, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chen:2021:ISS

[CP21a] Yushun Chen and Quinton Phelps. Introduction to a special section: 4th Mississippi–Yangtze River Basins Symposium. North American Journal of Fisheries Management, 41(4):889–890, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clarkson:2021:CFH

[CP21b] Emma Clarkson and Jennifer Beseres Pollack. Characterizing fish–habitat associations through paired fisheries-independent and in situ habitat assessments. North American Journal of Fisheries Management, 41(1):49–63, February 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cowie:2023:APP

[CR23] Emily D. Cowie and Mark S. Ridgway. Angling party persistence and visitation in a recreational lake trout fishery: Relative influence of travel distance and fuel costs. North American Journal of Fisheries Management, 43(6):1725–1734, December 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chhor:2022:TRC

[CRH^{+}22] Auston D. Chhor, Jessica L. Reid, Peter E. Holder, Liane B. Nowell, Jacob W. Brownscombe, Andy J. Danylchuk, and Steven J. Cooke. Temporary retention in cold water reduces postrelease behavioral impairment in angled rainbow trout. North American Journal of Fisheries Management, 42(1):140–150, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coffill-Rivera:2024:ULL

[CRNRFL24] Manuel E. Coffill-Rivera, J. Wesley Neal, Grisel Rodríguez-Ferrer, and Craig G. Lilyestrom. Using lessons learned from a multidecadal intercept survey of Puerto Rico spear fishers to improve data collection in the U.S. Caribbean. North American Journal of Fisheries Management, 44(2):507–519, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlson:2022:DAS

[CSLW22] Tanner L. Carlson, Benjamin J. Schall, David O. Lucchesi, and Jeff S. Wesner. Development and application of a size structure index for blue sucker. North American Journal of Fisheries Management, 42(1):200–206, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cheng:2021:STD

[CSQ^{+}21] Fei Cheng, Bjorn Schmidt, Jiao Qin, Songuang Xie, and Wei Li. Short-term distribution patterns of young-of-the-year fish assemblages between the main stem and affiliated lakes in the middle reaches of the Yangtze River, China. North American Journal of Fisheries Management, 41(4):904–915, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Camacho:2023:ICR

[CSWP23a] C. A. Camacho, C. J. Sullivan, M. J. Weber, and C. L. Pierce. Invasive carp reproduction phenology in tributaries of the Upper Mississippi River. North American Journal of Fisheries Management, 43(1):61–80, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Camacho:2023:SUM

[CSWP23b] Carlos A. Camacho, Christopher J. Sullivan, Michael J. Weber, and Clay L. Pierce. Suitability of an Upper Mississippi River tributary for invasive carp reproduction. North American Journal of Fisheries Management, 43(1):12–24, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Christiansen:2022:PIE

[CTJS^{+}22] Heather M. Christiansen, Amanda J. Tyler-Jedlund, Theodore S. Switzer, Sean F. Keenan, and Scott D. Stahl. Potential implications of excluding live-bottom habitat from a bottom-trawl survey in the Eastern Gulf of Mexico. North American Journal of Fisheries Management, 42(4):893–905, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caumartin:2020:RMA

[CVD^{+}20] Jean Caumartin, Richard Verdon, Denis Desrochers, Alexandre Guindon, François Lafleur, and Frédéric Guay. Response of outmigrating adult American shad guided by spill and ultrasound as they approach Rivière-des–Prairies Hydroelectric Generating Station (Québec, Canada). North American Journal of Fisheries Management, 40(3):773–788, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chin:2021:RAE

[CWB^{+}21] Sam Chew Chin, John Waldman, Mike Bednarski, Merry Camhi, Jake LaBelle, and S. Elizabeth Alter. Relating American eel abundance to environmental DNA concentration in the Bronx River. North American Journal of Fisheries Management, 41(4):1141– 1150, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2020:OBA

[DAS20] Daniel J. Daugherty, Allen H. Andrews, and Nathan G. Smith. Otolith-based age estimates of alligator gar assessed using bomb radiocarbon dating to greater than 60 years. North American Journal of Fisheries Management, 40(3):613–621, June 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dobos:2020:ULH

[DBCS20] Marika E. Dobos, Brett J. Bowersox, Timothy Copeland, and Eric J. Stark. Understanding life history diversity of a wild steelhead population and managing for resiliency. North American Journal of Fisheries Management, 40(5):1087–1099, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dattilo:2021:FDI

[DBS21] J. Dattilo, S. K. Brewer, and D. E. Shoup. Flow dynamics influence fish recruitment in hydrologically connected river–reservoir landscapes. North American Journal of Fisheries Management, 41(6):1752–1763, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Delomas:2022:GIG

[DC22] Thomas A. Delomas and Matthew R. Campbell. Grandparent inference from genetic data: The potential for parentage-based tagging programs to identify offspring of hatchery strays. North American Journal of Fisheries Management, 42(1):85–95, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dutterer:2020:ERF

[DDT^{+}20] Andrew C. Dutterer, Jason R. Dotson, Brandon C. Thompson, Christopher J. Paxton, and William F. Pouder. Estimating recreational fishing effort using autonomous cameras at boat ramps versus creel surveys. North American Journal of Fisheries Management, 40(6):1367–1378, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Doll:2020:PES

[DEO20] Jason C. Doll, Luke Etchison, and Dylan Owensby. Population estimate of the state and federally threatened spotfin chub using underwater observations. North American Journal of Fisheries Management, 40(2):342–353, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauwalter:2022:DHU

[DGJ^{+}22] Daniel C. Dauwalter, Tim Gatewood, Zachary J. Jackson, Jean Barney, and Zachary S. Beard. Digital hydrography underestimates stream length and leads to underestimates of trout population size. North American Journal of Fisheries Management, 42 (4):994–1002, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Delomas:2021:NEC

[DH21] Thomas A. Delomas and Jon E. Hess. A new estimator to correct for bias from tag rate expansion on natural-origin fish attributes in mixed-stock analysis using parentage-based tagging. North American Journal of Fisheries Management, 41(2):421–433, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dean:2021:ETC

[DHWP21] Quintin J. Dean, Martin J. Hamel, Jake P. Werner, and Mark A. Pegg. Efficacy and temporal capture patterns of bank poles in the Kansas River: a novel sampling tool for catfish. North American Journal of Fisheries Management, 41(S1):S379–S387, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dembkowski:2020:EEP

[DKEI20] Daniel J. Dembkowski, Janice A. Kerns, Emma G. Easterly, and Daniel A. Isermann. Electrofishing encounter probability, survival, and dispersal of stocked age-0 muskellunge in Wisconsin lakes. North American Journal of Fisheries Management, 40(2): 383–393, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dobos:2023:AAL

[DMBC23] Marika E. Dobos, Joshua L. McCormick, Brett J. Bowersox, and Timothy Copeland. Applications of age–length keys for assigning ages to juvenile steelhead. North American Journal of Fisheries Management, 43(3):758–771, June 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dembkowski:2022:LUE

[DNW^{+}22] Daniel J. Dembkowski, Daniel T. Nelson, Melissa R. Wuellner, Brian G. Blackwell, David O. Lucchesi, and David W. Willis. Limited utility of early life-history indices for predicting yellow perch year-class strength in Eastern South Dakota glacial lakes. North American Journal of Fisheries Management, 42(2):343– 351, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dick:2020:ASS

[DPR^{+}20] M. Dick, D. A. Patterson, K. A. Robinson, E. J. Eliason, S. G. Hinch, and S. J. Cooke. Adult sockeye salmon gastrically tagged near spawning grounds exhibit lower survival rates throughout the spawning period than externally tagged conspecifics. North American Journal of Fisheries Management, 40(4):939–951, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Donohoe:2021:SLH

[DRPW21] Christopher J. Donohoe, David E. Rundio, Devon E. Pearse, and Thomas H. Williams. Straying and life history of adult steelhead in a small California coastal stream revealed by otolith natural tags and genetic stock identification. North American Journal of Fisheries Management, 41(3):711–723, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dennis:2020:CCI

[DS20] Clark E. Dennis III and Peter W. Sorensen. Common carp are initially repelled by a broadband outboard motor sound in a lock chamber but habituate rapidly. North American Journal of Fisheries Management, 40(6):1499–1509, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Durhack:2024:CTR

[DSW^{+}24] Travis C. Durhack, Holly A. Simpson, Douglas A. Watkinson, Mark A. Pegg, and Eva C. Enders. A comparison of tag retention and mortality from two tagging methods for internal tag placement in channel catfish. North American Journal of Fisheries Management, 44(2):428–437, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Devine:2020:PEM

[DTK20] Thomas E. Devine, Sara J. Tripp, and Nicholas W. Kramer. Paddlefish exploitation and movements within the Mississippi River Basin. North American Journal of Fisheries Management, 40(2): 406–414, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

deVlaming:2020:FIC

[dVB20a] Catherine M. de Vlaming and Kevin R. Bestgen. Factors influencing capture and retention of razorback sucker in light traps. North American Journal of Fisheries Management, 40(5):1055– 1076, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

deVlaming:2020:LTE

[dVB20b] Catherine M. de Vlaming and Kevin R. Bestgen. Light trap effectiveness to capture razorback sucker larvae in a large floodplain wetland. North American Journal of Fisheries Management, 40 (6):1435–1442, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Doll:2021:IBM

[DWGR21] Jason C. Doll, Chris J. Wood, David W. Goodfred, and Jacob M. Rash. Incorporating batch mark–recapture data into an integrated population model of brown trout. North American Journal of Fisheries Management, 41(5):1390–1407, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DAmico:2021:RPI

[DWSM21] Timothy W. D’Amico, Dana L. Winkelman, Tyler R. Swarr, and Christopher A. Myrick. Retention of passive integrated transponder tags in a small-bodied catfish. North American Journal of Fisheries Management, 41(1):187–195, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eckelbecker:2022:ITL

[ECC22] Robert W. Eckelbecker, Tyler Steven Coleman, and Matthew J. Catalano. Incorporating time-lapse digital cameras into creel surveys at three Alabama reservoirs. North American Journal of Fisheries Management, 42(5):1349–1358, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elbassiouny:2023:GSS

[EFK^{+}23] Ahmed A. Elbassiouny, João Pedro Fontenelle, Patrick M. Kočovský, Nicholas E. Mandrak, and Nathan R. Lovejoy. Genetic structure of the silver chub indicates distinctiveness of Lake Erie population. North American Journal of Fisheries Management, 43(5):1180–1189, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ebel:2022:FBI

[ELCM22] Jonathan D. Ebel, Donald A. Larsen, Keats R. Conley, and Mollie A. Middleton. A fish out of basin: Increased stress physiology and reduced performance of Salmon River hatchery Chinook salmon. North American Journal of Fisheries Management, 42 (3):741–757, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eggleston:2020:IFC

[EMRK20] Michael R. Eggleston, Scott W. Milne, Maxwell Ramsay, and Kurt P. Kowalski. Improved fish counting method accurately quantifies high-density fish movement in dual-frequency identification sonar data files from a coastal wetland environment. North American Journal of Fisheries Management, 40(4):883–892, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eggleton:2021:BBG

[EP21] Michael A. Eggleton and Clint R. Peacock. Black bass growth patterns in relation to hydrology in the Arkansas River, Arkansas. North American Journal of Fisheries Management, 41(4):955– 966, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evans:2022:APJ

[EPB^{+}22] Allen F. Evans, Quinn Payton, Nathan Banet, Bradley M. Cramer, Caylen Kelsey, and David A. Hewitt. Avian predation on juvenile and adult Lost River and shortnose suckers: an updated multi-predator species evaluation. North American Journal of Fisheries Management, 42(6):1561–1574, December 2022. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Euclide:2024:WGF

[EPD^{+}24] Peter T. Euclide, Christian Perry, Steven B. Donabauer, Angie Grier, and Tomas O. Höök. What’s good for fish is good for fishes: Interspecific consistency of growth, condition, and abundance among glacial lakes. North American Journal of Fisheries Management, 44(2):319–334, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ehnert-Russo:2022:ANF

[ERCB22] Samantha L. Ehnert-Russo, Claire E. Crowley, and Katharine L. Becker. Assessing Northeast Florida offshore blue crab fishing effort within Florida’s state waters. North American Journal of Fisheries Management, 42(3):668–683, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elliott:2021:CAI

[ERT21] Connor W. Elliott, Alexander J. Row, and Bruce L. Tufts. Comparison of angler interventions for smallmouth bass experiencing barotrauma in catch-and-release angling tournaments. North American Journal of Fisheries Management, 41(3):834–843, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Erickson:2021:HEG

[ESCI21] Keith A. Erickson, Peter C. Sakaris, Hannah Conner, and Elise R. Irwin. Hydrologic effects on growth and hatching success of age-0 channel catfish in the Tallapoosa River Basin: Implications for management in regulated systems. North American Journal of Fisheries Management, 41(S1):S118–S132, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Erickson:2023:PSM

[ESF23] Tyler A. Erickson, George J. Schisler, and Eric R. Fetherman. Poststocking survival and myxospore evaluation of whirling disease-resistant rainbow trout strains. North American Journal of Fisheries Management, 43(2):586–599, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fisch:2020:DQD

[FB20] Nicholas C. Fisch and James R. Bence. Data quality, data quantity, and its effect on an applied stock assessment of cisco in Thunder Bay, Ontario. North American Journal of Fisheries Management, 40(2):368–382, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fuchs:2021:ODP

[FCMT21] Nathaniel T. Fuchs, Christopher C. Caudill, Andrew R. Murdoch, and Benjamin L. Truscott. Overwintering distribution and postspawn survival of steelhead in the Upper Columbia River Basin. North American Journal of Fisheries Management, 41(3): 757–774, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Franssen:2021:FCH

[FDG^{+}21] Nathan R. Franssen, Scott L. Durst, Eliza I. Gilbert, William K. Knight, and Manuel Ulibarri. Flow conditioning of hatcheryreared razorback sucker increases apparent survival in the wild. North American Journal of Fisheries Management, 41(2):545– 555, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Felts:2020:ERE

[FFG20] Eli A. Felts, Mark J. Fincel, and Brian D. S. Graeb. Effects of reservoir elevation and spatial synchrony on walleye recruitment in Lake Oahe, Missouri River. North American Journal of Fisheries Management, 40(5):1133–1145, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fredricks:2021:CCI

[FHA^{+}21] Kim T. Fredricks, Terrance D. Hubert, Jon J. Amberg, Aaron R. Cupp, and Verdel K. Dawson. Chemical controls for an integrated pest management program. North American Journal of Fisheries Management, 41(2):289–300, April 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Farrell:2024:ERS

[FHB^{+}24] Collin J. Farrell, Adam G. Hansen, Mandi M. Brandt, Christopher A. Myrick, and Brett M. Johnson. An evaluation of the relative size, body condition, and survival of triploid walleye in the wild. North American Journal of Fisheries Management, 44 (1):172–188, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frechette:2023:MAS

[FHK23] Danielle M. Frechette, James P. Hawkes, and John F. Kocik. Managing for Atlantic salmon smolt run timing variability in a changing climate. North American Journal of Fisheries Management, 43(2):517–538, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fitzsimons:2021:ODC

[FHR21] John D. Fitzsimons, Dale C. Honeyfield, and Scott Rush. Ontogenetic dietary changes and thiamine status of Lake Ontario Chinook salmon, 2005–2006. North American Journal of Fisheries Management, 41(5):1499–1513, October 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Farrell:2022:CMA

[FJH^{+}22] Collin J. Farrell, Brett M. Johnson, Adam G. Hansen, Christopher A. Myrick, Eric C. Anderson, Thomas A. Delomas, Andrea D. Schreier, and Joel P. Van Eenennaam. Cytological and molecular approaches for ploidy determination: Results from a wild walleye population. North American Journal of Fisheries Management, 42(4):849–856, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flannery:2021:MVY

[FLW21] Blair G. Flannery, Randal G. Loges, and John K. Wenburg. Microsatellite variation in Yukon River Coho salmon: Population structure and application to mixed-stock analysis. North American Journal of Fisheries Management, 41(5):1408–1422, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Faulkner:2020:ASS

[FM20] April V. Faulkner and Suzanne L. Maxwell. Adult sockeye salmon assessment in a tidal, turbid river: a comparison of sonar and test fishing methods. North American Journal of Fisheries Management, 40(4):852–864, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fraley:2020:IMT

[FMM20] Kevin M. Fraley, Peter A. McHugh, and Angus R. McIntosh. Influence of maternally-transferred nitrogen and carbon on stable isotope ratios in juvenile Chinook salmon. North American Journal of Fisheries Management, 40(1):175–181, February 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fischer:2022:UST

[FNB^{+}22] Jason L. Fischer, Lucas R. Nathan, John Buszkiewicz, Julia Colm, D. Andrew R. Drake, Mark R. DuFour, Patrick M. Kočovský, David Marson, Eric R. B. Smyth, Ryan Young, and Kelly F. Robinson. Using surrogate taxa to inform response methods for invasive grass carp in the Laurentian Great Lakes. North American Journal of Fisheries Management, 42(1):151– 163, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fabrizio:2021:IBC

[FNT21] Mary C. Fabrizio, Vaskar Nepal, and Troy D. Tuckey. Invasive blue catfish in the Chesapeake Bay region: a case study of competing management objectives. North American Journal of Fisheries Management, 41(S1):S156–S166, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Finlay:2020:DDB

[FPR20] Ross Finlay, Russell Poole, and Thomas E. Reed. Development of a double-breakaway pass-through PIT tag antenna system for flood-prone rivers. North American Journal of Fisheries Management, 40(4):952–958, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fritts:2023:PMS

[FTLR23] Anthony L. Fritts, Gabriel M. Temple, Cade Lillquist, and Dan Rawding. Postrelease mortality of spring Chinook salmon from a mark-selective recreational fishery in the Yakima River, Washington. North American Journal of Fisheries Management, 43(6): 1702–1712, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gilham:2021:PSW

[GBJ21] Andrew T. Gilham, Michael L. Brown, and George R. Jordan. Proposed standard weight (W_{s}) equations for Arctic grayling. North American Journal of Fisheries Management, 41(3):739– 745, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gostiaux:2022:RBA

[GBJ^{+}22] Jason C. Gostiaux, Hadley I. A. Boehm, Nathan J. Jaksha, Daniel J. Dembkowski, Joseph M. Hennessy, and Daniel A. Isermann. Recruitment bottlenecks for age-0 walleye in northern Wisconsin lakes. North American Journal of Fisheries Management, 42(3):507–522, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Griffeth:2023:STF

[GBM23] Steven P. Griffeth, Krista D. Baker, and Darrell R. J. Mullowney. Short-term fishery gains mask long-term resource pains: Spatial fisheries management changes promote hyperstable CPUE in Labrador snow crab Chionoecetes opilio during a period of heavy exploitation. North American Journal of Fisheries Management, 43(5):1359–1371, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gulak:2021:LST

[GC21] Simon J. B. Gulak and John K. Carlson. Less soak time saves those upon the line: Capture times and hooking mortality of sharks caught on bottom longlines. North American Journal of Fisheries Management, 41(3):791–808, June 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grooms:2022:RBW

[GDS^{+}22] Bennett Grooms, Ashley A. Dayer, Marc J. Stern, Jessica Barnes, and Jonathan D. Rutter. The relationship between wildlife recreationists’ familiarity with state fish and wildlife agencies, recreation identity, and future financial contributions. North American Journal of Fisheries Management, 42(6):1595–1608, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gephard:2023:TPA

[Gep23] Stephen Gephard. Transplantation of prespawn adult Sea Lampreys as an important restoration strategy. North American Journal of Fisheries Management, 43(6):1584–1595, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geist:2022:CGD

[GET22] Jeremy A. Geist, Diana Ethaiya, and Scott D. Tiegs. Coupling gear decontamination trials and angler surveys to minimize spread of invasive New Zealand mud snails Potamopyrgus antipodarum. North American Journal of Fisheries Management, 42 (4):1081–1093, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grausgruber:2021:IIA

[GGW21] Stephen J. Grausgruber, Emily E. Grausgruber, and Michael J. Weber. Influence of ice angler culling practices on bluegill physiological stress responses and mortality. North American Journal of Fisheries Management, 41(6):1703–1713, December 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gregory:2022:ART

[GGWV22] Jim S. Gregory, Bart L. Gamett, Caselle L. Wood, and Robert W. Van Kirk. Abundance, restoration, and transport of unanchored large wood in a small river in Central Idaho. North American Journal of Fisheries Management, 42(4):1041–1053, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Glade:2023:DPN

[GHA^{+}23] Kamden C. Glade, Brian R. Herwig, Tyler D. Ahrenstorff, Jeffrey R. Reed, and Andrew W. Hafs. Diet patterns and niche overlap of muskellunge and co-occurring piscivores in Minnesota lakes. North American Journal of Fisheries Management, 43(3): 656–676, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gaikowski:2021:ISS

[GK21] Mark Gaikowski and Patrick M. Kočovský. Introduction to a special section: Integrated pest management — extending a terrestrial paradigm to aquatic environments. North American Journal of Fisheries Management, 41(2):261–263, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gatch:2020:ESC

[GKRH20] Alexander J. Gatch, Scott T. Koenigbauer, Edward F. Roseman, and Tomas O. Höök. The effect of sediment cover and female characteristics on the hatching success of walleye. North American Journal of Fisheries Management, 40(1):293–302, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gatch:2021:ATT

[GKRH21] Alexander J. Gatch, Scott T. Koenigbauer, Edward F. Roseman, and Tomas O. Höök. Assessment of two techniques for remediation of Lacustrine rocky reef spawning habitat. North American Journal of Fisheries Management, 41(2):484–497, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Griswold:2022:DDT

[GKT^{+}22] Robert G. Griswold, Andre E. Kohler, Kurt A. Tardy, Kitty E. Griswold, and Doug Taki. Density dependence in three Snake River sockeye salmon nursery lakes in Central Idaho. North American Journal of Fisheries Management, 42(6):1477–1493, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gardner:2021:STV

[GLSF21] Jennifer L. Gardner, W. Christopher Long, Katherine M. Swiney, and Robert J. Foy. Spatial and temporal variation in annual versus biennial reproductive cycles in Eastern Bering Sea snow crab Chionoecetes opilio. North American Journal of Fisheries Management, 41(6):1878–1891, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goff:2022:EAT

[GRA^{+}22] Shelby Goff, Madeline E. Redd, J. Conner Almond, Reid Pearlman, and Peter J. Allen. Effects of acute temperature change associated with handling practices on metabolic rate stabilization in two centrarchids. North American Journal of Fisheries Management, 42(1):115–122, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallaway:2021:AAE

[GRM^{+}21] Benny J. Gallaway, Scott W. Raborn, Kyle A. McCain, R. Taylor Beyea, Susan Dufault, Will Heyman, Nathan F. Putman, and Jack Egerton. Absolute abundance estimates for red snapper, greater amberjack, and other federally managed fish on offshore petroleum platforms in the Gulf of Mexico. North American Journal of Fisheries Management, 41(6):1665–1690, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goldsmith:2023:IER

[GSML23] William M. Goldsmith, Andrew M. Scheld, Christopher McGuire, and Carl Lobue. Improving electronic reporting rates in the U.S. recreational fishery for Atlantic bluefin tuna. North American Journal of Fisheries Management, 43(4):893–907, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Golebie:2021:RIS

[GvRSS21] Elizabeth Golebie, Carena J. van Riper, Cory Suski, and Richard Stedman. Reducing invasive species transport among recreational anglers: The importance of values and risk perceptions. North American Journal of Fisheries Management, 41(6):1812–1825, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grausgruber:2020:BBE

[GW20] Emily Elise Grausgruber and Michael John Weber. Is bigger better? Evaluation of size-selective predation on age-0 walleye. North American Journal of Fisheries Management, 40(3):726–732, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grausgruber:2021:UBE

[GW21] Emily Elise Grausgruber and Michael John Weber. Using bioenergetics to estimate consumption of stocked age-0 walleye by a suite of piscivores. North American Journal of Fisheries Management, 41(2):383–398, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grausgruber:2021:ICW

[GWM21] Stephen J. Grausgruber, Michael J. Weber, and Joseph E. Morris. Importance of coarse woody habitat complexity to yellow perch egg skein deposition and survival. North American Journal of Fisheries Management, 41(6):1892–1902, December 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goode:2023:ERF

[GWMP23] Katherine Goode, Michael J. Weber, Aaron Matthews, and Clay L. Pierce. Evaluation of a random forest model to identify invasive carp eggs based on morphometric features. North American Journal of Fisheries Management, 43(1):46–60, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gravenhof:2024:EPR

[GWRF24] D. A. Gravenhof, M. R. Wuellner, E. A. Renner, and M. J. Fincel. Estimating predation rates of stocked juvenile Chinook salmon using novel acoustic predation transmitters. North American Journal of Fisheries Management, 44(2):438–448, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hajas:2024:ESP

[HA24] Wayne Hajas and Mike Atkins. Estimating seasonal Pacific geoduck clam Panopea generosa show-factors from long-term observation. North American Journal of Fisheries Management, 44 (1):162–171, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hassrick:2022:FAS

[HAP^{+}22] Jason L. Hassrick, Arnold J. Ammann, Russell W. Perry, Sara N. John, and Miles E. Daniels. Factors affecting spatiotemporal variation in survival of endangered winter-run Chinook salmon outmigrating from the Sacramento River. North American Journal of Fisheries Management, 42(2):375–395, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2020:AAB

[Har20] Julianne E. Harris. Assessing accuracy and bias of protocols to estimate age of Pacific salmon using scales. North American Journal of Fisheries Management, 40(4):1007–1022, August 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holbrook:2022:AMC

[HBBS22] Beth V. Holbrook, Bethany J. Bethke, Marc D. Bacigalupi, and David F. Staples. Assessing Minnesota’s changing yellow perch populations using length-based metrics. North American Journal of Fisheries Management, 42(3):642–658, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayer:2023:GCR

[HBR^{+}23] Cari-Ann Hayer, Michael A. Bayless, Catherine A. Richter, Amy E. George, and Duane C. Chapman. Grass carp reproduction in small tributaries of Harry S. Truman Reservoir, Missouri: Implications for establishment in novel habitats. North American Journal of Fisheries Management, 43(1):154–163, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hungerford:2021:RCC

[HBT^{+}21] Thomas J. Hungerford, Kristopher A. Bodine, John E. Tibbs, Randall A. Myers, David Prangnell, Daniel J. Daugherty, and J. Warren Schlechte. Relative catchability of channel catfish and blue catfish × channel catfish hybrids by anglers in put-and-take urban fisheries. North American Journal of Fisheries Management, 41(S1):S293–S297, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hummel:2022:EJL

[HBTE22] Stephanie Longstaff Hummel, Amber Bellamy, William A. Tucker, and Orey T. Eckes. Exposure of juvenile lake sturgeon to contaminants of emerging concern (CECs), including polybrominated diphenyl ethers (PBDEs): Location differences and effects on thyroid hormones. North American Journal of Fisheries Management, 42(1):123–139, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2021:MSU

[HBV^{+}21] Cleyo Harris, Travis O. Brenden, Christopher S. Vandergoot, Matthew D. Faust, Seth J. Herbst, John T. Buszkiewicz, Lucas R. Nathan, Jason L. Fischer, and Charles C. Krueger. Movement and space use of grass carp in the Sandusky River, Ohio: Implications for Lake Erie eradication efforts. North American Journal of Fisheries Management, 41(2):513–530, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansell:2021:SAL

[HCC^{+}21] Alexander C. Hansell, Tobey H. Curtis, John Carlson, Enric Cortés, Gavin Fay, and Steven X. Cadrin. Stock assessment of the lemon shark off the Southeast United States. North American Journal of Fisheries Management, 41(1):35–48, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hine:2022:ANE

[HCM^{+}22] Eric C. Hine, Robert E. Colombo, Scott J. Meiners, Cassi J. Moody-Carpenter, and Anabela Maia. An assessment of nighttime electrofishing in the Lower Wabash River. North American Journal of Fisheries Management, 42(1):63–72, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:2024:CEA

[HCO^{+}24] C. Ryan Hill, Abigale Culberson, Antóin M. O’Sullivan, Joke Adesola, R. Allen Curry, Tommi Linnansaari, and Philip M. Harrison. A controlled experimental assessment of intraperitoneal PIT tag injection on survival of blueback herring. North American Journal of Fisheries Management, 44(1):93–100, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hessler:2023:WHS

[HCP^{+}23] Tyler M. Hessler, Duane C. Chapman, Craig P. Paukert, Jeffrey C. Jolley, and Michael E. Byrne. Winter habitat selection and efficacy of telemetry to aid grass carp removal efforts in a large reservoir. North American Journal of Fisheries Management, 43(1):189–202, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herigan:2021:CTF

[HCS^{+}21] Garrett M. Herigan, Derek P. Crane, Mark C. Scott, Fred C. Rohde, and Dustin W. Smith. Comparison of two fish sampling techniques for low-conductivity, lowland headwater streams. North American Journal of Fisheries Management, 41(6):1781–1788, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harried:2020:PCL

[HDH^{+}20] Brittany L. Harried, Daniel J. Daugherty, David J. Hoeinghaus, Aaron P. Roberts, Barney J. Venables, Trent M. Sutton, and Brianne K. Soulen. Population contributions of large females may be eroded by contaminant body burden and maternal transfer: a case study of alligator gar. North American Journal of Fisheries Management, 40(3):566–579, June 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hargrove:2022:LBH

[HDP^{+}22] John S. Hargrove, Kevin J. Dockendorf, Kathryn M. Potoka, Christopher A. Smith, Valentina Alvarez, and James D. Austin. Largemouth bass hatchery contributions quantified via parentagebased tagging. North American Journal of Fisheries Management, 42(3):758–774, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heller:2022:OAM

[HDQ22] Megan R. Heller, Jeffrey C. Dillon, and Michael C. Quist. Occurrence, abundance, movement, and habitat associations of Bonneville cutthroat trout in tributaries to Bear Lake, Idaho–Utah. North American Journal of Fisheries Management, 42(3):684– 700, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heller:2022:PDH

[HDT^{+}22] Megan R. Heller, Jeffrey C. Dillon, Scott A. Tolentino, Carson J. Watkins, and Michael C. Quist. Population dynamics and harvest management of the Bonneville cutthroat trout fishery in Bear Lake, Idaho–Utah. North American Journal of Fisheries Management, 42(3):701–712, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hostetter:2023:RFA

[HEP^{+}23] Nathan J. Hostetter, Allen F. Evans, Quinn Payton, Daniel D. Roby, Donald E. Lyons, and Ken Collis. A review of factors affecting the susceptibility of juvenile salmonids to avian predation. North American Journal of Fisheries Management, 43(1): 244–256, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2023:SEI

[HFJ23] Adam G. Hansen, Collin J. Farrell, and Brett M. Johnson. Simulated effects of imperfect sterile sport fish stocking on persistence of fertile fish in new exploited populations. North American Journal of Fisheries Management, 43(4):908–934, August 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harrison:2022:BCU

[HG22] Hannah L. Harrison and Julie Grønsleth Gould. Big catch, undecided risks: Perspectives of risk, reward, and trade-offs in Alaska’s salmon enhancement program. North American Journal of Fisheries Management, 42(6):1433–1453, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hedden:2021:QCN

[HGH^{+}21] S. C. Hedden, K. B. Gido, C. K. Hedden, C. A. Pennock, B. R. Duran, B. A. Hines, E. I. Gilbert, M. C. McKinstry, S. L. Durst, and N. R. Franssen. Quantifying consumption of native fishes by nonnative channel catfish in a desert river. North American Journal of Fisheries Management, 41(S1):S82–S94, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hedden:2023:MSD

[HGH^{+}23] Skyler C. Hedden, Keith B. Gido, Crosby K. Hedden, Brian T. Hickerson, and William T. Stewart. Movement, not survival, differs between wild and hatchery-reared imperiled desert fishes. North American Journal of Fisheries Management, 43(5):1310– 1321, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henesy:2022:SMP

[HGM22] Joshua Henesy, Daniel Goetz, and John E. Mullican. Seasonal movement patterns and summertime use of thermal refuge areas by muskellunge in the nontidal Potomac River, Maryland. North American Journal of Fisheries Management, 42(5):1144– 1154, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hickerson:2021:SRN

[HGMR21] Brian T. Hickerson, Elizabeth R. Grube, Kent R. Mosher, and Anthony T. Robinson. Successful restoration of a native fish assemblage in the Blue River, Arizona. North American Journal of Fisheries Management, 41(3):746–756, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hermus:2020:GKA

[HGSS20] John W. Hermus, Jeremiah A. Gorne, Stephanie L. Shaw, and Greg G. Sass. Growth of known-age muskellunge stocked into Northwestern Wisconsin lakes. North American Journal of Fisheries Management, 40(5):1344–1350, October 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Horn:2022:MHB

[HHH^{+}22] Rebekah L. Horn, Maureen Hess, Stephanie Harmon, Jon Hess, Thomas A. Delomas, Matthew R. Campbell, and Shawn Narum. Monitoring hatchery broodstock composition and genetic variation of spring/summer Chinook salmon in the Columbia River Basin with multigeneration pedigrees. North American Journal of Fisheries Management, 42(2):307–333, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Horn:2023:MPM

[HHH^{+}23] Rebekah L. Horn, Maureen Hess, Stephanie Harmon, Jon Hess, Thomas A. Delomas, Matthew R. Campbell, and Shawn Narum. Multigeneration pedigrees to monitor hatchery broodstock composition and genetic variation of spring/summer Chinook salmon in the Columbia River Basin. North American Journal of Fisheries Management, 43(3):794–820, June 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoff:2021:PCR

[HIF21] Michael H. Hoff, Kevin S. Irons, and Jeffery E. Forester. Policy challenges relating to integrated pest management of freshwater aquatic invasive animals. North American Journal of Fisheries Management, 41(2):276–288, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holmes:2021:JCS

[HJ21] Eric J. Holmes and Carson A. Jeffres. Juvenile Chinook salmon weight prediction using image-based morphometrics. North American Journal of Fisheries Management, 41(2):446–454, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilling:2021:GDI

[HJB^{+}21] Corbin D. Hilling, Yan Jiao, Aaron J. Bunch, Robert S. Greenlee, Joseph D. Schmitt, and Donald J. Orth. Growth dynamics of invasive blue catfish in four subestuaries of the Chesapeake Bay, USA. North American Journal of Fisheries Management, 41(S1): S167–S179, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilling:2020:SSE

[HJBP20] Corbin D. Hilling, Yan Jiao, Aaron J. Bunch, and Quinton E. Phelps. A simulation study to evaluate biases in population characteristics estimation associated with varying bin numbers in sizebased age subsampling. North American Journal of Fisheries Management, 40(3):675–690, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huntsman:2020:LRA

[HKPP20] Brock M. Huntsman, Hae Kim, Quinton Phelps, and J. Todd Petty. Limestone remediation of an acidic stream creates a microchemical batch mark for brook trout within an Appalachian watershed. North American Journal of Fisheries Management, 40 (6):1523–1531, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Helminen:2023:CIS

[HL23] J. Helminen and T. Linnansaari. Combining imaging sonar counting and underwater camera species apportioning to estimate the number of Atlantic salmon and striped bass in the Miramichi River, New Brunswick, Canada. North American Journal of Fisheries Management, 43(3):743–757, June 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hager:2021:RFG

[HLDK21] Christian Hager, Juan C. Levesque, R. Jason Dickey, and Jason E. Kahn. Raised-footrope gill-net modification significantly reduces subadult Atlantic sturgeon bycatch. North American Journal of Fisheries Management, 41(1):19–25, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hess:2023:RAP

[HLJ^{+}23] J. E. Hess, R. T. Lampman, A. D. Jackson, T. Sween, L. Jim, N. McClain, G. Silver, L. Porter, and S. R. Narum. The return of the adult Pacific Lamprey offspring from translocations to the Columbia River. North American Journal of Fisheries Management, 43(6):1531–1552, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henderson:2023:EPD

[HLM^{+}23] Mark J. Henderson, Chris M. Loomis, Cyril J. Michel, Joe M. Smith, Ilysa S. Iglesias, Brendan M. Lehman, Nicholas J. Demetras, and David D. Huff. Estimates of predator densities using mobile DIDSON surveys: Implications for survival of Central Valley Chinook salmon. North American Journal of Fisheries Management, 43(3):628–645, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2023:SUE

[HLSW23] Julianne E. Harris, Theresa L. Liedtke, Joseph J. Skalicky, and Lisa K. Weiland. Salvage using electrofishing methods caused minimal mortality of burrowed and emerged larval lampreys in dewatered habitats. North American Journal of Fisheries Management, 43(6):1553–1566, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hubert:2021:BII

[HMB21] Terrance D. Hubert, James Miller, and Dale Burkett. A brief introduction to integrated pest management for aquatic systems. North American Journal of Fisheries Management, 41(2):264– 275, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hyatt:2021:ASI

[HMO^{+}21] Kim D. Hyatt, Donald J. McQueen, Athena D. Ogden, Ryan Benson, and Howie Wright. Age-structured interactions among reintroduced sockeye salmon, resident Kokanee, invasive mysids, and their zooplankton prey in Skaha Lake, British Columbia. North American Journal of Fisheries Management, 41(5):1246– 1273, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Healy:2022:MMF

[HMP22] Brian D. Healy, Jennifer F. Moore, and William E. Pine III. Monitoring and managing fishes that are invisible and keep moving around: Influences of an invasive species and environmental factors on capture probability. North American Journal of Fisheries Management, 42(2):396–409, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hemmelgarn:2022:CGS

[HMQM22] Grace L. Hemmelgarn, Benjamin Marcy-Quay, and J. Ellen Marsden. Contemporary growth and survival of stocked and wild lake trout in Lake Champlain evaluated using maxillary age estimates. North American Journal of Fisheries Management, 42(6): 1541–1549, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haley:2021:ESC

[HN21] Bryant M. Haley and J. Wesley Neal. Evaluation of supplemental crappie stocking in small southern reservoirs. North American Journal of Fisheries Management, 41(5):1514–1521, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2022:UOW

[HNM^{+}22] Michael J. Hansen, Nancy A. Nate, Andrew M. Muir, Louise Chavarie, Kimberly L. Howland, and Charles C. Krueger. Usefulness of otolith weight for estimating age-based life history metrics of lake trout. North American Journal of Fisheries Management, 42(5):1359–1371, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heckel:2020:DAW

[HQWD20] John W. Heckel IV, Michael C. Quist, Carson J. Watkins, and Andrew M. Dux. Distribution and abundance of westslope cutthroat trout in relation to habitat characteristics at multiple spatial scales. North American Journal of Fisheries Management, 40 (4):893–909, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2022:CID

[HRK^{+}22] Gretchen J. A. Hansen, Jenna K. Ruzich, Corey A. Krabbenhoft, Holly Kundel, Shad Mahlum, Christopher I. Rounds, Amanda O. Van Pelt, Lawrence D. Eslinger, Dale E. Logsdon, and Daniel A. Isermann. It’s complicated and it depends: a review of the effects of ecosystem changes on walleye and yellow perch populations in North America. North American Journal of Fisheries Management, 42(3):484–506, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hardy:2022:NRL

[HRM^{+}22] Ryan S. Hardy, Tyler J. Ross, Kevin N. McDonnell, Michael C. Quist, Charlie Holderman, and Bryan S. Stevens. Nutrient restoration of a large, impounded, ultra-oligotrophic western river to recover declining native fishes. North American Journal of Fisheries Management, 42(4):977–993, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunter:2020:ELC

[HRS^{+}20] Robert D. Hunter, Edward F. Roseman, Nick M. Sard, Daniel B. Hayes, Travis O. Brenden, Robin L. DeBruyne, and Kim T. Scribner. Egg and larval collection methods affect spawning adult numbers inferred by pedigree analysis. North American Journal of Fisheries Management, 40(2):307–319, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Healy:2024:QCT

[HS24] Brian D. Healy and Emily Omana Smith. Quantifying the contributions of tributaries to large-river fish populations through mark-recapture modeling. North American Journal of Fisheries Management, 44(2):299–318, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Howell:2022:CSB

[HSF^{+}22] Bradley E. Howell, Erin M. C. Stewart, Vince R. Frasca, Chris C. Wilson, and Graham D. Raby. Capture of spawning brook trout by electrofishing does not impair embryo survival. North American Journal of Fisheries Management, 42(1):228–235, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hogberg:2021:NRP

[HSH21] Nicholas P. Hogberg, Joseph A. Skorupski, Jr., and Samuel J. Hochhalter. Natural recruitment potential of a reintroduced shovelnose sturgeon population in the Bighorn River, Wyoming. North American Journal of Fisheries Management, 41(5):1288–1298, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamel:2021:CPC

[HSP21] Martin J. Hamel, Jonathan J. Spurgeon, and Mark A. Pegg. Catfish population characteristics among river segments with altered fluvial geomorphic conditions in the Missouri River, Nebraska, USA. North American Journal of Fisheries Management, 41(S1): S133–S145, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Healy:2020:ERP

[HSS^{+}20] Brian D. Healy, Emily C. Omana Smith, Robert C. Schelly, Melissa A. Trammell, and Clay B. Nelson. Establishment of a reproducing population of endangered humpback chub through translocations to a Colorado River tributary in Grand Canyon, Arizona. North American Journal of Fisheries Management, 40 (1):278–292, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harnish:2020:SCE

[HSTH20] Ryan Harnish, John Skalski, Rich Townsend, and Kenneth Ham. In search of a cost-effective approach for estimating dam passage survival. North American Journal of Fisheries Management, 40 (4):865–882, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heaston:2021:DEA

[HSW21] Emily D. Heaston, Catalina Segura, and Dana R. Warren. Do electrofishing activities disrupt stream biofilm standing stocks? An assessment from two headwater streams in Western Oregon. North American Journal of Fisheries Management, 41(2):466– 473, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hafen:2021:ECA

[HTH^{+}21] Thomas Hafen, Andrew T. Taylor, Dean A. Hendrickson, David R. Stewart, and James M. Long. Environmental conditions associated with occurrences of the threatened Yaqui catfish in the Yaqui River Basin, Mexico. North American Journal of Fisheries Management, 41(S1):S54–S63, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hooley-Underwood:2021:RSS

[HUTB21] Zachary E. Hooley-Underwood, Kevin G. Thompson, and Kevin R. Bestgen. Razorback sucker spawning in an intermittent Colorado tributary. North American Journal of Fisheries Management, 41(4):1151–1158, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hessenauer:2022:DBT

[HW22] Jan-Michael Hessenauer and Kevin Wehrly. Developing a brown trout stocking screening tool from a statewide random sampling program and landscape-scale predictor variables. North American Journal of Fisheries Management, 42(1):217–227, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heim:2024:TEA

[HWCS24] Kurt C. Heim, Jonah Withers, and Theodore Castro-Santos. Tagger effects in aquatic telemetry: Short-term and delayed effects of surgery in Atlantic salmon smolts. North American Journal of Fisheries Management, 44(2):262–275, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2022:ATR

[HYM^{+}22] Les. N. Harris, David J. Yurkowski, Brendan K. Malley, Samantha F. Jones, Brent G. T. Else, Ross F. Tallman, Aaron T. Fisk, and Jean-Sébastien Moore. Acoustic telemetry reveals the complex nature of mixed-stock fishing in Canada’s largest Arctic char commercial fishery. North American Journal of Fisheries Management, 42(5):1250–1268, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Izzo:2023:SLM

[IDH^{+}23] Lisa K. Izzo, Daniel Dembkowski, Todd Hayden, Tom Binder, Christopher Vandergoot, Steven Hogler, Michael Donofrio, Troy Zorn, Charles C. Krueger, and Daniel Isermann. Spawning locations, movements, and potential for stock mixing of walleye in Green Bay, Lake Michigan. North American Journal of Fisheries Management, 43(3):695–714, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Izzo:2021:SFR

[IPZK21] Lisa K. Izzo, Donna L. Parrish, Gayle Barbin Zydlewski, and Ryan Koenigs. Second fin ray shows promise for estimating ages of juvenile but not adult lake sturgeon. North American Journal of Fisheries Management, 41(1):217–228, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2021:NMU

[JCB^{+}21] Jennifer L. Johnson, Justin A. Chiotti, Andrew S. Briggs, James C. Boase, Jan-Michael Hessenauer, and Edward F. Roseman. Northern madtom use of artificial reefs in the St. Clair– Detroit River system. North American Journal of Fisheries Management, 41(S1):S42–S53, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:2020:RNR

[JDO^{+}20] Krista L. Jones, Jason B. Dunham, Jim E. O’Connor, Mackenzie K. Keith, Joseph F. Mangano, Kelly Coates, and Travis Mackie. River network and reach-scale controls on habitat for lamprey larvae in the Umpqua River Basin, Oregon. North American Journal of Fisheries Management, 40(6):1400–1416, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jensen:2023:GPR

[JJD^{+}23] B. L. Jensen, R. C. Johnson, J. J. Duda, C. O. Ostberg, T. J. Code, J. H. Mclean, K. D. Stenberg, K. A. Larsen, M. S. Hoy, and D. A. Beauchamp. Growth performance of rainbow trout in reservoir tributaries and implications for steelhead growth potential above Skagit River dams. North American Journal of Fisheries Management, 43(5):1427–1446, October 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johansen:2022:DDH

[JOH^{+}22] Kurt Johansen, Esben Moland Olsen, Tormod Haraldstad, Robert Arlinghaus, and Erik Höglund. Digital data help explain drivers of angler satisfaction: an example from Southern Norway.

North American Journal of Fisheries Management, 42(5):1165– 1172, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jensen:2021:AGS

[JSH^{+}21] Alexander J. Jensen, Carl B. Schreck, Jon E. Hess, Sandra Bohn, Kathleen G. O’Malley, and James T. Peterson. Application of genetic stock identification and parentage-based tagging in a mixedstock recreational Chinook salmon fishery. North American Journal of Fisheries Management, 41(1):130–141, February 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Josephson:2021:PHF

[JWG^{+}21] Ronald Josephson, Alex Wertheimer, David Gaudet, E. Eric Knudsen, Benjamin Adams, David R. Bernard, Steven C. Heinl, Andrew W. Piston, and William D. Templin. Proportions of hatchery fish in escapements of summer-run chum salmon in Southeast Alaska, 2013–2015. North American Journal of Fisheries Management, 41(3):724–738, June 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kasper:2020:UHS

[KBC^{+}20] Jacob M. Kasper, Jeffrey Brust, Amanda Caskenette, Jason Mc-Namee, Jason C. Vokoun, and Eric T. Schultz. Using harvest slot limits to promote stock recovery and broaden age structure in marine recreational fisheries: a case study. North American Journal of Fisheries Management, 40(6):1451–1471, December 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koemle:2021:HDR

[KBDA21] Dieter Koemle, Ben Beardmore, Malte Dorow, and Robert Arlinghaus. The human dimensions of recreational anglers targeting freshwater species in coastal ecosystems, with implications for management. North American Journal of Fisheries Management, 41(5):1572–1590, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krebs:2021:RED

[KBDI21] Jared E. Krebs, Ethan J. Brandt, Daniel J. Dembkowski, and Daniel A. Isermann. Relative effectiveness of D-frame dip nets, quatrefoil light traps, and towed ichthyoplankton nets for larval muskellunge. North American Journal of Fisheries Management, 41(5):1334–1340, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kreiser:2020:SAC

[KDB^{+}20] Brian R. Kreiser, Daniel J. Daugherty, David L. Buckmeier, Nathan G. Smith, and Eleanor B. Newsome. Sibship analysis to characterize alligator gar reproductive contributions in two Texas systems. North American Journal of Fisheries Management, 40 (3):555–565, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koel:2023:AAO

[KDM^{+}23] Todd M. Koel, Philip D. Doepke, Drew J. MacDonald, Nathan A. Thomas, Cody W. Vender, Hayley C. Glassic, Alex S. Poole, Christopher S. Guy, and Alexander V. Zale. Aerial application of organic pellets eliminates lake trout recruitment from a primary spawning reef in Yellowstone Lake. North American Journal of Fisheries Management, 43(2):505–516, April 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krogman:2022:MWR

[KDO^{+}22] Rebecca M. Krogman, Benjamin J. Dodd, Andy Otting, Michael J. Weber, and Robert E. Weber. Modeled walleye regulation outcomes in a Midwestern reservoir with varying forage and walleye density. North American Journal of Fisheries Management, 42(3):597–611, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keplinger:2021:AMP

[Kep21] Brandon J. Keplinger. Abundance, movement, population dynamics, and management implications of channel catfish in a mid-Atlantic River. North American Journal of Fisheries Management, 41(S1):S221–S233, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kock:2021:ASS

[KES^{+}21] Tobias J. Kock, Scott D. Evans, Brian P. Saluskin, Andrew P. Matala, Richard Visser, Mark V. Johnston, Peter F. Galbreath, and Adam C. Pope. Adult sockeye salmon responses to transplanting upstream of an impassable dam. North American Journal of Fisheries Management, 41(6):1640–1651, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kinziger:2022:PAI

[KFG22] Andrew P. Kinziger, Steven R. Fong, and John Carlos Garza. Pedigree analysis of an integrated hatchery steelhead program from the Mad River, California, provides insight into life history patterns and informs management. North American Journal of Fisheries Management, 42(5):1285–1295, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keplinger:2022:TTM

[KHB22] Brandon Keplinger, James Hedrick, and Vicki S. Blazer. Temporal trends in macroscopic indicators of fish health in the South Branch Potomac River. North American Journal of Fisheries Management, 42(2):277–294, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kroboth:2024:PFC

[KHBF24] P. T. Kroboth, A. M. Herndon, C. L. Broaddus, and J. R. Fischer. Precision of four calcified structures for age estimation of black carp. North American Journal of Fisheries Management, 44(1): 235–243, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knoth:2022:RCU

[KHD^{+}22] Brian A. Knoth, John S. Hargrove, Marika Dobos, Timothy Copeland, and Brett J. Bowersox. Rapid colonization of upstream habitats by Oncorhynchus mykiss following culvert modification. North American Journal of Fisheries Management, 42(5):1173– 1184, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kambestad:2022:ISM

[KHJS22] Marius Kambestad, Bjart Are Hellen, Knut Helge Midtbø Jensen, and Harald Sægrov. Improving salmonid monitoring by nocturnal counting in rivers. North American Journal of Fisheries Management, 42(5):1185–1195, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kanigan:2024:ATT

[KHL^{+}24] Adam M. Kanigan, Scott G. Hinch, Andrew G. Lotto, Kamil Szlachta, Stephen D. Johnston, and Stephanie A. Lingard. Acoustic telemetry tracking of Coho salmon smolts released from a community-run hatchery into a marine inlet reveals low early ocean survival. North American Journal of Fisheries Management, 44(2):359–376, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kinter:2022:EWS

[Kin22] Bryan T. Kinter. Estimating walleye and saugeye gill-net encounter probability in a reservoir with simple morphometry. North American Journal of Fisheries Management, 42(3):795– 812, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krause:2022:WEF

[KJB^{+}22] Jacob R. Krause, Eric C. Janney, Summer M. Burdick, Alta C. Harris, and Brian S. Hayes. Water and endangered fish in the Klamath River Basin: Do Upper Klamath Lake surface elevation and water quality affect adult Lost River and shortnose sucker survival? North American Journal of Fisheries Management, 42 (6):1414–1432, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koob:2023:UOM

[KKMA23] Elise R. Koob, Lisa A. Kerr, John W. Mandelman, and Michael P. Armstrong. Using otolith microchemistry to determine natal origin of Black Sea Bass off the coast of Maine. North American Journal of Fisheries Management, 43(4):1093–1110, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kazyak:2021:DGB

[KLR^{+}21] David C. Kazyak, Barbara A. Lubinski, Jacob M. Rash, Thomas C. Johnson, and Timothy L. King. Development of genetic baseline information to support the conservation and management of wild brook trout in North Carolina. North American Journal of Fisheries Management, 41(3):626–638, June 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knapp:2021:MAF

[KMS21] Michael L. Knapp, Steven W. Mero, and David F. Staples. Are muskellunge affecting fish communities in waters where they have been introduced? A re-examination of Minnesota’s stocked muskellunge waters. North American Journal of Fisheries Management, 41(1):229–241, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefer:2020:RMH

[KNC^{+}20] Matthew L. Keefer, Christopher J. Noyes, Tami S. Clabough, Daniel C. Joosten, and Christopher C. Caudill. Rapid migration and high survival of adult Pacific lampreys in reservoirs. North American Journal of Fisheries Management, 40(2):354–367, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klein:2022:RBG

[KNK22] Zachary B. Klein, Ben C. Neely, and Jeff D. Koch. Relationships between gizzard shad, impoundment characteristics, and sympatric fish species. North American Journal of Fisheries Management, 42(2):447–454, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kirsch:2024:IDM

[KPD^{+}24] Joseph E. Kirsch, James T. Peterson, Adam Duarte, Denise Goodman, Andrew Goodman, Sara Hugentobler, Mariah Meek, Russell W. Perry, Corey Phillis, Lori Smith, and Jeffrey Stuart. Imperfect detection and misidentification affect inferences from data informing water operation decisions. North American Journal of Fisheries Management, 44(2):335–358, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klein:2023:SIF

[KQG23] Zachary B. Klein, Michael C. Quist, and Christopher S. Guy. Suppression of invasive fish in the West: Synthesis and suggestions for improvement. North American Journal of Fisheries Management, 43(2):369–383, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klein:2020:IPD

[KQS^{+}20] Zachary B. Klein, Michael C. Quist, Daniel J. Schill, Andrew M. Dux, and Matthew P. Corsi. Influence of population density and length structure on angler catch rate in Kokanee fisheries. North American Journal of Fisheries Management, 40(1):182– 189, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krall:2023:ELE

[KR23] Michelle Krall and Philip Roni. Effects of livestock exclusion on stream habitat and aquatic biota: a review and recommendations for implementation and monitoring. North American Journal of Fisheries Management, 43(2):476–504, April 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keplinger:2024:EPR

[KR24] Brandon Keplinger and Christopher T. Rota. Effects of poor recruitment on riverine smallmouth bass population dynamics. North American Journal of Fisheries Management, 44(2):247– 261, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

King:2020:EAD

[KS20] Sarah M. King and Jeffrey A. Stein. Effective anesthetic dosage and recovery from specialized surgical methods in shortnose gar. North American Journal of Fisheries Management, 40(6):1486– 1498, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krogman:2021:SPC

[KS21] Rebecca M. Krogman and Tyler J. Stubbs. Strategic planning for catfish angling in the future: An example from urban Iowa. North American Journal of Fisheries Management, 41(S1):S298– S313, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Khalsa:2020:IUI

[KSJ^{+}20] Noah S. Khalsa, Justin Smith, Kim A. Jochum, Garrett Savory, and J. Andrés López. Identifying under-ice overwintering locations of juvenile Chinook salmon by using environmental DNA. North American Journal of Fisheries Management, 40(3):762– 772, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kissinger:2024:EBT

[KSP^{+}24] Benjamin C. Kissinger, Michael G. Sullivan, Andrew J. Paul, Adrian Meinke, and John Post. Establishment of bull trout in a previously fishless subalpine lake by translocation. North American Journal of Fisheries Management, 44(2):520–531, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kramer:2022:PBR

[KSPP22] Randy J. Kramer, Cory D. Suski, Anthony P. Porreca, and Joseph J. Parkos. Physiological and behavioral responses of age-0 muskellunge during simulated stocking in elevated pH water. North American Journal of Fisheries Management, 42(2):361– 368, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kendall:2023:LCM

[KUV^{+}23] Neala W. Kendall, Julia Unrein, Carol Volk, David A. Beauchamp, Kurt L. Fresh, and Thomas P. Quinn. Life cycle model reveals sensitive life stages and evaluates recovery options for a dwindling Pacific salmon population. North American Journal of Fisheries Management, 43(1):203–230, February 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lochmann:2021:MMU

[LBH21] Steve Lochmann, Eric L. Brinkman, and Dylan A. Hann. Movements and macrohabitat use of alligator gar in relation to a low head lock and dam system. North American Journal of Fisheries Management, 41(1):204–216, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leblanc:2020:RFR

[LCF20] John Paul Leblanc, Anna Conklyn, and John M. Farrell. Rapid fin regeneration of age-0 northern pike and implications of fin clips as a marking protocol. North American Journal of Fisheries Management, 40(5):1294–1301, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lenaerts:2023:EBC

[LCIL23a] Allison W. Lenaerts, Alison A. Coulter, Kevin S. Irons, and James T. Lamer. Examination of bigheaded carp ovaries indicates batch spawning. North American Journal of Fisheries Management, 43(1):25–34, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lenaerts:2023:PRP

[LCIL23b] Allison W. Lenaerts, Alison A. Coulter, Kevin S. Irons, and James T. Lamer. Plasticity in reproductive potential of bigheaded carp along an invasion front. North American Journal of Fisheries Management, 43(1):92–100, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lenaerts:2023:QME

[LCIL23c] Allison W. Lenaerts, Alison A. Coulter, Kevin S. Irons, and James T. Lamer. A quick method for estimating batch fecundity in bigheaded carp. North American Journal of Fisheries Management, 43(1):3–11, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lewis:2023:RFE

[LCM^{+}23] Madeline C. Lewis, W. Robert Cope, Thomas P. Miles, Claire Rude, Richard E. Bruesewitz, Benjamin J. Dodd, Mark K. Flammang, Kevin S. Page, Robert Weber, Michael J. Weber, and Max Wolter. Reservoir fish escapement in North America: a historical review and future directions. North American Journal of Fisheries Management, 43(2):352–368, April 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Litts:2023:BEC

[LCVS23] Richard Litts, Benjamin J. Clemens, Gary Vonderohe, and John Schaefer. Backpack electrofishing can be used to collect adult lamprey. North American Journal of Fisheries Management, 43 (6):1623–1630, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loher:2022:EPL

[LDH^{+}22] Timothy Loher, Claude L. Dykstra, Allan C. Hicks, Ian J. Stewart, Nathan Wolf, Bradley P. Harris, and Josep V. Planas. Estimation of postrelease longline mortality in Pacific halibut using acceleration-logging tags. North American Journal of Fisheries Management, 42(1):37–49, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lewis:2021:GEI

[LEHP21] Matthew R. Lewis, Phil Ekema, Mike Holley, and Eric J. Peatman. Genetic evidence of introduced redeye bass and Alabama bass and hybridization with native Micropterus spp. in Town Creek, Alabama, USA. North American Journal of Fisheries Management, 41(1):78–85, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Losee:2020:EVI

[LFS20] James P. Losee, Hannah S. Faulkner, and Todd R. Seamons. Evaluation of visible implant elastomer tags in wild coastal cutthroat trout in the marine environment. North American Journal of Fisheries Management, 40(6):1443–1450, December 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lewis:2022:TVC

[LGOM22] Madeline C. Lewis, Christopher S. Guy, Eric W. Oldenburg, and Thomas E. McMahon. Temporal variation in capture efficiency underrepresents spring out-migrating bull trout in a trap-and-haul program. North American Journal of Fisheries Management, 42(5):1237–1249, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Liu:2021:FDE

[LGQ^{+}21] Han Liu, Chuanbo Guo, Xiao Qu, Fangyuan Xiong, Craig P. Paukert, Yushun Chen, and Wei Su. Fish diversity, endemism, threats, and conservation in the Jinsha River Basin (Upper Yangtze River), China. North American Journal of Fisheries Management, 41(4):967–984, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Liedtke:2023:SLL

[LHB^{+}23] Theresa L. Liedtke, Julianne E. Harris, Monica R. Blanchard, Joseph J. Skalicky, and Ann B. Grote. Synthesis of larval lamprey responses to dewatering: State of the science, critical uncertainties, and management implications: Native lampreys: Research and conservation of ancient fishes. North American Journal of Fisheries Management, 43(6):1475–1491, December 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

LaRochelle:2024:DLW

[LHB^{+}24] Luc LaRochelle, Laura Haniford, Declan Burton, John F. Bieber, Jessica A. Robichaud, Cory D. Suski, Andy J. Danylchuk, and Steven J. Cooke. Do live-well additives influence the physiological and behavioral recovery of largemouth bass? North American Journal of Fisheries Management, 44(1):189–203, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Longo:2024:PGC

[LHPS^{+}24] Gary C. Longo, Melissa A. Head, Sandy L. Parker-Stetter, Ian G. Taylor, Vanessa J. Tuttle, Alicia A. Billings, Stéphane Gauthier, Michelle M. McClure, and Krista M. Nichols. Population genomics of coastal Pacific hake. North American Journal of Fisheries Management, 44(1):222–234, February 2024. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larson:2020:UGD

[LHR^{+}20] Wesley A. Larson, Thompson Hill, David Rowe, Daniel Oele, Joseph Gerbyshak, and Jennifer Bergman. Using genetic data to estimate capture rate of Wisconsin and Leech Lake strains of muskellunge stocked in four Wisconsin lakes. North American Journal of Fisheries Management, 40(5):1302–1312, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Liedtke:2023:BPT

[LHWS23] Theresa L. Liedtke, Julianne E. Harris, Christina J. Wang, and Trent M. Sutton. Bringing partners together: a symposium on native lampreys and the Pacific Lamprey Conservation Initiative. North American Journal of Fisheries Management, 43(6): 1449–1457, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lubenau:2024:ERC

[LJB^{+}24] William J. Lubenau, Timothy R. Johnson, Brett J. Bowersox, Timothy Copeland, Joshua L. McCormick, and Michael C. Quist. Encounter rates and catch-and-release mortality of steelhead in the Snake River basin. North American Journal of Fisheries Management, 44(1):3–20, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lampton:2023:EOT

[LKD23] Jared Lampton, Ryan Kovach, and James L. Dunnigan. Efficacy of one-time rotenone application for Westslope cutthroat trout restoration in Boulder Creek, Montana. North American Journal of Fisheries Management, 43(4):1081–1092, August 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lackmann:2021:TSO

[LKP^{+}21] Alec R. Lackmann, Storm Kettelhut, Brandon L. Paulson, Cody M. Anderson, Shawn C. Goodchild, Kevin D. Guadalupe, and Craig A. Stockwell. Thin-sectioned otoliths reveal sexual dimorphism and a 10-year lifespan in the endangered pahrump poolfish. North American Journal of Fisheries Management, 41 (6):1631–1639, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lamglait:2021:CSE

[LL21] Benjamin Lamglait and Stéphane Lair. Comparative study on electric fish handling gloves and immersion anesthesia for the surgical implantation of transmitters in brook trout. North American Journal of Fisheries Management, 41(1):103–114, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lederman:2022:UPF

[LLA^{+}22] Nathan J. Lederman, Seth A. Love, Rebekah L. Anderson, Jason A. DeBoer, and James T. Lamer. Using pectoral fin ridges to assign sex to bighead carp. North American Journal of Fisheries Management, 42(4):1003–1010, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lawson:2022:SPL

[LLE22] Zach J. Lawson, Alexander W. Latzka, and Lawrence Eslinger. Stocking practices and lake characteristics influence probability of stocked walleye survival in Wisconsin’s ceded territory lakes. North American Journal of Fisheries Management, 42(3):523– 534, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loeppky:2020:IHR

[LMA20] A. R. Loeppky, C. A. McDougall, and W. G. Anderson. Identification of hatchery-reared lake sturgeon using natural elemental signatures and elemental marking of fin rays. North American Journal of Fisheries Management, 40(1):61–74, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lusk:2023:EPC

[LMA23] Sean C. Lusk, Christopher R. Middaugh, and Allison M. Asher. Evaluation of population characteristic differences between the largemouth bass and its backcross following a long-term Florida bass stocking effort. North American Journal of Fisheries Management, 43(3):603–617, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lacho:2020:EDR

[LMN^{+}20] Christine D. Lacho, Craig A. McDougall, Patrick A. Nelson, Michael M. Legge, Mark A. Gillespie, Yhana Michaluk, Cheryl N. Klassen, and Don Macdonald. Evaluation of a deepwater release method for hatchery-reared lake sturgeon. North American Journal of Fisheries Management, 40(4):828–839, August 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lusk:2021:EPM

[LMO21] Sean C. Lusk, Christopher R. Middaugh, and Derek H. Ogle. Evaluating the performance of methods used to estimate growth parameters from subsampled age data. North American Journal of Fisheries Management, 41(3):570–584, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lowman:2023:EBA

[LOC23] Brooke A. Lowman, Catherine E. O’Keefe, and Steven X. Cadrin. Evaluating bycatch avoidance in the U.S. Atlantic sea scallop Placopecten magellanicus fishery. North American Journal of Fisheries Management, 43(6):1713–1724, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lamy:2023:GAR

[LOV^{+}23] Jared B. Lamy, Brigid C. O’Donnell, Amy M. Villamagna, Tyson R. Morrill, Ben J. Nugent, and Joshua C. Hoekwater. Genetic analysis reveals a complex mosaic of admixture in brook trout in a historically fragmented watershed. North American Journal of Fisheries Management, 43(4):947–961, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lawson:2020:SPE

[LRM^{+}20] Katelyn M. Lawson, Josey L. Ridgway, Andy T. Mueller, Jacob D. A. Faulkner, and Robin D. Calfee. Semiautomated process for enumeration of fishes from recreational-grade side-scan sonar imagery. North American Journal of Fisheries Management, 40 (1):75–83, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lewis:2021:FPM

[LSB^{+}21] Matthew R. Lewis, Katherine E. Silliman, Benjamin Beck, Steven M. Sammons, and Eric J. Peatman. Failure of phenotypic markers to accurately identify black bass species and associated hybrids in the Mobile River Basin, Alabama. North American Journal of Fisheries Management, 41(5):1591–1601, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Long:2020:ETH

[LSP20] James M. Long, Richard A. Snow, and Michael J. Porta. Effects of temperature on hatching rate and early development of alligator gar and spotted gar in a laboratory setting. North American Journal of Fisheries Management, 40(3):661–668, June 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Long:2023:VCA

[LSSB23] James M. Long, Richard A. Snow, Daniel E. Shoup, and Jory B. Bartnicki. Validation and comparison of age estimates for smallmouth buffalo in Oklahoma based on otoliths, pectoral fin rays, and opercula. North American Journal of Fisheries Management, 43(3):618–627, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

LaRochelle:2022:CDT

[LTB^{+}22] Luc LaRochelle, Alexandria Trahan, Jacob W. Brownscombe, Andy J. Danylchuk, and Steven J. Cooke. A comparison of different tournament weigh-in formats on the short-term postrelease behavior of black bass assessed with biologgers. North American Journal of Fisheries Management, 42(2):250–259, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

LaHood:2023:LTR

[LTL^{+}23] Boone M. La Hood, Tyler C. Thomsen, Allison W. Lenaerts, Madeline G. Tomczak, Emily A. Szott, Zebadiah Woiak, Kyle M. Von Ruden, Katherine D. Bockrath, Kevin S. Irons, and James T. Lamer. Light trapping reveals multiple bigheaded carp spawns upstream of Lock and Dam 19 in the Upper Mississippi River. North American Journal of Fisheries Management, 43(1):81–91, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Midway:2020:ESR

[MAB^{+}20] Stephen R. Midway, Jason Adriance, Patrick Banks, Sepp Haukebo, and Rex Caffey. Electronic self-reporting: Angler attitudes and behaviors in the recreational red snapper fishery. North American Journal of Fisheries Management, 40(5):1119– 1132, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moore:2021:EVI

[MB21] Desiree M. Moore and Shannon K. Brewer. Evaluation of visual implant elastomer, PIT, and p-chip tagging methods in a smallbodied minnow species. North American Journal of Fisheries Management, 41(4):1066–1078, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meeuwig:2022:ITC

[MBCH22] Michael H. Meeuwig, Elizabeth J. Bailey, Shaun P. Clements, and Brett L. Hodgson. Inferred trophic characteristics and ecological roles of native and nonnative fishes in Odell Lake, Oregon. North American Journal of Fisheries Management, 42(5):1301– 1323, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKenna:2023:APS

[MBF^{+}23] Jorden R. McKenna, Anjanette Bowen, John R. Farver, James M. Long, Jeffrey G. Miner, Nathan D. Stott, and Patrick M. Kočovský. Assessing potential spawning locations of silver chub in Lake Erie. North American Journal of Fisheries Management, 43(5):1166–1179, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKinney:2022:NGT

[MBP^{+}22] Garrett J. McKinney, Patrick D. Barry, Carita Pascal, James E. Seeb, Lisa W. Seeb, and Megan V. McPhee. A new genotyping-in-thousands-by-sequencing single nucleotide polymorphism panel for mixed-stock analysis of chum salmon from Coastal Western Alaska. North American Journal of Fisheries Management, 42(5):

1134–1143, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2021:ETL

[MBW21] Joshua L. McCormick, Arnold F. Brimmer, and Carson J. Watkins. Evaluating target levels of American white pelican abundance to achieve management objectives for adfluvial Yellowstone cutthroat trout. North American Journal of Fisheries Management, 41(5):1274–1287, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Molina:2020:ACV

[MCB20] Joseph T. Molina, Derek P. Crane, and Lindsey Bell. Accuracy of concert versus independent age estimates made from anal fin rays of muskellunge. North American Journal of Fisheries Management, 40(5):1195–1201, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKenna:2023:CDS

[McK23] James E. McKenna, Jr. Conservation decision support for silver chub habitat in Lake Erie. North American Journal of Fisheries Management, 43(5):1151–1165, October 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2021:UAG

[MDA21] Paul H. Michaletz, H. Ross Dames, and Steffanie L. Abel. Use of angling gear to evaluate the potential benefits of minimum length limits on a highly mobile Midwestern blue catfish population. North American Journal of Fisheries Management, 41(S1):S234– S246, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mainguy:2021:IME

[MdAM21] Julien Mainguy and Rafael de Andrade Moral. An improved method for the estimation and comparison of mortality rates in fish from catch-curve data. North American Journal of Fisheries Management, 41(5):1436–1453, October 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2021:EMS

[MDBC21] Joshua L. McCormick, Marika E. Dobos, Brett J. Bowersox, and Timothy Copeland. Evaluation of management strategies for an incidental catch-and-release steelhead fishery. North American Journal of Fisheries Management, 41(2):498–512, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:2023:LTP

[MDD^{+}23] Charles P. Madenjian, Patricia M. Dieter, Timothy J. Desorcie, Stephen A. Lengnick, Timothy P. O’Brien, Lynn M. Benes, Steven A. Farha, and Benjamin S. Leonhardt. Lake trout population dynamics in the Northern Refuge of Lake Michigan: Importance of stocking rate. North American Journal of Fisheries Management, 43(4):1017–1034, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCarrick:2022:STD

[MDHQ22] Darcy K. McCarrick, Jeffrey C. Dillon, Brett High, and Michael C. Quist. Spatial and temporal distribution and habitat selection of native Yellowstone cutthroat trout and nonnative Utah chub. North American Journal of Fisheries Management, 42(4):939–951, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marschall:2021:MLA

[MGMP21] Elizabeth A. Marschall, David C. Glover, Martha E. Mather, and Donna L. Parrish. Modeling larval American shad recruitment in a large river. North American Journal of Fisheries Management, 41(4):939–954, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mokdad:2022:UDD

[MGNP22] A. I. Mokdad, S. R. Garner, B. D. Neff, and T. E. Pitcher. Upstream and downstream dispersal behavior of hard- and softreleased juvenile Atlantic salmon. North American Journal of Fisheries Management, 42(2):438–446, April 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2020:BPA

[Mic20a] Paul H. Michaletz. Bolstering piscivore abundance to restructure small impoundment fish communities. North American Journal of Fisheries Management, 40(5):1276–1293, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2020:VLA

[Mic20b] Paul H. Michaletz. Variable length at age for black crappies and white crappies in Missouri small impoundments. North American Journal of Fisheries Management, 40(5):1216–1224, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlbauer:2021:EPR

[MK21] Savannah R. Muhlbauer and Rebecca M. Krogman. Evaluation of potential regulations for improving flathead catfish size structure in Iowa’s large reservoirs. North American Journal of Fisheries Management, 41(S1):S268–S276, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Melstrom:2023:VAR

[MKC^{+}23] Richard T. Melstrom, Mark A. Kaemingk, Nicholas W. Cole, John C. Whitehead, Christopher J. Chizinski, and Kevin L. Pope. Valuing angling on reservoirs using benefit transfer. North American Journal of Fisheries Management, 43(2):400–416, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:2022:EHE

[MKG^{+}22] Charles P. Madenjian, Kevin M. Keeler, Lyle T. Grivicich, Paul E. Haver, Roger R. Gordon Jr, Donald E. Tillitt, and Jacques Rinchard. Egg hatchability and egg thiamine concentrations in lake trout from the northern refuge of Lake Michigan. North American Journal of Fisheries Management, 42(5):1155–1164, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morgan:2021:GSM

[MKK^{+}21] Raymond P. Morgan II, David C. Kazyak, Tim L. King, Barbara A. Lubinski, Matthew T. Sell, Alan A. Heft, and Jess W. Jones. Genetic structure of Maryland brook trout populations: Management implications for a threatened species. North American Journal of Fisheries Management, 41(4):1097–1119, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miazga:2022:EHR

[MKKN22] James R. Miazga, Zachary B. Klein, Jeff D. Koch, and Ben C. Neely. The effect of harvest regulations on white crappie fisheries in Kansas impoundments. North American Journal of Fisheries Management, 42(6):1623–1634, December 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mueller:2021:UCP

[MLB21] Robert P. Mueller, Ralph T. Lampman, and Tyler E. Beals. Using a customized portable deepwater electrofisher to assess larval lamprey populations in irrigation canals. North American Journal of Fisheries Management, 41(4):1124–1130, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madden:2023:CPM

[MLB^{+}23] Jamie C. Madden, Luc LaRochelle, Declan Burton, Andy J. Danylchuk, Sean J. Landsman, and Steven J. Cooke. Condition and postrelease mortality of angled Northern Pike temporarily retained on stringers. North American Journal of Fisheries Management, 43(5):1395–1406, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maine:2023:PIS

[MMJW23] Alexa N. Maine, Mary L. Moser, Aaron D. Jackson, and Frank Wilhelm. Probiotics improve survival and growth of larval Pacific Lamprey in laboratory culture. North American Journal of Fisheries Management, 43(6):1648–1663, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McEachran:2022:PLB

[MML^{+}22] Margaret C. McEachran, Alicia Hofelich Mohr, Thomas Lindsay, David C. Fulton, and Nicholas B. D. Phelps. Patterns of live baitfish use and release among recreational anglers in a regulated landscape. North American Journal of Fisheries Management, 42(2):295–306, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moser:2023:OCS

[MMSJ23] Mary Moser, Alexa Maine, Tristan Shonat, and Aaron Jackson. Oxygen consumption of sexually mature adult, first-feeding larval, and yearling Pacific Lampreys. North American Journal of Fisheries Management, 43(6):1664–1672, December 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McDougall:2020:HRL

[MNA^{+}20] Craig A. McDougall, Patrick A. Nelson, James K. Aiken, Duncan C. Burnett, Cameron C. Barth, Donald S. MacDonell, Yhana Michaluk, Cheryl N. Klassen, and Don Macdonald. Hatchery rearing of lake sturgeon to age 1 prior to stocking: a path forward for species recovery in the Upper Nelson River, Manitoba, Canada. North American Journal of Fisheries Management, 40 (4):807–827, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moore:2021:ELS

[MPM21] Michael J. Moore, Craig P. Paukert, and Travis L. Moore. Effects of latitude, season, and temperature on lake sturgeon movement. North American Journal of Fisheries Management, 41(4):916– 928, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McClure:2020:MDS

[MQK^{+}20] Conor McClure, Michael C. Quist, Joseph R. Kozfkay, Michael P. Peterson, and Daniel J. Schill. Movement dynamics of smallmouth bass in a large Western river system. North American Journal of Fisheries Management, 40(1):154–162, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maitland:2021:NFS

[MR21a] Bryan M. Maitland and Frank J. Rahel. Nonlethal fin sampling of North American freshwater fishes for food web studies using stable isotopes. North American Journal of Fisheries Management, 41(2):410–420, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Markham:2021:IST

[MR21b] James L. Markham and Jason M. Robinson. Influence of steelhead trout stocking size and location on emigration and adult returns. North American Journal of Fisheries Management, 41(6): 1652–1664, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mallory:2022:BLA

[MRK^{+}22] Mark L. Mallory, Gregory J. Robertson, Shane Keegan, Ingrid L. Pollet, Les N. Harris, Tyler Jivan, and Jennifer F. Provencher. Bycatch of loons assessed in coastal Arctic char fisheries in the Canadian Arctic. North American Journal of Fisheries Management, 42(5):1215–1225, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mock:2021:ERS

[MRM^{+}21] Alan J. Mock, Carl R. Ruetz III, James N. McNair, Dan Mays, and Archie Martell. Evaluating remote site incubators in Michigan streams: Implications for Arctic grayling reintroduction. North American Journal of Fisheries Management, 41(2):434– 445, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2024:PSW

[MRW^{+}24] Brett T. Miller, Elizabeth A. Renner, Kyle R. Winders, Juju C. Wellemeyer, and Hae H. Kim. Proposed standard weight (W_{s}) equation and standard length categories for goldeye. North American Journal of Fisheries Management, 44(1):216–221, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Montague:2021:TDA

[MS21a] Graham F. Montague and Daniel E. Shoup. Two decades of advancement in flathead catfish research. North American Journal of Fisheries Management, 41(S1):S11–S26, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moran:2021:ECB

[MS21b] Zachary S. Moran and Joseph N. Stoeckel. Effects of chase boat use, electrofishing duration, and water velocity on sample efficiency and size structure of blue catfish. North American Journal of Fisheries Management, 41(S1):S388–S399, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Montague:2022:APO

[MS22] Graham F. Montague and Daniel E. Shoup. Accuracy, precision, and optimal sampling duration of low-frequency electrofishing for sampling reservoir flathead catfish populations. North American Journal of Fisheries Management, 42(5):1269–1284, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mussmann:2022:RMD

[MSD^{+}22] Steven M. Mussmann, Melody J. Saltzgiver, Thomas A. Delomas, Matthew R. Campbell, Tracy A. Diver, and Wade D. Wilson. Rapid molecular determination of sex for Western North American chub (Gila spp.). North American Journal of Fisheries Management, 42(1):73–84, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michel:2020:LAR

[MSL^{+}20] Cyril J. Michel, Joseph M. Smith, Brendan M. Lehman, Nicholas J. Demetras, David D. Huff, Patricia L. Brandes, Joshua A. Israel, Thomas P. Quinn, and Sean A. Hayes. Limitations of active removal to manage predatory fish populations. North American Journal of Fisheries Management, 40(1):3–16, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Midway:2022:BCB

[MSL22] Stephen R. Midway, Allen Schaefer, and Julie A. Lively. Bycatch in a coastal black drum trotline fishery. North American Journal of Fisheries Management, 42(5):1372–1378, October 2022. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMillan:2022:HRR

[MSLP22] John R. McMillan, Matthew R. Sloat, Martin Liermann, and George Pess. Historical records reveal changes to the migration timing and abundance of winter steelhead in Olympic Peninsula rivers, Washington State, USA. North American Journal of Fisheries Management, 42(1):3–23, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2021:SUY

[MSM21] Joshua L. McCormick, Daniel J. Schill, and Kevin A. Meyer. Simulated use of YY male stocking and suppression for eradicating common carp populations. North American Journal of Fisheries Management, 41(2):366–382, April 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moran:2021:EVH

[MSP21] Zachary S. Moran, Joseph N. Stoeckel, and Brandon K. Peoples. Environmental variation and habitat use affect blue catfish lowfrequency electrofishing catch rates. North American Journal of Fisheries Management, 41(S1):S103–S117, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McDonald:2022:GFA

[MSP^{+}22] Dusty L. McDonald, J. Warren Schlechte, Donovan A. Patterson, Gregory R. Binion, and Amanda Boyles. Ghost fishing by abandoned trotlines in a simulated and actual reservoir. North American Journal of Fisheries Management, 42(4):839–848, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Montague:2023:BSR

[MSSS23] Graham F. Montague, Jason D. Schooley, Dennis L. Scarnecchia, and Richard A. Snow. Bowfishing shoot and release: High short-term mortality of nongame fishes and its management implications. North American Journal of Fisheries Management, 43 (4):962–983, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murdoch:2022:AMS

[MST22] Andrew R. Murdoch, Kevin See, and Benjamin L. Truscott. Abundance and migration success of overshoot steelhead in the Upper Columbia River. North American Journal of Fisheries Management, 42(4):1066–1080, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mrnak:2021:HSU

[MSV^{+}21] Joseph T. Mrnak, Logan W. Sikora, M. Jake Vander Zanden, Thomas R. Hrabik, and Greg G. Sass. Hydroacoustic surveys underestimate yellow perch population abundance: The importance of considering habitat use. North American Journal of Fisheries Management, 41(4):1079–1087, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maahs:2022:ELB

[MSW22] Brandon Maahs, Andrea Sylvia, and Michael J. Weber. Effects of length and bag limits on largemouth bass tournament capture and mortality. North American Journal of Fisheries Management, 42 (6):1513–1529, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Montague:2023:SRT

[MSZ^{+}23] Graham F. Montague, Daniel E. Shoup, Douglas L. Zentner, Richard A. Snow, and Patrick B. Cooney. Surfacing response thresholds of blue catfish exposed to low-frequency electrofishing. North American Journal of Fisheries Management, 43(3): 677–686, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mundahl:2022:ACH

[Mun22] Neal D. Mundahl. Analysis of a catch-and-harvest fishing contest for brown trout spanning 50+ years: Long-term trends influenced by fisheries management actions and angler behavior. North American Journal of Fisheries Management, 42(6):1653–1664, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2022:IMH

[MVH^{+}22] Joshua L. McCormick, Jennifer Vincent, Brett High, Darcy K. McCarrick, and Michael C. Quist. Informing management of Henrys Lake, Idaho, using an integrated catch-at-age model. North American Journal of Fisheries Management, 42(4):857–873, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McLaren:2023:DHK

[MVM^{+}23] John S. McLaren, Robert W. Van Kirk, Arthur J. Mabaka, Soren Brothers, and Phaedra Budy. Drawdown, habitat, and Kokanee populations in a Western U.S. reservoir. North American Journal of Fisheries Management, 43(2):339–351, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meerbeek:2020:MSI

[MW20] Jonathan R. Meerbeek and Michael J. Weber. Muskellunge survival, interbasin movement, and emigration in a simple and a complex interconnected glacial lake. North American Journal of Fisheries Management, 40(5):1146–1160, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McDermid:2020:GIL

[MWASW20] Jenni L. McDermid, Jordan Walker, Mohammed Al-Shamlih, and Chris C. Wilson. Genetic integrity of lake trout in Cold Lake, Alberta, despite decades of supplemental stocking. North American Journal of Fisheries Management, 40(2):459–474, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McGrath:2022:PHP

[MWMH22] Patrick E. McGrath, Brian E. Watkins, Ashleigh Magee, and Eric J. Hilton. Patterns of hatchery-produced returns of American shad in the James River, Virginia. North American Journal of Fisheries Management, 42(4):906–914, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mack:2021:UAT

[MWR21] Kevin Mack, Holly White, and Fred C. Rohde. Use of acoustic telemetry to identify spawning river and spawning migration patterns of American shad in the Albemarle Sound, North Carolina. North American Journal of Fisheries Management, 41(1): 242–251, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nepal:2021:RCD

[NF21] Vaskar Nepal and Mary C. Fabrizio. Reproductive characteristics differ in two invasive populations of blue catfish. North American Journal of Fisheries Management, 41(S1):S180–S194, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Norris:2020:AEG

[NHC^{+}20] David M. Norris, Hunter R. Hatcher, Michael E. Colvin, Giancarlo Coppola, Marcus A. Lashley, and Leandro E. Miranda. Assessing establishment and growth of agricultural plantings on reservoir mudflats. North American Journal of Fisheries Management, 40(2):394–405, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Naughton:2023:PMS

[NKC^{+}23] George P. Naughton, Matthew L. Keefer, Tami S. Clabough, Matthew J. Knoff, Timothy J. Blubaugh, Mark R. Morasch, Cameron S. Sharpe, and Christopher C. Caudill. Prespawn mortality of spring Chinook salmon in three Willamette River populations. North American Journal of Fisheries Management, 43(3): 715–729, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neely:2023:ELI

[NKG23] Ben C. Neely, Jeff D. Koch, and Keith B. Gido. Effects of liveimaging sonar on Blue Catfish angler success, perception, and behavior. North American Journal of Fisheries Management, 43 (6):1765–1771, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neely:2021:RMT

[NKK21] Ben C. Neely, Jeff D. Koch, and Nicholas W. Kramer. A review of marking and tagging methods for blue catfish, channel catfish, and flathead catfish. North American Journal of Fisheries Management, 41(S1):S415–S427, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nemec:2021:DEH

[NLB21] Zach C. Nemec, Larissa N. Lee, and Scott A. Bonar. Development and evaluation of habitat suitability criteria for native fishes in three Arizona streams. North American Journal of Fisheries Management, 41(3):661–677, June 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neely:2021:CGB

[NLK21] Ben C. Neely, Sean T. Lynott, and Jeff D. Koch. Comparison of growth, body weight, and contribution to recreational creel of channel catfish and channel catfish × blue catfish hybrids in two public put–grow–take fisheries in Kansas. North American Journal of Fisheries Management, 41(S1):S314–S322, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nesslage:2020:SBE

[NLW^{+}20] Geneviève M. Nesslage, Robert T. Leaf, Michael J. Wilberg, Raymond M. Mroch III, and Amy M. Schueller. A simulation-based evaluation of commercial port sampling programs for the Gulf and Atlantic menhaden fisheries. North American Journal of Fisheries Management, 40(4):995–1006, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nisbet:2021:UCD

[NMBM21] Mitchell T. Nisbet, Randall A. Myers, Greg R. Binion, and Dusty McDonald. Using creel data to evaluate blue and channel catfish aggregate harvest regulations for Texas reservoirs. North American Journal of Fisheries Management, 41(S1):S323–S333, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nguyen:2023:DDP

[NMS^{+}23] Erin Nguyen, Kevin B. Mayes, Ryan Smith, Joe Trungale, and Joshuah S. Perkin. The duality of drought: Pelagic- and benthicspawning stream fishes show opposing responses to drought in the southern Great Plains. North American Journal of Fisheries Management, 43(5):1276–1293, October 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nordberg:2021:HDP

[NMW^{+}21] Brittany J. Nordberg, Elizabeth G. Mandeville, Annika W. Walters, Jason C. Burckhardt, and Catherine E. Wagner. Historical data provide important context for understanding declines in cutthroat trout. North American Journal of Fisheries Management, 41(3):809–819, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Naman:2020:EVG

[NRJ^{+}20] Sean M. Naman, Jordan S. Rosenfeld, Eva Jordison, Melanie Kuzyk, and Brett C. Eaton. Exploitation of velocity gradients by sympatric stream salmonids: Basic insights and implications for instream flow management. North American Journal of Fisheries Management, 40(2):320–329, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Oswald:2020:SDM

[OBR^{+}20] Kenneth J. Oswald, Sophia Beery, Kalyn Rossiter, Yong Wang, and Marc R. Kibbey. Species distribution models for a native imperiled minnow and a nonnative sport fish in a Western Ohio River. North American Journal of Fisheries Management, 40(5): 1225–1238, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Oliver:2022:EET

[OMAA22] Devon C. Oliver, Ryan D. Mann, Caitlyn G. Aymami, and Lorraine D. Avenetti. An evaluation of the efficacy of a targeted supplemental stocking using a multistate capture–recapture model. North American Journal of Fisheries Management, 42(4):952– 965, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Osborne:2022:APO

[ORL^{+}22] Christopher A. Osborne, Jason M. Robinson, Brian F. Lantry, Brian C. Weidel, Ian Harding, and Michael J. Connerton. Accuracy and precision of otolith-derived age interpretations for known-age lake trout. North American Journal of Fisheries Management, 42(1):207–216, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Oliver:2021:ERI

[ORWS21] Devon C. Oliver, Neil P. Rude, Gregory W. Whitledge, and Daniel S. Stich. Evaluation of recently implemented harvest regulations in a data-limited catfish fishery with Bayesian estimation. North American Journal of Fisheries Management, 41(S1):S364– S378, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olsen:2020:SMS

[OS20] Noelle A. Olsen and Bradley G. Stevens. Size at maturity, shell conditions, and morphometric relationships of male and female Jonah crabs in the Middle Atlantic Bight. North American Journal of Fisheries Management, 40(6):1472–1485, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ohms:2023:FPT

[OTBB23] Haley A. Ohms, J. Vincent Tranquilli, Gabriel Brooks, and David A. Boughton. Flexible-pass-through antennas for halfduplex passive integrated transponder tags in medium to large rivers used to study size-selective reservoir passage. North American Journal of Fisheries Management, 43(3):687–694, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olsen:2022:DET

[OTW^{+}22] Zachary Olsen, Darin Topping, Tom Wagner, Joel Anderson, Carey Gelpi, Glen Sutton, and Nicole Carrillo. Differential effects of three crab trap escape ring sizes on retention of male and female blue crabs. North American Journal of Fisheries Management, 42(2):270–276, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Owens:2021:SFT

[Owe21] Stephen J. Owens. Supplemental feeding as a tool in the development of a quality wild rainbow trout fishery. North American Journal of Fisheries Management, 41(2):531–536, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pendleton:2021:LTT

[PA21] Richard M. Pendleton and Robert D. Adams. Long-term trends in juvenile Atlantic sturgeon abundance may signal recovery in the Hudson River, New York, USA. North American Journal of Fisheries Management, 41(4):1170–1181, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perry:2022:SMB

[PAC22] Harriet M. Perry, John Anderson, and Lillian Collins. Status and management of the blue crab fishery in the Gulf of Mexico. North American Journal of Fisheries Management, 42(1):164– 179, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perkin:2023:FRR

[PAE^{+}23] Joshuah S. Perkin, Matthew R. Acre, Johnathan K. Ellard, Anthony W. Rodger, Joeseph F. Trungale, Kirk O. Winemiller, and Lauren E. Yancy. Flow–recruitment relationships for shoal chub and implications for managing environmental flows. North American Journal of Fisheries Management, 43(5):1260–1275, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pan:2021:MEY

[Pan21] Minling Pan. Maximum economic yield and nonlinear catchability. North American Journal of Fisheries Management, 41(5):

1229–1245, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porreca:2022:FPR

[PBD^{+}22] Anthony P. Porreca, Kyle J. Broadway, Thomas M. Detmer, Dakota S. Radford, Carly C. Fenstermacher, and Joseph J. Parkos III. Fish and prey resources in reservoir coves with and without offshore artificial habitat structures. North American Journal of Fisheries Management, 42(4):966–976, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perkin:2023:AMF

[PBEK23] Joshuah S. Perkin, Shannon K. Brewer, Anthony A. Echelle, and Patrick M. Kočovský. Avoiding a macabre future for Macrhybopsis — a special section on improving management and conservation of chubs. North American Journal of Fisheries Management, 43(5):1145–1150, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parkos:2023:SVM

[PBK^{+}23] Joseph J. Parkos III, Steven E. Butler, Gregory D. King, Anthony P. Porreca, David P. Coulter, Ruairi MacNamara, and David H. Wahl. Spatiotemporal variation in the magnitude of reproduction by invasive, pelagically spawning carps in the Illinois waterway. North American Journal of Fisheries Management, 43 (1):112–125, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Polivka:2020:BRS

[PC20] Carlos M. Polivka and Shannon M. Claeson. Beyond redistribution: In-stream habitat restoration increases capacity for youngof-the-year Chinook salmon and steelhead in the Entiat River, Washington. North American Journal of Fisheries Management, 40(2):446–458, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paul:2021:ANP

[PCM^{+}21] Andrew J. Paul, Christopher L. Cahill, Laura MacPherson, Michael G. Sullivan, and Myles R. Brown. Are Alberta’s northern pike populations at risk from walleye recovery? North American Journal of Fisheries Management, 41(2):399–409, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2021:IPT

[PE21] Douglas P. Peterson and Rod O. Engle. Ingestion of PIT tags by hatchery-reared juvenile steelhead and subsequent detection at main-stem dams in the Columbia River Basin. North American Journal of Fisheries Management, 41(3):871–885, June 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Payton:2023:ECS

[PFGE23] Quinn Payton, Jeffrey Fryer, Tommy Garrison, and Allen F. Evans. Estimating cause-specific mortality and survival of juvenile fall Chinook salmon: an investigation of avian predation across large spatial scales. North American Journal of Fisheries Management, 43(2):569–585, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Putman:2020:UCA

[PG20] Nathan F. Putman and Benny J. Gallaway. Using common age units to communicate the relative catch of red snapper in recreational, commercial, and shrimp fisheries in the Gulf of Mexico. North American Journal of Fisheries Management, 40(1):232– 241, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parker:2023:EUP

[PHG^{+}23] Charles R. Parker, Kevin M. Hunt, Stephen C. Grado, James E. Henderson, and Matthew A. Freeman. Exploring the use of price indices to adjust angler expenditure profiles over time. North American Journal of Fisheries Management, 43(2):417–430, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pelekai:2023:EPL

[PHW^{+}23] K. N. Pelekai, J. E. Hess, L. A. Weitkamp, R. T. Lampman, and J. A. Miller. Evaluation of Pacific Lamprey statoliths for age estimation across their life cycle. North American Journal of Fisheries Management, 43(6):1610–1622, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peoples:2021:MDE

[PJD^{+}21] Brandon K. Peoples, Emily Judson, Tanya L. Darden, Daniel J. Farrae, Kevin Kubach, Jean Leitner, and Mark C. Scott. Modeling distribution of endemic Bartram’s bass Micropterus sp. cf. coosae: Disturbance and proximity to invasion source increase hybridization with invasive Alabama bass. North American Journal of Fisheries Management, 41(5):1309–1321, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2024:MGB

[PK24] Miles I. Peterson and Satoshi Kitano. Male guarding behavior and brood predators of invasive bluegill in a Japanese lake. North American Journal of Fisheries Management, 44(1):204– 215, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porath:2021:CCV

[PKNS21] Mark T. Porath, Thomas J. Kwak, Ben C. Neely, and Daniel E. Shoup. Catfish 2020, a clear vision of the future. North American Journal of Fisheries Management, 41(S1):S1–S10, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perkin:2023:WLF

[PKSG23] Joshuah S. Perkin, Patrick M. Kočovský, Zachary D. Steffensmeier, and Keith B. Gido. Why are larger fish farther upstream? Testing multiple hypotheses using silver chub in two Midwestern United States riverscapes. North American Journal of Fisheries Management, 43(5):1225–1245, October 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Poole:2020:BSI

[PKTZ20] Alex S. Poole, Todd M. Koel, Nathan A. Thomas, and Alexander V. Zale. Benthic suffocation of invasive lake trout embryos by fish carcasses and sedimentation in Yellowstone Lake. North American Journal of Fisheries Management, 40(5):1077–1086, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pothoven:2022:BAF

[PO22] Steven Pothoven and Chris Olds. Biotic and abiotic factors affecting feeding success of early juvenile lake whitefish in Lakes Michigan and Huron. North American Journal of Fisheries Management, 42(1):180–188, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pearsons:2024:CDS

[PO24] Todd N. Pearsons and Rolland R. O’Connor. Comparisons of donor stray percentages between hatchery- and natural-origin Chinook salmon and steelhead in the upper Columbia watershed. North American Journal of Fisheries Management, 44(2):276– 298, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Putt:2022:MEA

[PREBK22] Annika E. Putt, Daniel Ramos-Espinoza, Douglas C. Braun, and Josh Korman. Methods for estimating abundance and associated uncertainty from passive count technologies. North American Journal of Fisheries Management, 42(1):96–108, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Payton:2021:EAH

[PS21] Quinn Payton and Nicholas A. Som. Evaluating alternative hierarchical modeling approaches for the estimation of salmonid smolt abundance. North American Journal of Fisheries Management, 41(4):1182–1193, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porta:2020:EEE

[PSS20] Michael J. Porta, David R. Stewart, and Richard A. Snow. Estimating the effects of environmental variables and gear type on panfish catch and bycatch in reservoirs. North American Journal of Fisheries Management, 40(5):1174–1186, October 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Porta:2021:CTM

[PSS21] Michael J. Porta, Richard A. Snow, and Daniel E. Shoup. A comparison of two methods for sampling bluegill and redear sunfish in small impoundments. North American Journal of Fisheries Management, 41(1):196–203, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Petrosky:2020:APR

[PST^{+}20] Charles E. Petrosky, Howard A Schaller, Eric S. Tinus, Timothy Copeland, and Adam J. Storch. Achieving productivity to recover and restore Columbia River stream-type Chinook salmon relies on increasing smolt-to-adult survival. North American Journal of Fisheries Management, 40(3):789–803, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prull:2023:PBT

[PSW23] Marcus Prull, Martin A. Simonson, and Michael J. Weber. Precision and bias of three bigmouth buffalo aging structures. North American Journal of Fisheries Management, 43(3):646–655, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pritt:2022:ESV

[PTCZ22] Jeremy J. Pritt, Stephen M. Tyszko, Joseph D. Conroy, and Richard D. Zweifel. Evaluating single- versus chase-boat lowfrequency electrofishing for assessing flathead catfish and blue catfish populations in a large river. North American Journal of Fisheries Management, 42(4):915–925, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2023:ASL

[PTFZ23] Erin Peterson, Rex Thors, Danielle Frechette, and Joseph D. Zydlewski. Adult sea lamprey approach and passage at the Milford Dam fishway, Penobscot River, Maine, United States. North American Journal of Fisheries Management, 43(4):1052–1065, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Page:2021:OVA

[PWLH21] Kevin S. Page, Curtis P. Wagner, Edward Lewis, and R. Scott Hale. An online voluntary angler catch reporting program for managing muskellunge fisheries in Ohio: Trends in catch data and angler participation. North American Journal of Fisheries Management, 41(4):1194–1206, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quintana:2021:EAF

[QBA21] Yasmín Quintana, Christian A. Barrientos, and Micheal Allen. Evaluation of an artisanal freshwater fishery in Guatemala finds underfished conditions. North American Journal of Fisheries Management, 41(6):1731–1743, December 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radomski:2020:AMS

[RAB^{+}20] Paul Radomski, Charles S. Anderson, Richard E. Bruesewitz, Andrew J. Carlson, and Brian D. Borkholder. An assessment model for a standard gill net incorporating direct and indirect selectivity applied to walleye. North American Journal of Fisheries Management, 40(1):105–124, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rachels:2021:ELD

[Rac21] Kyle T. Rachels. Exploring legacy data sets to infer spatial and temporal trends in the ictalurid assemblage of an Atlantic Slope river. North American Journal of Fisheries Management, 41(S1): S195–S204, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ridgway:2023:SCH

[RAH^{+}23] Josey L. Ridgway, Matthew R. Acre, Tyler M. Hessler, Dustin W. Broaddus, Jessica Morris, and Robin D. Calfee. Silver carp herding: a telemetry evaluation of efficacy and implications for design and application. North American Journal of Fisheries Management, 43(6):1750–1764, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:2020:PIH

[RAHS20] H. Eve Robinson, Julie D. Alexander, Sascha L. Hallett, and Nicholas A. Som. Prevalence of infection in hatchery-origin Chinook salmon correlates with abundance of Ceratonova shasta spores: Implications for management and disease risk. North American Journal of Fisheries Management, 40(4):959–972, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roth:2020:PEE

[RBFQ20] Curtis J. Roth, Zachary S. Beard, Jon M. Flinders, and Michael C. Quist. Population ecology and evaluation of suppression scenarios for introduced Utah chub. North American Journal of Fisheries Management, 40(1):133–144, February 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ruggirello:2020:UUV

[RBFS20] Jack E. Ruggirello, Scott A. Bonar, Olin G. Feuerbacher, and Lee H. Simons. Use of underwater videography to quantify conditions utilized by endangered moapa dace while spawning. North American Journal of Fisheries Management, 40(1):17–28, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rehm:2020:EDR

[RCD20] Travis R. Rehm, Steve R. Chipps, and Jacob L. Davis. Effects of density reduction on age-specific growth of stream-dwelling brown trout. North American Journal of Fisheries Management, 40(6): 1355–1366, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reinhardt:2022:VSD

[RCD22] Leslie Reinhardt, Timothy Copeland, and Micah Davison. Validation of scale-derived ages in wild juvenile and adult steelhead using parental-based tagging. North American Journal of Fisheries Management, 42(2):260–269, April 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ryan:2021:CIW

[RCR21] Robert G. Ryan, Matthew P. Corsi, and Pete Rust. Characteristics of an introduced walleye population with implications for suppression. North American Journal of Fisheries Management, 41(6):1863–1877, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richer:2020:MFP

[RFK^{+}20] Eric E. Richer, Eric R. Fetherman, Elizabeth A. Krone, F. Boyd Wright III, and Matt C. Kondratieff. Multispecies fish passage evaluation at a rock-ramp fishway in a Colorado transition zone stream. North American Journal of Fisheries Management, 40 (6):1510–1522, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reid:2020:NRP

[RG20] Stewart B. Reid and Damon H. Goodman. Natural recolonization by Pacific lampreys in a Southern California coastal drainage: Implications for their biology and conservation. North American Journal of Fisheries Management, 40(2):335–341, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosinski:2023:NOD

[RGHF23] Caroline L. Rosinski, Jeff Glaid, Matt Hahn, and William W. Fetzer. Natal origin differentiation using eye lens stable isotope analysis. North American Journal of Fisheries Management, 43 (2):547–555, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ridgway:2020:DIS

[RGMH20] Josey L. Ridgway, Jason M. Goeckler, Jessica Morris, and Jeremy J. Hammen. Diel influences on silver carp catch rates using an electrified paupier in lentic habitats. North American Journal of Fisheries Management, 40(4):1023–1031, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Renik:2020:EBB

[RH20] Kathryn M. Renik and Andrew W. Hafs. Effect of beaver on brook trout habitat in North Shore, Lake Superior, streams. North American Journal of Fisheries Management, 40(2):427–445, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:2023:IMR

[RH23] Kevin B. Rogers and Brian W. Hodge. An inexpensive method for reliable recovery of stream temperature data. North American Journal of Fisheries Management, 43(5):1349–1358, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:2024:QSV

[RHFU24] Tyler J. Robinson, Andrew W. Hafs, Shannon J. Fisher, and Jeffrey S. Ueland. Quantifying seasonal variations in space use within the home range of burbot. North American Journal of Fisheries Management, 44(2):492–506, April 2024. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riepe:2023:IDB

[RHUM^{+}23] Tawni B. Riepe, Zachary E. Hooley-Underwood, Rebecca E. McDevitt, Alan Sralik, and Pete Cadmus. Increased density of bluehead sucker larvae decreases critical thermal maximum. North American Journal of Fisheries Management, 43(4):1135– 1142, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riggers:2022:PSA

[RJ22] Brian L. Riggers and Michelle K. Jones. Paradigm shift: Applying capture–recapture techniques to electronic licensing system data to estimate Chinook salmon harvest. North American Journal of Fisheries Management, 42(4):815–827, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richer:2022:GMG

[RKP^{+}22] Eric E. Richer, Matt C. Kondratieff, Greg Policky, Matt D. Robinson, Michael Atwood, and Madison R. Myers. From gold mining to gold medal fishery: Evaluating the fishery response to stream restoration on the Upper Arkansas River, Colorado. North American Journal of Fisheries Management, 42(1):24–36, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ridgway:2023:AFH

[RLS^{+}23] Josey L. Ridgway, Katelyn M. Lawson, Stephen A. Shier, Robin D. Calfee, and Duane C. Chapman. An assessment of fish herding techniques: Management implications for mass removal and control of silver carp. North American Journal of Fisheries Management, 43(1):176–188, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reilly:2021:PCZ

[RM21] Jessica R. Reilly and Joshua M. Miller. A phylogeographic contact zone for Arctic grayling in Alberta, Canada. North American Journal of Fisheries Management, 41(6):1619–1630, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Runde:2020:AFT

[RMB^{+}20] Brendan J. Runde, Théo Michelot, Nathan M. Bacheler, Kyle W. Shertzer, and Jeffrey A. Buckel. Assigning fates in telemetry studies using hidden Markov models: an application to deepwater groupers released with descender devices. North American Journal of Fisheries Management, 40(6):1417–1434, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roth:2022:TPT

[RMR^{+}22] Curtis J. Roth, Kevin A. Meyer, Ronald V. Roberts, Brock A. Lipple, Eric J. Stark, and Timothy Copeland. Temporal population trends and habitat associations for mountain whitefish in Central Idaho. North American Journal of Fisheries Management, 42(5):1226–1236, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ruoss:2023:RCH

[RMWK23] Jenna P. Ruoss, Brian E. Mason, Melissa R. Wuellner, and Keith D. Koupal. Resetting the clock: How a reconnection event reshaped cove fish assemblages in a Nebraska reservoir. North American Journal of Fisheries Management, 43(2):327–338, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rossi:2023:FAC

[ROP^{+}23] Gabriel J. Rossi, Mariska Obedzinski, Shelley Pneh, Sarah Nossaman Pierce, William T. Boucher, Weston M. Slaughter, Keane M. Flynn, and Theodore E. Grantham. Flow augmentation from off-channel storage improves salmonid habitat and survival. North American Journal of Fisheries Management, 43 (6):1772–1788, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roth:2023:CAP

[RPEF^{+}23] Daniel R. Roth, Jordan J. Pesik, Eden L. Effert-Fanta, David H. Wahl, and Robert E. Colombo. Comparison of active and passive larval sampling gears in monitoring reproduction of invasive bigheaded carps in large-river tributaries. North American Journal of Fisheries Management, 43(1):35–45, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rechisky:2021:ETW

[RPJ^{+}21] Erin L. Rechisky, Aswea D. Porter, Stephen D. Johnston, Christine F. Stevenson, Scott G. Hinch, Brian P. V. Hunt, and David W. Welch. Exposure time of wild, juvenile sockeye salmon to open-net-pen Atlantic salmon farms in British Columbia, Canada. North American Journal of Fisheries Management, 41 (3):650–660, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:2023:UCO

[RR23] Kerrick W. Robinson and Jacob M. Rash. Using catch orientations to identify clusters of North Carolina trout anglers. North American Journal of Fisheries Management, 43(4):984–999, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rudershausen:2020:EVD

[RRB20] P. J. Rudershausen, B. J. Runde, and J. A. Buckel. Effectiveness of venting and descender devices at increasing rates of postrelease survival of black sea bass. North American Journal of Fisheries Management, 40(1):125–132, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Runde:2022:DDR

[RRB^{+}22] Brendan J. Runde, Paul J. Rudershausen, Nathan M. Bacheler, Ryan M. Tharp, and Jeffrey A. Buckel. Depredation of demersal reef fishes released with descender devices is uncommon off North Carolina, USA. North American Journal of Fisheries Management, 42(5):1196–1201, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rogers:2021:CBT

[RRL^{+}21] Karli M. Rogers, Shawn M. Rummel, Kathleen M. Lavelle, Joseph E. Duchamp, Jonathan M. Niles, and David J. Janetski. A comparison of brook trout passage at road culverts to broadscale assessment criteria in a Pennsylvania headwater stream. North American Journal of Fisheries Management, 41(5):1351– 1359, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roth:2021:PDT

[RSK^{+}21] Curtis J. Roth, Eric J. Stark, Lucas D. Koenig, Brian S. Ayers, and Kevin A. Meyer. Population dynamics and temporal trends of bull trout in the East Fork Salmon River, Idaho. North American Journal of Fisheries Management, 41(2):455–465, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rudershausen:2023:EVR

[RSM^{+}23] Paul J. Rudershausen, Hailey M. Schmidt, Jeffery H. Merrell, Brendan J. Runde, and Jeffrey A. Buckel. Effectiveness of venting and recompression for increasing postrelease survival of barotraumatized black sea bass across a range of depths. North American Journal of Fisheries Management, 43(1):257–267, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roper:2023:RBD

[RSO23] Brett Roper, W. Carl Saunders, and Jeffrey V. Ojala. The relationship between disturbance events and substantial changes in stream conditions on public lands in the inland Pacific Northwest. North American Journal of Fisheries Management, 43(1): 268–290, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rodger:2021:SDH

[RWS^{+}21] Anthony W. Rodger, Skylar L. Wolf, Trevor A. Starks, Jim P. Burroughs, and Shannon K. Brewer. Seasonal diet and habitat use of large, introduced rainbow trout in an Ozark highland stream. North American Journal of Fisheries Management, 41 (6):1764–1780, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rackovan:2021:EIT

[RZA^{+}21] Jenna Rackovan, Joe Zottoli, Stephen V. Amaral, Jonathan Black, and Brian McMahon. Effect of intermittent traveling water screen operation on fish survival. North American Journal of Fisheries Management, 41(6):1714–1730, December 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rutherford:2021:PEB

[RZK^{+}21] Edward S. Rutherford, Hongyan Zhang, Yu-Chun Kao, Doran M. Mason, Ali Shakoor, Keith Bouma-Gregson, Jason T. Breck, David M. Lodge, and W. Lindsay Chadderton. Potential effects of bigheaded carps on four Laurentian Great Lakes food webs. North American Journal of Fisheries Management, 41(4):999– 1019, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sakaris:2021:CTO

[SB21a] Peter C. Sakaris and Timothy F. Bonvechio. Comparison of two otolith processing methods for estimating age of three catfish species. North American Journal of Fisheries Management, 41 (S1):S428–S439, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shoup:2021:ESD

[SB21b] Daniel E. Shoup and Kristopher A. Bodine. Effect of sample duration on catch rate and size structure data for blue catfish collected by low-frequency electrofishing. North American Journal of Fisheries Management, 41(S1):S400–S405, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2023:PEA

[SBB23] Kevin M. Sullivan, Michael M. Bailey, and David L. Berlinsky. Passage efficiency of alewife in a Denil fishway using passive integrated transponder tags. North American Journal of Fisheries Management, 43(3):772–785, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2020:HCR

[SBD^{+}20] Nathan G. Smith, David L. Buckmeier, Daniel J. Daugherty, Daniel L. Bennett, Peter C. Sakaris, and Clinton R. Robertson. Hydrologic correlates of reproductive success in the alligator gar. North American Journal of Fisheries Management, 40(3):595– 606, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2023:SVF

[SBF^{+}23] Nathan G. Smith, David L. Buckmeier, B. Paul Fleming, Archis R. Grubh, Michael D. Homer, Jr., and Sarah M. Robertson. Spatial variability in the fish assemblage of a large river– reservoir ecosystem. North American Journal of Fisheries Management, 43(2):313–326, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2023:OSS

[SBH^{+}23] David R. Stewart, Joseph C. Barron II, Tasha Harden, Elizabeth R. Grube, Manuel Ulibarri, Andrew T. Taylor, Edward J. Heist, Elizabeth Mandeville, Brian T. Hickerson, Nathan Berg, Lacrecia A. Johnson, Matthew J. Butler, and Grant M. Harris. The optimal stocking strategy for Yaqui catfish. North American Journal of Fisheries Management, 43(5):1407–1426, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffensmeier:2023:CNN

[SBW^{+}23] Zachary D. Steffensmeier, Shannon K. Brewer, Maeghen Wedgeworth, Trevor A. Starks, Anthony W. Rodger, Erin Nguyen, and Joshuah S. Perkin. Conservation at the nexus of niches: Multidimensional niche modeling to improve management of Prairie Chub. North American Journal of Fisheries Management, 43(5): 1205–1224, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spanos:2023:EVU

[SCDS23] Mary C. Spanos, Curry J. Cunningham, Katie A. Drew, and Trent M. Sutton. Evaluating the viability of the use of T-bar and radiotelemetry tags on prespawn Arctic Lampreys. North American Journal of Fisheries Management, 43(6):1631–1647, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sweezey:2020:EDM

[SCL^{+}20] B. B. Sweezey, C. W. Capizzano, J. A. Langan, H. P. Benoît, E. W. Hutchins II, J. W. Mandelman, W. Y. Koh, M. J. Dean, B. N. Anderson, and J. A. Sulikowski. Estimating the discard mortality of Atlantic cod in the Southern Gulf of Maine commercial lobster fishery. North American Journal of Fisheries Management, 40(5):1252–1262, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schumann:2020:OCO

[SCMJF20] David A. Schumann, Michael E. Colvin, Leandro E. Miranda, and D. Todd Jones-Farrand. Occurrence and co-occurrence patterns of gar in river–floodplain habitats: Methods to leverage species coexistence to benefit distributional models. North American Journal of Fisheries Management, 40(3):622–637, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepard:2021:ERS

[SCN^{+}21] Bradley B. Shepard, Patrick Clancey, M. Lee Nelson, Carter G. Kruse, Robert Al-Chokhachy, Daniel Drinan, Mark L. Taper, and Alexander V. Zale. Evaluation of remote site incubators to incubate wild- and hatchery-origin westslope cutthroat trout embryos. North American Journal of Fisheries Management, 41(3): 844–855, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sinnatamby:2023:NES

[SCP^{+}23] R. Niloshini Sinnatamby, Ariane Cantin, Andrew J. Paul, Jennifer Earle, and John R. Post. No evidence of sustained recovery of native trout in response to angling suppression of invasive brook trout. North American Journal of Fisheries Management, 43(5): 1294–1309, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shertzer:2024:FMS

[SCW^{+}24] Kyle Shertzer, Scott Crosson, Erik Williams, Jie Cao, Rick De-Victor, Chris Dumas, and Geneviève Nesslage. Fishery management strategies for red snapper in the southeastern U.S. Atlantic: a spatial population model to compare approaches. North American Journal of Fisheries Management, 44(1):113–131, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2020:ACM

[SDB^{+}20] N. G. Smith, D. J. Daugherty, E. L. Brinkman, M. G. Wegener, B. R. Kreiser, A. M. Ferrara, K. D. Kimmel, and S. R. David. Advances in conservation and management of the alligator gar: a synthesis of current knowledge and introduction to a special section. North American Journal of Fisheries Management, 40 (3):527–543, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scyphers:2021:UEA

[SDF^{+}21] Steven B. Scyphers, J. Marcus Drymon, Kelsi L. Furman, Elizabeth Conley, Yvette Niwa, Amanda E. Jefferson, and Gregory W. Stunz. Understanding and enhancing angler satisfaction with fisheries management: Insights from the “Great Red Snapper Count”. North American Journal of Fisheries Management, 41 (3):559–569, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2023:ABA

[SDL^{+}23] Steven M. Sammons, Lawrence G. Dorsey, C. Scott Loftis, Preston Chrisman, Mark Scott, John Hammonds, Michael Jolley, Hunter Hatcher, John Odenkirk, John Damer, Matthew R. Lewis, and Eric J. Peatman. Alabama bass alter reservoir black bass species assemblages when introduced outside their native range.

North American Journal of Fisheries Management, 43(2):384– 399, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2024:AEE

[SDR24] Steven M. Sammons, Jason E. Dattilo, and Steven J. Rider. Assessing electrofishing effort required to estimate species richness of fishes in nonwadeable rivers of Alabama, USA. North American Journal of Fisheries Management, 44(1):39–56, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2021:SVF

[SDWM21] David R. Smith, Michael A. Dance, Joe West, and Stephen R. Midway. Spatiotemporal variability of fishery-dependent indices for the declining Louisiana southern flounder fishery. North American Journal of Fisheries Management, 41(6):1826–1837, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stratton:2021:APS

[SFFW21] M. E. Stratton, H. Finkle, J. A. Falke, and P. A. H. Westley. Assessing potential stock structure of adult Coho salmon in a small Alaska watershed: Quantifying run timing, spawning locations, and holding areas with radiotelemetry. North American Journal of Fisheries Management, 41(5):1423–1435, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swarr:2023:DCP

[SFM23] Tyler R. Swarr, Ryan M. Fitzpatrick, and Christopher A. Myrick. Design, construction, and preliminary hydraulic evaluation of a model rock ramp fishway. North American Journal of Fisheries Management, 43(4):935–946, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2021:PIT

[SG21] Charles G. Snow and Benjamin J. Goodman. Passive integrated transponder tag monitoring reveals complex migration and life history patterns in mountain whitefish in Upper Columbia River tributaries. North American Journal of Fisheries Management, 41(6):1691–1702, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sliger:2022:FDD

[SG22] Ridge Sliger and Gary D. Grossman. Foraging dynamics of a domesticated strain of brook trout. North American Journal of Fisheries Management, 42(5):1340–1348, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Somers:2021:CNP

[SGH^{+}21] Christopher M. Somers, Una Goncin, Shayna Hamilton, Michelle Chupik, and Ryan Fisher. Chasing northern pike under ice: Long-distance movements following catch-and-release ice angling. North American Journal of Fisheries Management, 41(5):1341– 1350, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Siemiantkowski:2022:CAT

[SGK^{+}22] Michael J. Siemiantkowski, Christopher S. Guy, Todd M. Koel, Lusha M. Tronstad, Carter R. Fredenberg, and Leo R. Rosenthal. Combination of acoustic telemetry and side-scan sonar advances suppression efforts for invasive lake trout in a submontane lake. North American Journal of Fisheries Management, 42(6): 1609–1622, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swanson:2020:OMC

[SGM^{+}20] Reid G. Swanson, Joel E. Gagnon, Loren M. Miller, Justine D. Dauphinais, and Peter W. Sorensen. Otolith microchemistry of common carp reflects capture location and differentiates nurseries in an interconnected lake system of the North American Midwest. North American Journal of Fisheries Management, 40(5):1100– 1118, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stell:2020:ALC

[SHFP20] E. Stell, J. J. Hoover, L. Fuller, and G. R. Parsons. Analyzing leap characteristics and water escape velocities of silver carp using in situ video analysis. North American Journal of Fisheries Management, 40(1):163–174, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sheffer:2022:MRM

[SHI22] Robert J. Sheffer, Steven R. Hogler, and Daniel A. Isermann. Mark–recapture models accurately predict growth trajectories of known-age muskellunge in Green Bay, Lake Michigan. North American Journal of Fisheries Management, 42(2):410–424, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2022:DPC

[SHWW22] Dustin M. Smith, Corbin D. Hilling, Stuart A. Welsh, and David I. Wellman, Jr. Differences in population characteristics and modeled response to harvest regulations in reestablished Appalachian walleye populations. North American Journal of Fisheries Management, 42(3):612–629, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sindt:2021:CCC

[Sin21] Anthony R. Sindt. Comparing channel catfish populations in the fragmented Minnesota River. North American Journal of Fisheries Management, 41(S1):S146–S155, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2023:ORO

[SK23] Steve M. Sammons and Rebecca M. Krogman. Old reservoirs, old and new challenges: Introduction to the 2021 Reservoir Symposium. North American Journal of Fisheries Management, 43(2): 287–289, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

See:2021:SSM

[SKA21] Kevin E. See, Ryan N. Kinzer, and Michael W. Ackerman. Statespace model to estimate salmon escapement using multiple data sources. North American Journal of Fisheries Management, 41 (5):1360–1374, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shaffer:2024:CED

[SKBB24] Jason T. Shaffer, Andrew P. Kinziger, Eric P. Bjorkstedt, and Andre Buchheister. Comparison of environmental DNA and underwater visual count surveys for detecting juvenile Coho salmon in small rivers. North American Journal of Fisheries Management, 44(2):449–462, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spurgeon:2022:TRS

[SKGL22] Jonathan Spurgeon, Joseph Kaiser, Christy Graham, and Steve Lochmann. Trout responses to stocking rates and river discharge within a southeastern U.S. hydropeaking tailwater. North American Journal of Fisheries Management, 42(4):926–938, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schriener:2023:AHA

[SKWR23] William Schriener, Keith Koupal, Melissa Wuellner, and Dustin H. Ranglack. Assessing habitat availability in a hydrodynamic reservoir system. North American Journal of Fisheries Management, 43(2):304–312, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schall:2021:PDS

[SL21] Benjamin J. Schall and David O. Lucchesi. Population dynamics and simulated effects of length-based trophy regulations for flathead and channel catfish in the Lower James River, South Dakota. North American Journal of Fisheries Management, 41(S1):S277– S292, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Staples:2023:AAD

[SLD^{+}23] Amber Staples, Henry D. Legett, Jessica L. Deichmann, Keira Heggie, and Matthew B. Ogburn. Automated acoustic detection of river herring (alewife and blueback herring) spawning activity. North American Journal of Fisheries Management, 43(3):869– 881, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2021:EAE

[SLE21] Troy W. Smith, Brad W. Liermann, and Lisa A. Eby. Ecological assessment and evaluation of potential eradication approaches for introduced redside shiners in a montane lake. North American Journal of Fisheries Management, 41(5):1473–1489, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shaw:2023:CRA

[SLG^{+}23] Stephanie L. Shaw, Zachary Lawson, Joseph Gerbyshak, Nathan Nye, and Michael Donofrio. Catch-and-release angling effects on lake sturgeon in Wisconsin, USA. North American Journal of Fisheries Management, 43(2):451–464, April 2023. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shuai:2021:SDP

[SLL^{+}21] Fangmin Shuai, Qianfu Liu, Yaqiu Liu, Zhi Wu, Sovan Lek, and Yushun Chen. Spatial distribution patterns of Japanese eels Anguilla japonica in a large subtropical river (Pearl River), China. North American Journal of Fisheries Management, 41(4):929– 938, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2020:MOA

[SLP20] Richard A. Snow, James M. Long, and Michael J. Porta. Marking otoliths of alligator gar by immersion in oxytetracycline. North American Journal of Fisheries Management, 40(3):669–674, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sun:2020:MDL

[SLZ^{+}20] Ming Sun, Yunzhou Li, Chongliang Zhang, Binduo Xu, Yiping Ren, and Yong Chen. Management of data-limited fisheries: Identifying informative data to achieve sustainable fisheries. North American Journal of Fisheries Management, 40(3):733–751, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stott:2022:ECS

[SM22] Nathan Stott and Jeffrey Miner. Environmental cues of spawning migration into a confined wetland by northern pike and common carp in Lake Erie: Identifying fine-scale patterns. North American Journal of Fisheries Management, 42(2):239–249, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Swain-Menzel:2023:LLI

[SMB23] Hannah N. Swain-Menzel and Eric J. Billman. Low-level infection of parasitic copepods on rainbow trout does not affect vulnerability to angling or short-term survival following catch-and-release angling. North American Journal of Fisheries Management, 43 (6):1789–1798, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schaick:2023:BCS

[SMCEF^{+}23] Samuel J. Schaick, Cassi J. Moody-Carpenter, Eden L. Effert-Fanta, Kellie N. Hanser, Daniel R. Roth, and Robert E. Colombo. Bigheaded carp spatial reproductive dynamics in Illinois and Wabash River tributaries. North American Journal of Fisheries Management, 43(1):101–111, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Strussmann:2020:NEM

[SMM20] Carlos Augusto Strüssmann, Kaho Miyoshi, and Shota Mitsui. A novel, efficient method for otolith specimen preparation using UV-Cured resins. North American Journal of Fisheries Management, 40(5):1187–1194, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Slongo:2022:CFR

[SMM^{+}22] Stephen A. Slongo, Brian E. McLaren, Robert W. Mackereth, Sohail Siddiqui, Raymond Tyhuis, Daniel Gibson, and Michael D. Rennie. Characterizing the flow regime in brook trout incubation habitats and implications for management in a hydro-regulated river. North American Journal of Fisheries Management, 42(5): 1097–1110, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2021:REF

[SMP^{+}21] Geoffrey D. Smith, Danielle L. Massie, Joseph Perillo, Tyler Wagner, and Daryl Pierce. Range expansion and factors affecting abundance of invasive flathead catfish in the Delaware and Susquehanna Rivers, Pennsylvania, USA. North American Journal of Fisheries Management, 41(S1):S205–S220, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scarnecchia:2021:DMI

[SNLD21] Dennis L. Scarnecchia, Cody J. Nagel, Youngtaik Lim, and Steven R. Dalbey. Demographics and management of the isolated Upper Missouri River paddlefish stock of Fort Peck Reservoir, Montana. North American Journal of Fisheries Management, 41 (6):1903–1919, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smokorowski:2021:WRL

[SPGP21] Karen E. Smokorowski, Jennie L. Pearce, William D. Geiling, and Thomas C. Pratt. Wood removals from lakes may not necessarily elicit fish population responses. North American Journal of Fisheries Management, 41(1):142–157, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepta:2024:LBI

[SPM^{+}24] Erin Shepta, Joshuah Perkin, Kevin B. Mayes, Monica E. Mc-Garrity, Christopher M. Schalk, and Carmen G. Montaña. Livebait industry as a pathway for movement of nonnative and invasive species: Implications for conservation of native Texas fishes. North American Journal of Fisheries Management, 44(2):394– 406, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shane:2021:SDP

[SPP^{+}21] Keith D. Shane, Kevin S. Page, Jeremy J. Pritt, Joseph D. Conroy, and Elizabeth A. Marschall. Season and discharge predict downstream emigration rates for reservoir sport fish populations. North American Journal of Fisheries Management, 41(6):1798– 1811, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spurgeon:2020:RBA

[SPSH20] Jonathan J. Spurgeon, Mark A. Pegg, Stephen F. Siddons, and Henry H. Hansen. Retention of T-bar anchor tags for channel catfish in the Red River of the North. North American Journal of Fisheries Management, 40(2):330–334, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stone:2020:EDW

[SPY^{+}20] Dennis M. Stone, Michael J. Pillow, Kirk L. Young, David R. Van Haverbeke, and James D. Walters. Effects of disparate water temperatures and food bases on humpback chub growth rates within the Little Colorado River, Arizona. North American Journal of Fisheries Management, 40(2):475–487, April 2020. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Starks:2020:OSB

[SR20] Trevor A. Starks and Anthony W. Rodger. Otolith and scalebased growth standards for lotic smallmouth bass. North American Journal of Fisheries Management, 40(4):986–994, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sass:2023:SHL

[SSF^{+}23] Greg G. Sass, Stephanie L. Shaw, Carly C. Fenstermacher, Anthony P. Porreca, and Joseph J. Parkos III. Structural habitat in lakes and reservoirs: Physical and biological considerations for implementation. North American Journal of Fisheries Management, 43(2):290–303, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skolte:2022:ESL

[SSH22] Alicia Skolte, Casey W. Schoenebeck, and Andrew W. Hafs. Effects of a shallow lake condition shift on habitat, zooplankton, and yellow perch dynamics. North American Journal of Fisheries Management, 42(3):659–667, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2024:SSW

[SSM^{+}24] David R. Smith, J. Warren Schlechte, Randall A. Myers, Michael A. Dance, Jacob D. Norman, and Mitchell T. Nisbet. Seasonal, spatial, and water level predictors of angler catch and effort within Texas black bass reservoir fisheries. North American Journal of Fisheries Management, 44(1):79–92, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2020:FEA

[SSPL20] Richard A. Snow, David R. Stewart, Michael J. Porta, and James M. Long. Feeding ecology of age-0 gar at Texoma Reservoir inferred from analysis of stable isotopes. North American Journal of Fisheries Management, 40(3):638–650, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2024:SFM

[SSR^{+}24] Quinnlan C. Smith, Stephanie L. Shaw, Joshua K. Raabe, Greg G. Sass, and Thomas R. Hrabik. Sport fish movement responses to a littoral zone coarse woody habitat addition. North American Journal of Fisheries Management, 44(2):463–491, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sass:2022:PAD

[SSS^{+}22] Greg G. Sass, Stephanie L. Shaw, Logan W. Sikora, Matthew Lorenzoni, and Mark Luehring. Plasticity in abundance and demographic responses of walleye to elevated exploitation in a north temperate lake. North American Journal of Fisheries Management, 42(3):535–548, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2022:MPR

[SSSP22] Richard S. Snow, David R. Stewart, Nathan G. Smith, and Michael J. Porta. Modeling the population response of alligator gar in Texoma Reservoir to harvest and discard mortality. North American Journal of Fisheries Management, 42(6):1635–1652, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schlechte:2021:IPA

[STB^{+}21] J. Warren Schlechte, John B. Taylor, David L. Buckmeier, Clifford P. Hutt, and Kevin M. Hunt. Identifying potential anglers and customer segments of Texas catfish anglers to guide management actions. North American Journal of Fisheries Management, 41(S1):S345–S363, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shoup:2021:UGD

[STP^{+}21] Daniel E. Shoup, Jeff B. Tibbits, Clayton P. Porter, Richard A. Snow, and Michael J. Porta. Use of gizzard diameter to estimate weight and length of partly digested gizzard shad in piscivore diets. North American Journal of Fisheries Management, 41(1): 252–258, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sikora:2021:FCC

[SVS^{+}21] Logan W. Sikora, Justin A. VanDeHey, Greg G. Sass, Gregory Matzke, and Michael Preul. Fish community changes associated with bullhead removals in four Northern Wisconsin lakes. North American Journal of Fisheries Management, 41(S1):S71–S81, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sylvia:2022:ASD

[SW22] Andrea Sylvia and Michael J. Weber. Assessing size-dependent population-level effects of largemouth bass tournament mortality. North American Journal of Fisheries Management, 42(5):1324– 1339, October 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stegens:2024:MRS

[SWA^{+}24] Erica Stegens, Daniel D. Wiegmann, Lisa M. Angeloni, Jeffrey R. Baylis, Robert A. S. Laroche, Steven P. Newman, Scott P. Egan, Greg G. Sass, and Kelly L. Weinersmith. Mark–recapture surveys affect nest site fidelity but not reproductive timing of male smallmouth bass. North American Journal of Fisheries Management, 44(1):132–144, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sowards:2021:UNA

[SWJ^{+}21] Bryan Sowards, Micah Waters, Craig Johnson, Ben C. Neely, and Daniel E. Shoup. Use of the North American standard gill net for sampling the invasive white perch: Information from three Kansas reservoirs. North American Journal of Fisheries Management, 41 (4):1207–1214, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simonson:2022:HEC

[SWM22] Martin A. Simonson, Michael J. Weber, and Audrey McCombs. Hyperstability in electrofishing catch rates of common carp and bigmouth buffalo. North American Journal of Fisheries Management, 42(2):425–437, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skalski:2021:PSJ

[SWTH21] John R. Skalski, Steven L. Whitlock, Richard L. Townsend, and Ryan A. Harnish. Passage and survival of juvenile salmonid smolts through dams in the Columbia and Snake Rivers, 2010– 2018. North American Journal of Fisheries Management, 41(3): 678–696, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Samuel:2024:VMN

[SYHF24] William T. Samuel, Lauren E. Yancy, Elizabeth G. Hinkle, and Jeffrey A. Falke. Validating morphometrics as a nonlethal tool to determine Arctic grayling sex. North American Journal of Fisheries Management, 44(1):70–78, February 2024. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trudeau:2022:SFT

[TBG^{+}22] Ashley Trudeau, Ben Beardmore, Gretchen A. Gerrish, Greg G. Sass, and Olaf P. Jensen. Social fish-tancing in Wisconsin: The effects of the COVID-19 pandemic on statewide license sales and fishing effort in northern inland lakes. North American Journal of Fisheries Management, 42(6):1530–1540, December 2022. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

TenHarmsel:2021:PCN

[TBIJ21] Hailey J. TenHarmsel, B. Bynum Boley, Brian J. Irwin, and Cecil A. Jennings. Perceived constraints and negotiations to trout fishing in Georgia based on angler specialization level. North American Journal of Fisheries Management, 41(1):115– 129, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Turner:2021:EEF

[TCB^{+}21] Sara M. Turner, Bradford C. Chase, Michael S. Bednarski, Scott P. Elzey, and Matt H. Ayer. Evaluating the effectiveness of fyke-net sampling to characterize an anadromous rainbow smelt population in the Fore River, Massachusetts, USA. North American Journal of Fisheries Management, 41(6):1789–1797, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trahan:2021:DCB

[TCL^{+}21] Alexandria T. Trahan, Auston D. Chhor, Michael J. Lawrence, Jacob W. Brownscombe, Daniel M. Glassman, Connor H. Reid, Alice E. I. Abrams, Andy J. Danylchuk, and Steven J. Cooke. Do carbonated beverages reduce bleeding from gill injuries in angled northern pike? North American Journal of Fisheries Management, 41(3):639–649, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2020:MBT

[TCM20] Timothy N. Taylor, Benjamin K. Cross, and Barry C. Moore. Modeling brook trout carrying capacity in Owhi Lake, Washington, using bioenergetics. North American Journal of Fisheries Management, 40(1):84–104, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2022:ASM

[TJ22] Matthew D. Taylor and Daniel D. Johnson. Adaptive spatial management to deal with postflood inshore bycatch in a penaeid trawl fishery. North American Journal of Fisheries Management, 42(2):334–342, April 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2020:HPG

[TLSP20] Andrew T. Taylor, James M. Long, Richard A. Snow, and Michael J. Porta. Hybridization and population genetics of alligator gar in Lake Texoma. North American Journal of Fisheries Management, 40(3):544–554, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2021:EPS

[TM21] Paul D. Thompson and Matthew D. McKell. Evaluating a permeable-style fish barrier to prevent the upstream recolonization of nonnative trout during an eradication project. North American Journal of Fisheries Management, 41(2):537–544, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2023:MPD

[TO23] Jamie Nicholas Thompson and Nathanael Christian Overman. Multiscale population demographics and growth responses of trout in the upper Snoqualmie River watershed, Washington. North American Journal of Fisheries Management, 43(4):1111– 1134, August 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Turecek:2021:RRF

[TPA^{+}21] Aaron Turecek, Quinn Payton, Julie D. Alexander, Damon Goodman, Allen F. Evans, and Nicholas A. Som. Reducing river flows to control a parasitic salmonid disease in the Klamath River: Simulations question the efficacy of desiccation as a management tool. North American Journal of Fisheries Management, 41(4):1215– 1224, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyszko:2021:ECC

[TPC21a] Stephen M. Tyszko, Jeremy J. Pritt, and Joseph D. Conroy. Explaining channel catfish population characteristics in Ohio reservoirs. North American Journal of Fisheries Management, 41(S1): S334–S344, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyszko:2021:IRC

[TPC21b] Stephen M. Tyszko, Jeremy J. Pritt, and Joseph D. Conroy. Indexing reservoir channel catfish population density and size structure with tandem, baited hoop nets. North American Journal of Fisheries Management, 41(S1):S406–S414, October 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2022:MRM

[TPC^{+}22] Andrew T. Taylor, Alexis M. Peeper, Binod Chapagain, Omkar Joshi, and James M. Long. Modern reporting methods for angler tag-return studies: Trends in data quality, choice of method, and future considerations. North American Journal of Fisheries Management, 42(1):189–199, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2020:SAF

[TPPC20] Brandon C. Thompson, Wesley F. Porak, William F. Pouder, and Edward V. Camp. Survival of advanced-fingerlings of Florida largemouth bass stocked in small Florida lakes. North American Journal of Fisheries Management, 40(6):1532–1544, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:2021:PSF

[TRB^{+}21] Kenneth F. Tiffan, Tobyn N. Rhodes, Brad K. Bickford, Dalton D. Lebeda, William P. Connor, and Frank L. Mullins. Performance of subyearling fall Chinook salmon tagged with 8-, 9-, and 12-mm passive integrated transponder tags in the Snake River. North American Journal of Fisheries Management, 41(1):176– 186, February 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2020:ISM

[TSBJ20] Matthew D. Taylor, Jack Silburn, David J. Booth, and Daniel D. Johnson. Impact of spatial management on nontarget species in an oceanic penaeid trawl fishery. North American Journal of Fisheries Management, 40(2):509–520, April 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tuohy:2020:MCF

[TSJ20] Adrian M. Tuohy, John R. Skalski, and Aaron T. Jorgenson. Modified commercial fish trap to help eliminate salmonid bycatch mortality. North American Journal of Fisheries Management, 40(5): 1239–1251, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2021:UAV

[TVG^{+}21] Paul D. Thompson, Emanuel A. Vasquez, Ian Gowing, Tim Edgar, Ann Neville, and Allison Jones. Unmanned aerial vehicle technology proves an effective and efficient technique for identifying critical native fish habitat. North American Journal of Fisheries Management, 41(3):616–625, June 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Utrup:2023:BIE

[UHB^{+}23] Brad E. Utrup, Jan-Michael Hessenauer, Andrew S. Briggs, Kim T. Scribner, Jeannette Kanefsky, and Todd C. Wills. Biological investigation of the endangered northern madtom in the North Channel St. Clair River. North American Journal of Fisheries Management, 43(3):730–742, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vivian:2021:ZCR

[VF21] Mary K. Vivian and Daniel Frazer. Zooplankton community response to the introduction of cisco in Tiber Reservoir, Montana. North American Journal of Fisheries Management, 41(6): 1838–1849, December 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

vanPoorten:2020:UDA

[vPM20] Brett T. van Poorten and Cameron J. A. MacKenzie. Using decision analysis to balance angler utility and conservation in a recreational fishery. North American Journal of Fisheries Management, 40(1):29–47, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wildhaber:2023:LSP

[WA23] Mark L. Wildhaber and Janice L. Albers. Laboratory studies of potential competition for food and substrate among early juvenile Missouri River sturgeon and sympatric chub species. North American Journal of Fisheries Management, 43(5):1190–1204, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williford:2021:PPS

[WAKO21] Damon Williford, Joel Anderson, William Karel, and Zachary Olsen. Phylogeography, population structure, and historical demography of black drum in North America. North American Journal of Fisheries Management, 41(4):1020–1039, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wolf:2021:SMP

[WB21] Skylar L. Wolf and Shannon K. Brewer. Survival and movement patterns of rainbow trout stocked in a groundwater-influenced warmwater stream. North American Journal of Fisheries Management, 41(3):600–615, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilson:2024:WET

[WB24] Jeremy T. Wilson and Thomas W. Buehrens. Weirs: an effective tool to reduce hatchery–wild interactions on the spawning grounds? North American Journal of Fisheries Management, 44 (1):21–38, February 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilberg:2022:EHP

[WCSL22] Michael J. Wilberg, Sara Coleman, H. Ward Slacum, and Dong Liang. Efficiency of hydraulic patent tongs for surveying restored eastern oyster reefs in Harris Creek, Maryland. North American Journal of Fisheries Management, 42(4):1011–1019, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watkins:2021:ALB

[WD21] Carson J. Watkins and Andrew M. Dux. Assessment of largemouth bass growth in Northern Idaho lakes: Spatiotemporal patterns and extrinsic associations. North American Journal of Fisheries Management, 41(1):86–102, February 2021. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Werner:2022:SVS

[WDPH22] Jacob P. Werner, Quintin J. Dean, Mark A. Pegg, and Martin J. Hamel. Spatial variability of silver carp population demographics in a large tributary river. North American Journal of Fisheries Management, 42(4):882–892, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2022:QSS

[WGBK22] Jacob R. Williams, Christopher S. Guy, Patricia E. Bigelow, and Todd M. Koel. Quantifying the spatial structure of invasive lake trout in Yellowstone Lake to improve suppression efficacy. North American Journal of Fisheries Management, 42(1):50–62, February 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2020:TAT

[WGKB20] Jacob R. Williams, Christopher S. Guy, Todd M. Koel, and Patricia E. Bigelow. Targeting aggregations of telemetered lake trout to increase gillnetting suppression efficacy. North American Journal of Fisheries Management, 40(1):225–231, February 2020. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williford:2023:EDC

[WHA23] Damon Williford, Polly Hajovsky, and Joel Anderson. Environmental DNA compliments traditional sampling for monitoring fish communities in a Texas estuary. North American Journal of Fisheries Management, 43(5):1372–1394, October 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weissman:2021:DDM

[WKM^{+}21] Amelia Weissman, Ryan Knotek, John Mandelman, David Rudders, Sally Roman, and James Sulikowski. Determining discard mortality of monkfish in a sea scallop dredge fishery. North American Journal of Fisheries Management, 41(3):856–870, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winders:2021:SSE

[WM21] Kyle Robert Winders and Joseph A. McMullen. Size-specific exploitation of flathead catfish and blue catfish by recreational and commercial fishers in the Missouri and Mississippi Rivers, Missouri. North American Journal of Fisheries Management, 41(S1): S247–S254, October 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wedgeworth:2023:VPC

[WMB23] Maeghen Wedgeworth, Robert Mollenhauer, and Shannon K. Brewer. Variation in prairie chub hatch relationships across wet and dry years in the upper Red River basin. North American Journal of Fisheries Management, 43(5):1246–1259, October 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whitman:2024:FGI

[WMJ^{+}24] George Whitman, Jordan Messner, Rachel C. Johnson, Malte Willmes, Brian P. Kennedy, and Carson Jeffres. Forensic geochemistry identifies the illegal introduction of walleye into Lake Cascade, Idaho. North American Journal of Fisheries Management, 44(2):407–414, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Webster:2021:NGA

[WPGC21] N. Webster, S. Price, C. Gowan, and S. S. Coster. A nonlethal genetic assay for discriminating sex in brook trout. North American Journal of Fisheries Management, 41(4):1120–1123, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic). See erratum [Ano22a].

White:2023:LSF

[WPM^{+}23] Michael S. White, Paul H. Patrick, Elaine Mason, Nichole Wiemann, and Joe Tetreault. Laboratory studies of fish entrainment sampling using a screw centrifugal impeller pump. North American Journal of Fisheries Management, 43(3):882–888, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whitmore:2021:HBR

[WRH^{+}21] Margaret Whitmore, Shannon Richardson, Austin Huff, Kevin Goodson, Thomas P. Quinn, Andrew H. Dittman, Marc A. Johnson, Maryam Kamran, and David L. G. Noakes. Homeward bound: In-river movements of adult hatchery- and natural-origin Chinook salmon in the Elk River, Oregon. North American Journal of Fisheries Management, 41(4):1088–1096, August 2021.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watson:2023:JAP

[WSG^{+}23] Sterling M. Watson, Abraham D. Schneider, Lynette C. Gardner, Bryan R. Apell, Paige C. Thompson, Gregor B. Cadman, Ian F. Gagnon, Chris R. Frese, and Jesse F. Wechsler. Juvenile alewife passage through a compact hydropower turbine designed for fish safety. North American Journal of Fisheries Management, 43(2): 465–475, April 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winner:2022:HBF

[WSK^{+}22] Brent L. Winner, Theodore S. Switzer, Sean F. Keenan, Caleb H. Purtlebaugh, Heather Christiansen, and John Davis. A habitatbased, fishery-independent survey using actively fished hooked gear successfully characterizes reef fish populations in the Eastern Gulf of Mexico. North American Journal of Fisheries Management, 42(6):1575–1594, December 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weitkamp:2023:PLW

[WTW^{+}23] Laurie A. Weitkamp, Vanessa Tuttle, Eric J. Ward, Daniel Kamikawa, Alicia Billings, John Buchanan, and Jon E. Hess. Pacific Lamprey and Western River Lamprey marine ecology: Insight from new ocean collections. North American Journal of Fisheries Management, 43(6):1492–1510, December 2023. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2020:YMP

[WW20] Robert E. Weber and Michael J. Weber. Yearling muskellunge poststocking behavior and survival in two Central Iowa reservoirs with contrasting habitat availability. North American Journal of Fisheries Management, 40(5):1263–1275, October 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2021:EHN

[WW21] Robert E. Weber and Michael J. Weber. Effects of harvest, natural mortality, and escapement on reservoir walleye populations. North American Journal of Fisheries Management, 41(3):775– 790, June 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2020:URE

[WWBM20] Michael J. Weber, Robert E. Weber, Emily E. Ball, and Jonathan R. Meerbeek. Using radiotelemetry to evaluate poststocking survival and behavior of large fingerling walleye in three Iowa, USA, lakes. North American Journal of Fisheries Management, 40(1):48–60, February 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams:2023:ASC

[WWK^{+}23] Jesse A. Williams, Gregory W. Whitledge, Brent C. Knights, Nicholas C. Bloomfield, and James T. Lamer. Age-0 silver carp otolith microchemistry and microstructure reveal multiple early life environments and protracted spawning in the Upper Mississippi River. North American Journal of Fisheries Management, 43(1):141–153, February 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whitinger:2022:SIM

[WZG22] John A. Whitinger, Troy G. Zorn, and Brandon S. Gerig. Stable isotopes and movement of walleye change following ecological shifts driven by dreissenid mussels. North American Journal of Fisheries Management, 42(3):572–584, June 2022. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2023:FEE

[YBB^{+}23] Ryan T. Young, James C. Boase, John T. Buszkiewicz, Jan C. Dean, and Jesse T. McCarter. Field evaluation of electrofishing response thresholds for adult grass carp. North American Journal of Fisheries Management, 43(3):859–868, June 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yeager:2024:PDA

[YBH24] Joel W. Yeager, Timothy F. Bonvechio, and Martin J. Hamel. Population demographics and angler exploitation of Suwannee bass in two Georgia rivers. North American Journal of Fisheries Management, 44(2):415–427, April 2024. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yule:2020:DFC

[YDM^{+}20] Daniel L. Yule, Jamie A. Dobosenski, Jared T. Myers, Mark P. Ebener, Randall M. Claramunt, James D. McKenna, H. George Ketola, and Owen T. Gorman. Does fecundity of cisco vary in the Upper Great Lakes? North American Journal of Fisheries Management, 40(4):973–985, August 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

York:2022:OBV

[YSS22] Betsey York, Jason D. Schooley, and Dennis L. Scarnecchia. Oklahoma bowfishing values and perspectives toward nongame fishes and their management. North American Journal of Fisheries Management, 42(4):1020–1040, August 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yang:2021:RFA

[YZJ^{+}21] Zhi Yang, Qiguang Zhu, Yao Jin, Huiyuan Tang, Hong Liu, Chengyan Wan, and Xiaojuan Chen. Response of fish assemblages to habitat changes and fishing activity in a tributary of the Jinsha River in Southwest China. North American Journal of Fisheries Management, 41(4):985–998, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zendt:2023:STO

[ZAK^{+}23] Joseph S. Zendt, M. Brady Allen, Tobias J. Kock, Russell W. Perry, and Adam C. Pope. Spatial and temporal overlap between hatchery- and natural-origin steelhead and Chinook salmon during spawning in the Klickitat River, Washington, USA. North American Journal of Fisheries Management, 43(6):1687–1701, December 2023. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zheng:2021:GMA

[ZJL^{+}21] Chaochen Zheng, Tao Jiang, Renjun Luo, Xiubao Chen, Hongbo Liu, and Jian Yang. Geometric morphometric analysis of the Chinese mitten crab Eriocheir sinensis: a potential approach for geographical origin authentication. North American Journal of Fisheries Management, 41(4):891–903, August 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zorn:2022:CHC

[ZK22] Troy G. Zorn and Darren R. Kramer. Changes in habitat conditions, fish populations, and the fishery in Northern Green Bay, Lake Michigan, 1989—2019. North American Journal of Fisheries Management, 42(3):549–571, June 2022. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhang:2021:IAS

[ZLM21] Xiaohan Zhang, Adam C. Landon, and Craig A. Miller. The influence of aggregate sociodemographic characteristics on fishing license sales in Cook County, Illinois. North American Journal of Fisheries Management, 41(2):344–354, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zorn:2020:EWS

[ZMA20] Troy G. Zorn, Mark S. Mylchreest, and Arnold W. Abrahamson. Effects of white sucker removal and stocking on growth of fishes in northern lakes. North American Journal of Fisheries Management, 40(3):718–725, June 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zentner:2021:TTL

[ZSLG21] Douglas L. Zentner, Jonathan J. Spurgeon, Steve E. Lochmann, and Christy L. Graham. Tag type and location-dependent retention impart varied levels of bias on mark–recapture parameter estimates. North American Journal of Fisheries Management, 41 (2):474–483, April 2021. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zentner:2021:RFA

[ZWBS21] Douglas L. Zentner, Skylar L. Wolf, Shannon K. Brewer, and Daniel E. Shoup. A review of factors affecting PIT tag detection using mobile arrays and use of mobile antennas to detect PIT-Tagged suckers in a wadeable Ozark stream. North American Journal of Fisheries Management, 41(3):697–710, June 2021. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhao:2020:BEG

[ZZWX20] Yingming Zhao, Liang Zhang, Chunfang Wang, and Congxin Xie. Biology and ecology of grass carp in China: a review and synthesis. North American Journal of Fisheries Management, 40(6): 1379–1399, December 2020. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).